### v3

# Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection

Abstract
The advancement of large language models (LLMs) has made it difficult to differentiate human-written text from AI-generated text. Several AI-text detectors have been developed in response, which typically utilize a fixed global threshold (e.g., $\theta=0.5$) to classify machine-generated text. However, we find that one universal threshold can fail to account for subgroup-specific distributional variations. For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text than longer, and more positive classifications on neurotic writing styles than open among long text. These discrepancies can lead to misclassification that disproportionately affects certain groups. We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization in AI-generated content classifiers. Our approach partitions data into subgroups based on attributes (e.g., text length and writing style) and learns decision thresholds for each group, which enables careful balancing of performance and fairness metrics within each subgroup. In experiments with four AI text classifiers on three datasets, FairOPT enhances overall F1 score and decreases balanced error rate (BER) discrepancy across subgroups. Our framework paves the way for more robust and fair classification criteria in AI-generated output detection. We release our data, code, and project information at URL.
detection, robustness, fairness, threshold optimization
[…]
mO May[#1]
1 Introduction
The large-scale adoption of large language models (LLMs) has led to the widespread dissemination of AI-generated text that is challenging to differentiate from human-written content (Wu et al., 2025). This has raised concerns around the spread of misinformation (Chen & Shu, 2024), the degradation of publication standards (Wu et al., 2023), potential cybersecurity threats (Yao et al., 2024), and breaches of academic integrity (Perkins, 2023). As a result, effective AI content detection systems are essential for mitigating the risks from generative models. Recently, there has been substantial development in AI text detection tools, such as RoBERTa-based models (Solaiman et al., 2019), DetectGPT (Mitchell et al., 2023), Fast-DetectGPT (Bao et al., 2023), GLTR (Gehrmann et al., 2019), RADAR (Hu et al., 2023), and efforts to use LLMs as detectors (Bhattacharjee & Liu, 2024). Generally, these methods estimate a probability $p$ that the text is generated by AI by comparing the patterns (e.g., token probabilities or logits) of the given text against AI-generated text, and use a single universal threshold of $p>0.5$ for classification (Freeman & Moisen, 2008).
Figure 1: We find that using one universal probability threshold (often $\theta>0.5$) to classify AI and human-generated text can fail to account for subgroup-specific distributional variations, such as text length or writing style, leading to higher errors for certain demographics. We propose FairOPT to learn multiple decision thresholds tailored to each subgroup, which leads to more robust AI-text detection.
Figure 2: Applying adaptive thresholds to different probability distributions by subgroup, as seen through cumulative density functions (CDFs). We partition texts by three length categories (short, medium, long) and five personality traits (extroversion, neuroticism, agreeableness, conscientiousness, openness), and use RADAR (Hu et al., 2023) to infer AI-generated probabilities on the test dataset. Observe that a static classification threshold ($0.5$, in black) and a single optimized threshold (in gray at the right end side) does not account for subgroup-specific distributional variations. We used the thresholds in Appendix C. The static threshold followed conventional $0.5$, AUROC method is noted in Appendix B.
However, a single universal threshold for AI content classification fails to account for text with different characteristics, such as length or writing style, which can have very different probability distributions. That is to say, one fixed threshold does not consider the variation of the probability. For example, when using a fixed threshold of 0.5, a writing style characterized by openness and positivity is more likely to be identified as AI-generated than one characterized by neuroticism, as illustrated in Figure 2. This can lead to misclassification and disproportionately affect certain groups. Similar concerns were raised regarding race (Alghamdi et al., 2022; Bendekgey & Sudderth, 2021) and gender (Weber et al., 2020; Jang et al., 2022; Bendekgey & Sudderth, 2021).
To address these limitations, we propose FairOPT to learn decision thresholds for distinct subgroups rather than using a standard threshold (see Figure 1). Our method balances performance (e.g., ACC, F1, precision) and fairness metrics (e.g., demographic parity, equality of odds, equal opportunity) in learning these thresholds, ensuring that predefined fairness criteria are met across all subgroups in question. In experiments with four AI text classifiers on three datasets, FairOPT increases overall F1 score and decreases balanced error rate (BER) discrepancy across subgroups, yielding better overall tradeoffs between performance and fairness. To summarize, our main contributions are: (a) identifying discrepancies in AI-generated text classifiers based on length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold method to enhance group performance and fairness, and (c) developing a model with a rule-based relaxed fairness criteria to achieve fast convergence even when perfect fairness is not possible.
2 Related Work
We review work in AI-generated text detection and methods to threshold probability distributions into classifier labels.
AI text detectors
AI text detectors have been developed to flag AI-generated text. RADAR utilizes adversarial learning techniques (Hu et al., 2023), GLTR employs statistical metrics, including entropy, distributions, and ranks, to discern anomalies suggestive of AI content (Gehrmann et al., 2019), DetectGPT examines the curvature of log probability distributions, concentrating on local likelihood curvature to detect AI-generated text (Mitchell et al., 2023), and RoBERTa-based detectors fine-tune pretrained language models to proficiently categorize text as either human-produced or AI-generated (Solaiman et al., 2019). We employed probabilistic classifiers among them which require a threshold-based decision mechanism. Unlike direct discrimination methods, thresholding enables post-processing for the balancing of performance and fairness.
Fixed universal thresholds for classification:
The most prevalent approach for making decision thresholds in classification involves employing a fixed universal threshold of $0.5$ to map predicted probabilities to class labels (Freeman & Moisen, 2008). It is commonly used in many fields like photogrammetry (Shao et al., 2016), ecology (Manel et al., 1999; Hanberry & He, 2013) and computer science (Lu et al., 2024). Nevertheless, as highlighted by Freeman & Moisen (2008), the dependence on $\theta=0.5$ as a default threshold, merely due to its general acceptance, is frequently unreliable owing to variations in data distributions.
Universal thresholds with optimization:
Threshold optimization techniques have been formulated to learn optimal thresholds across entire datasets. These methods optimize metrics such as the area under the receiver operating characteristic curve (AUROC) (Bradley, 1997) and derive optimized decision thresholds under constraints such as the false positive rate (FPR) (Krishna et al., 2024; Lipton et al., 2014). These approaches find applications across disciplines, including econometrics (Staňková, 2023), statistics (Esposito et al., 2021), and machine learning (OpenAI, 2023). While these methods yield a globally optimized threshold tailored to the chosen model, they can fail to adapt to the characteristics of individual instances, which can compromise the robustness of detection systems, particularly in significantly divergent data distributions.
Adaptive thresholds across groups:
Recent work has highlighted the necessity for adaptive thresholds across groups (Jang et al., 2022; Bakker et al., 2021). Menon & Williamson (2018) propose instance-dependent thresholds for inferred probabilities, resulting in optimal classifiers under cost-sensitive fairness objectives and enhancing overall performance. Similarly, Corbett-Davies et al. (2017) demonstrate that group-specific thresholds, informed by group-level statistics, can facilitate fairness-aware classification. Canetti et al. (2019) investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity. Jang et al. (2022) present threshold-adaptation methods to ensure fair classification, and Bakker et al. (2021) explore threshold tuning to maintain stable classification performance across subgroups. Our method, FairOPT is grounded in the aforementioned principles and introduce a novel approach of adaptive thresholding.
3 Conceptual Framework and Methodology
This section outlines our method for adaptive thresholds in AI-generated text detection, FairOPT, which iteratively adjusts classification thresholds for different subgroups to balance performance with inter-group fairness.
3.1 Objective
Considering a set of subgroups $G_{1},\ldots,G_{i}$ characterized by attributes (e.g., length and style) and acknowledging that texts with varying characteristics shows discrepancies in probability distributions, our objective is to refine the binary classification thresholds $\{\theta(G_{1}),\ldots,\theta(G_{i})\}$ applicable to $G_{1},\ldots,G_{i}$. By accounting for subgroup-specific variations, these tailored thresholds ensure that decision boundaries accurately reflect the specific probability distributions.
3.1.1 Preliminaries of AI Text Detection
The detection of AI-generated text is commonly formulated as a binary classification task, where the input space $\mathcal{X}$ consists of the sequence of tokenized indices derived from a given text, and the label space $\mathcal{Y}=\{0,1\}$ denotes the two distinct classes of 00 for human-written text and $1$ for AI-generated text. To address this task, existing work primarily utilizes a pretrained neural network as a probabilistic classifier $M_{\theta}:\mathcal{X}\rightarrow[0,1]$, which assigns a probability $p$ to the input reflecting the likelihood of it being AI-generated (Solaiman et al., 2019; Hu et al., 2023). This probability is then converted into a predicted label by applying a threshold, typically set at $0.5$. Thus, the final predicted label $\hat{Y}\in\{0,1\}$ is determined as follows:
$\hat{Y}=\begin{cases}1,&\text{if }M_{\theta}(x)\geq\text{threshold},\\ 0,&\text{otherwise.}\end{cases}$ (1)
3.1.2 Performance metrics
After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Correctly classifying AI-generated content as AI constitutes a true positive (TP). If human text is not classified as AI, it is true negative (TN). Flagging human text as AI is a false positive (FP), and flagging AI text as human is a false negative (FN). Details about the classification matrix and performance metrics are in Appendix A.1.
3.1.3 Fairness metrics
The objective of fair classification is to mitigate disparities while preserving performance. To quantify fair performance, we focus on Demographic Parity (DP) (Kim et al., 2020), a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$. This condition mandates that the probability of assigning a positive outcome is equitable across all groups, regardless of the true class labels $Y$. To measure the extent of disparity, we define the DP metric:
$\Delta_{\text{DP}}$ $=\left|P(\hat{Y}=1\mid S=a)-P(\hat{Y}=1\mid S=b)\right|$ (2)
By addressing an overall equal rate of positive predictions across groups, DP seeks to measure disparate impact, where one group receives a disproportionate share of positive or negative outcomes due to biases in the classifier model (Feldman et al., 2015). In applying DP, the perfect fairness is $\Delta_{\text{DP}}=0$, which means that there is no difference in positive rates between groups. Details about fairness and performance metrics are noted in Appendix A.2.
3.2 Relaxed Fairness
While achieving 100% fairness across groups is theoretically ideal, it is often impractical due to computational constraints and the difficulty of ensuring model convergence under strict fairness criteria (Dwork et al., 2012). The 80% fairness rule, often called relaxed fairness, offers a pragmatic approach to balance reasonable fairness and performance (Feldman et al., 2015), and can also mitigate overfitting risks and ensure convergence based on the rule-based constraints.
We formalize a relaxed fairness criterion requiring that the ratio of minimum to maximum value of each fairness metric across all subgroups does not fall below a specified threshold $\tau$, set to $0.8$. It can be mathematically represented as:
$\frac{\min\limits_{i\in\mathcal{S}}\,c_{i}^{(k)}(\theta)}{\max\limits_{i\in\mathcal{S}}\,c_{i}^{(k)}(\theta)}\geq\tau,\quad\forall k\in\mathcal{K},$ (3)
where:
  • $\mathcal{S}$ denotes the set of attributes (e.g. length, writing style, etc.)
  • $\mathcal{M}$ represents the set of fairness metrics (e.g., demographic parity, equalized odds, etc.).
  • $c_{i}^{(k)}(\theta)$ computes the $k$-th metric in $\mathcal{K}$ for each group using classifier parameters $\theta$.
  • $\tau\in[0,1]$ is the fairness criteria, with $\tau=0.8$ reflecting the 80% rule and $\tau=1$ reflects 100% fairness.
Note that we should consider limiting the number of fairness metrics because of the impossibility theorem (Dwork et al., 2012) as it is computationally impractical to satisfy all fairness criteria simultaneously. The generated thresholds for subgroups based on the training dataset is used for the classification criteria in the test dataset that has same subgroups.
3.3 Group-Adaptive Threshold Optimization
Process
Given these preliminaries, we now introduce our FairOPT algorithm, which is summarized in Algorithm 1. At a high level, FairOPT uses a gradient-based subgroup threshold update to jointly optimize classification performance (ACC and F1) and group-level fairness.
Formally, we partition the dataset $D$ into subgroups $\{G_{1},\dots,G_{i}\}$ based on features $\{S_{1},\dots,S_{n}\}$. Each subgroup $G_{i}$ is assigned an initial threshold $\theta_{\mathrm{init}}$ (e.g., 0.5 to all groups).
During each iteration, we compute predictions for each sample $x_{n}\in G_{i}$ by binarizing the predicted probability $p_{n}$ using the current threshold $\theta(G_{i})$. The model derives a contingency matrix based on the predictions $\hat{y}_{n}$ and actual values $y_{n}$.
For each subgroup $G_{i}$, we define a loss function that penalizes both low accuracy and insufficient F1 score:
$L_{i}(\theta)\;=\;-\,\mathrm{ACC}_{i}(\theta)\;+\;\kappa\,\bigl{[}\beta-\mathrm{F1}_{i}(\theta)\bigr{]}_{+},$ (4)
where $\beta$ is a minimum F1 threshold, $\kappa>0$ is a penalty weight, $\alpha$ is also given for the minimum ACC, and $[\cdot]_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero. The minimum thresholds are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
To optimize each $\theta(G_{i})$, we employ a finite-difference approximation of the gradient:
$\nabla L_{i}(\theta(G_{i}))\;\approx\;\frac{L_{i}(\theta(G_{i})+\delta)-L_{i}(\theta(G_{i})-\delta)}{2\,\delta}.$ (5)
The finite-difference approximation estimates the gradient of each subgroup’s threshold-based loss function by computing the difference between the loss values at $\theta(G_{i})+\delta$ and $\theta(G_{i})-\delta$, then dividing by $2\delta$ (Liu et al., 2020). This method is employed for an effective interaction with the discrete contingency matrix, as it allows gradient estimation without requiring an explicit analytic form of the loss function’s derivative.
We then perform a gradient descent step for each subgroup:
$\theta(G_{i})\;\leftarrow\;\theta(G_{i})-\eta\,\nabla L_{i}(\theta(G_{i})),$ (6)
followed by clipping to the allowable interval $\bigl{[}a,b\bigr{]}\subseteq[0,1]$ to make sure it is a valid probability.
Algorithm 1 FairOPT: Group-Specific Thresholding
1:  Input: $D$ dataset, $S_{n}$ features, predicted probabilities $p$, performance targets ($\alpha,\beta$), fairness gap limit $\delta_{\text{fair}}$, max iterations, step size $\eta$.
2:  Split $D$ into sub-datasets $D_{1},\ldots,D_{j}$ by $S_{n}$.
3:  for iteration = 1 to max iterations do
4:     for each subgroup $G_{i}$ do
5:        Convert $p$ in $G_{i}$ to binary predictions using $\theta(G_{i})$ and compute confusion matrix.
[…]
7:           return $\{\theta(G_{1}),\ldots,\theta(G_{j})\}$.
8:        else
9:           Adjust $\theta(G_{i})$
10:  Output: $\{\theta(G_{1}),\dots,\theta(G_{j})\}$.
Finally, the model calculates domain-relevant fairness metrics $\{M_{1},\ldots,M_{k}\}$ across all subgroups by measuring the maximum inter-group disparity:
$\Delta_{k}\;=\;\max_{i}\,M_{k}(G_{i})\;-\;\min_{i}\,M_{k}(G_{i}),$ (7)
and verifies whether $\Delta_{k}\leq\delta_{\mathrm{fair}}$ for each metric $M_{k}$. If any disparity exceeds $\delta_{\mathrm{fair}}$, it is considered unfair. The evaluation for fairness is not included in the loss function because it assesses performance in a binary manner, fair or unfair (e.g., pass or fail), similar to zero-one loss, rather than providing a continuous gradient that can guide optimization.
To determine termination, we monitor the largest threshold change among all groups,
$\Delta_{\theta}\;=\;\max_{G_{i}}\,\bigl{|}\theta_{\mathrm{new}}(G_{i})-\theta_{\mathrm{old}}(G_{i})\bigr{|}.$ (8)
[…]
Outcome
By iterating this process, the algorithm 1 produces an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that strike a balance between high classification performance and bounded inter-group fairness disparities. This adaptive thresholding method is particularly effective when probability distributions differ significantly across subgroups, as it tailors the decision boundary to each subgroup’s characteristics rather than relying on a single global threshold. Additional details about the algorithm, including learning and implementation details, are included in Appendix F.1.
4 Experiments
Our experiments aim to assess the performance of FairOPT in comparison to a static threshold and a single universally optimized threshold. For training, we use the RAID and MAGE datasets, and for testing, we utilized SemEval24, MAGE, and Human v. AI 500K dataset.
4.1 Datasets
4.1.1 Train dataset
The RAID dataset (Dugan et al., 2024) encompasses text samples generated by a wide array of language models, including GPT-4, GPT-3, Llama 2 70B, Cohere, MPT-30B, and Mistral 7B. These models produced text across diverse domains such as abstracts, recipes, reddit posts, book summaries, news articles, and poetry. MAGE (Li et al., 2024) includes human-written text covering opinion statements, news articles, question answering, stories, and scientific writing, with machine-generated texts produced by 27 LLMs, including OpenAI’s GPT variants, LLaMA series, FLAN-T5, and EleutherAI models. We randomly extracted 10,000 samples from each of these two datasets.
4.1.2 Test dataset
We also use three datasets for test: SemEval24, MAGE, and Human v. AI 500K. The SemEval24 dataset (Wang et al., 2024) was employed for its diverse representation of commonly used AI text generators balanced with human-written texts across various topics such as news, social media posts, and creative writing, from which we extracted 3,000 samples. We also extracted random 3,000 writings from the MAGE dataset except for the instances that used in the training, and randomly selected 2,000 samples from the Human v. AI 500K dataset, the most widely used dataset for AI text detection on Kaggle.
4.2 AI-Generated Text Classifiers
To detect AI-generated text, we employed a combination of openly accessible classifiers, RoBERTa-base (Solaiman et al., 2019), RoBERTa-large (Solaiman et al., 2019), and RADAR (Hu et al., 2023). These classifiers were selected based on their open-source availability and the ability to produce unified probabilistic outputs ranging from 0 to 1 with seven decimal points. GPT4o-mini was incorporated for its promising application in AI content detection tasks (Bhattacharjee & Liu, 2024). Unlike binary classification approaches, these models assign a probability score indicating the likelihood that a given text was generated by AI rather than directly making a discrimination. To maintain uniformity in our evaluation, GPT4o-mini was explicitly prompted to return probability scores within the same range as other detectors.
Figure 3: Histograms show how AI-generated probability distributions differ by text length (Short in red, Medium in green, Long in blue). If we use the threshold of 0.5 (black dash line), human-written medium-length text shows higher error rate than other lengths. Each kernel density estimation (KDE) curve reflects the probability scores assigned by RoBERTa-large on the training dataset.
4.3 Subgroup Attributes
We extracted subgroup attributes from each dataset to undercover the potential disparities present in AI classifiers. These attributes are shown in Table 1 and include both non-sensitive attributes like text length and sensitive attributes inferred from stylistic traits.
Attribute Characteristics Labels
Text Length Number of characters in a text Short, medium, long (n = 3)
Stylistic personality Inferred Big Five personality traits. Openness, conscientiousness, extroversion, agreeableness, neuroticism (n = 5)
Table 1: The subgroup attributes studied in our experiments.
Figure 4: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoBERTa-large detector probabilities. The visualization is based on KDE. The biggest discrepancy is observed with a KS statistic of $0.2078$ ($p<0.01$), while the smallest is $0.1001$ ($p<0.01$). These discrepancies indicate varying levels of divergence between groups based on the characteristics of the given text.
(a) RoBERTa-large
(b) RoBERTa-base
Figure 5: The FairOPT algorithm enables us to obtain a balanced tradeoff between performance and fairness, through a relaxed fairness criterion specified by a given fairness threshold. We improve the Pareto frontier of performance and fairness compared to static and ROCFPR baselines.
Text length measures the number of characters in a text. OpenAI’s evaluations (OpenAI, 2023) revealed performance disparities between short (under 1,000 characters) and long text when using their AI classifier. Building on this and using exploratory data analysis, we refined the thresholds to 1,000 and 2,500 characters, categorizing texts as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500).
The stylistic personality attribute includes five traits: extroversion, neuroticism, agreeableness, conscientiousness, and openness (Minej, 2023). We determine the personality trait by selecting the one with the highest probability among the five using a BERT-based personality prediction model (Minej, 2023), which is fine-tuned from BERT-base-uncased (Devlin et al., 2018). Details are noted in Appendix D.
4.4 Discovery of Subgroup Bias in Text Detection
Our first experimental analysis uncovers many discrepancies in AI-text detection performance with respect to different text lengths and writing styles. For example, as shown in Figure 3, medium-length human-written text is more likely to be flagged as AI-generated than short-length human-written text. Similarly, in Figure 2, long text with a neurotic style is also more likely to be detected as AI-generated.
To obtain a holistic view of these biases, we used the Kolmogorov-Smirnov (KS) test to evaluate distributional discrepancies across various text groups under the assumption of a non-parametric distribution. As shown in Figure 4, there are notable differences between certain subgroup pairs. For instance, the comparison between texts characterized by neuroticism and medium-length versus those with neuroticism and short-length yielded a KS statistic of 0.2078 ($p<0.01$). Extroversion and mid-length and neuroticism and short length, showed a KS statistic of 0.1959 ($p<0.01$). On the other hand, relatively smaller discrepancies were observed between groups such as neuroticism and short length versus openness and short length, with a KS statistic of 0.1155 ($p<0.001$), and between neuroticism and short length versus openness and mid-length, with a KS statistic of 0.1001 ($p<0.01$). These observed discrepancies show that varying subgroup characteristics can lead to diverging outcomes when using a fixed threshold for classification. Addressing these discrepancies through tailored threshold adjustments or other interventions is crucial to ensuring equitable treatment across all subgroups.
Detector Method SemEval Kaggle MAGE
F1 ACC FairLen FairPer FPR F1 ACC FairLen FairPer FPR F1 ACC FairLen FairPer FPR
RoBERTa-large Static 0.4382 0.2847 0.1444 0.1377 0.9892 0.4309 0.2823 0.0607 0.1161 0.9830 0.7147 0.6360 0.1817 0.1760 0.6464
ROCFPR 0.0740 0.2990 0.2197 0.1792 0.4854 0.3469 0.3900 0.1490 0.5782 0.6364 0.5201 0.6475 0.3471 0.2777 0.1031
FairOPT 0.4849 0.3223 0.1347 0.1075 0.9892 0.4542 0.2973 0.0291 0.0755 0.9835 0.7088 0.6080 0.1749 0.1209 0.6473
RADAR Static 0.5592 0.6373 0.3826 0.3399 0.2266 0.7097 0.7373 0.3088 0.2876 0.3360 0.4364 0.3930 0.5979 0.4599 0.6896
ROCFPR 0.0153 0.5293 0.1162 0.0475 0.0019 0.0351 0.6333 0.1199 0.1412 0.0005 0.3215 0.4935 0.4102 0.4811 0.2662
FairOPT 0.6307 0.6620 0.3536 0.4939 0.2228 0.6445 0.6183 0.2269 0.1566 0.3732 0.4541 0.3820 0.5705 0.4140 0.6965
RoBERTa-base Static 0.3345 0.2133 0.0976 0.0978 0.9703 0.3659 0.2457 0.0950 0.1294 0.9553 0.7020 0.6515 0.2023 0.2124 0.5265
ROCFPR 0.0740 0.2253 0.2463 0.1428 0.7633 0.3374 0.2943 0.0876 0.1276 0.8171 0.5814 0.6580 0.4275 0.3666 0.1739
FairOPT 0.3548 0.2253 0.1072 0.0866 0.9696 0.3706 0.2447 0.0876 0.1234 0.9553 0.7122 0.6485 0.1966 0.1517 0.5265
GPT4o-mini Static 0.3737 0.5497 0.1806 0.3137 0.2114 0.2949 0.3083 0.1042 0.2494 0.7390 0.2017 0.4340 0.2921 0.1990 0.2878
ROCFPR 0.0004 0.5247 0.0053 0.0130 0.0057 0.0000 0.6267 0.0005 0.0020 0.0005 0.0000 0.5080 0.0064 0.0039 0.0020
FairOPT 0.4371 0.5553 0.1907 0.6844 0.2139 0.3792 0.3223 0.1279 0.3471 0.7459 0.2802 0.4170 0.2578 0.1697 0.3006
Table 2: Comparison of three thresholding methods (Static, ROCFPR, FairOPT) across four probabilistic AI text classifiers (RoBERTa-large, RoBERTa-base, RADAR, GPT4o-mini) and three data sources (SemEval, Kaggle, MAGE). Metrics include F1, ACC, FairLen (BER discrepancy for Length), FairPer (BER discrepancy for Personality), and FPR. Compared to the Static method, FairOPT results in an average increase of 2.91% in F1 score and decrease of 1.77% in FairPer, while causing a decrease of 0.58% in ACC and an increase of 1.59% in FairLen., demonstrating its effectiveness in enhancing specific aspect of performance and fairness with minimal trade-offs in classifier performance. Although ROCFPR method showed the best result in decreasing the FPR of the model, it overly sacrificed performance.
4.5 Mitigating Subgroup Bias using FairOPT
Given these findings, we now describe the experimental setup using FairOPT for more robust AI-text detection across subgroups. We extracted the text length and personality features from every item in the dataset, which enabled the partition of 15 subgroups with different characteristics. For the training dataset, the AI text classifiers RoBERTa-base, RoBERTa-large, and RADAR were employed to generate predicted probabilities indicating the likelihood that a given text was AI-generated. These probabilities ranged from 0 to 1, with no direct classification decisions made at this stage. For the test dataset, additional predictions were generated using the GPT4o-mini API.
We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup. We include all hyperparameters and training details in Appendix F.3. Threshold optimization was performed on the training dataset based on the results only from RoBERTa-large rather than averaging the probabilities of three detectors. The relaxed fairness criteria in 3.2 is employed to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness. The RoBERTa-large model was also used to implement the AUROC-based approach (Method based on ROC curve B ) with $FPR<0.1$ constraint for comparison, and the static threshold 2 does not need a model for training in generating a decision threshold.
The optimized thresholds were then applied to the test dataset. Performance was evaluated using F1, ACC, BER, and BER calculated per subgroup. FairLen and FairPer metrics were utilized to quantify the maximum discrepancies in BER across groups defined by text length and personality traits. These metrics are defined as follows:
FairLen $=\max_{a\in L}\text{BER}(a)-\min_{a\in L}\text{BER}(g)$ (9)
FairPer $=\max_{b\in P}\text{BER}(b)-\min_{b\in P}\text{BER}(g)$ (10)
where $L$ includes three groups based on text length, $P$ includes five groups based on big five personality traits, and BER denotes the balanced error rate for each group. A higher value of FairLen or FairPer indicates a greater disparity in classification performance across subgroups.
4.6 Results with FairOPT for AI-Text Detection
Table 2 presents a comprehensive comparison of three thresholding methodologies—Static, ROCFPR, and FairOPT—applied to four AI text detectors (RoBERTa-base, RoBERTa-large, RADAR, GPT4o-mini) across three distinct datasets (MAGE, Kaggle, SemEval). The evaluation metrics include F1, ACC, FairLen (maximum BER discrepancy for text length), FairPer (maximum BER discrepancy for personality traits), and FPR. Overall across all datasets and tasks tested, using FairOPT yields an average increase of 2.91% in F1 Score and a decrease of 1.77% in FairPer, thereby strictly improving the performance-fairness tradeoff. Overall accuracy was only reduced by a marginal 0.58%, indicating that FairOPT does not affect performance accuracy. These results indicate that FairOPT effectively enhances the overall classification performance and reduces fairness disparities across subgroups, outperforming the Static threshold approach by achieving higher F1 scores and lower BER discrepancies related to stylistic features.
We now analyze the results of FairOPT on each specific dataset. On the SemEval task, overall F1 scores show significant improvement, with RADAR achieving a notable increase of 7% after running our FairOPT algorithm. These findings imply that environments exhibiting patterns similar to SemEval can greatly benefit from this method. Additionally, the RoBERTa-large model exhibits a substantial reduction in discrepancy across all datasets that it was tested on - in the MAGE dataset, the discrepancy related to writing style based on personality decreases by 5%. This implies that models that work similarly to RoBERTa-large can pursue better fairness through our method. For example, RoBERTa-based models also exhibit reduced disparities in text length and style. These results provide strong evidence to support the future application of FairOPT to other models, particularly detectors currently under development, as outlined in OpenAI (2023).
4.7 Tradeoff Analysis between Performance and Fairness
Figure 5 illustrates the accuracy-fairness tradeoffs achieved by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines. The relaxed fairness criterion from Section 3.2 effectively navigates the trade-off space, enhancing the Pareto frontier between classification performance and fairness metrics. Especially, the relaxed fairness criterion leads to early termination of the algorithm before overly sacrificing ACC while still satisfying a reasonable level of fairness. In contrast, the ROCFPR baseline, while achieving lower FPRs as noted in Table 2, significantly compromises TPR thus harming overall ACC, highlighting the superiority of FairOPT in maintaining a harmonious balance between constraints and model performance.
The ROCFPR method demonstrates a notable ability to minimize the FPR across various classifiers and datasets. However, this reduction in FPR comes at a significant cost to the TPR, leading to a sacrifice of ACC as illustrated in Appendix B. This trade-off is particularly problematic as it diminishes the detector’s reliability in correctly identifying AI-generated text.
Conversely, the adaptive thresholding approach employed by FairOPT offers advantages in F1 and group fairness over both static and ROCFPR methods by tailoring decision boundaries to subgroup-specific probability distributions. This strategy ensures that each subgroup, characterized by attributes such as text length and personality traits, is evaluated with an appropriate threshold, thereby balancing the trade-off between fairness and performance. FairOPT achieves this balance by improving F1 scores and reducing FairPer discrepancies without substantially sacrificing ACC, thereby enhancing the classifier’s overall performance and addressing fairness.
5 Conclusion
This paper proposed FairOPT, based on the observation that there is a statistically significant difference in AI-generated text classifier outputs across subgroups of text (e.g. length and personality) and that this discrepancy can lead to notably higher error rates in specific groups. To address this problem, we proposed FairOPT, a new method for group-specific threshold optimization in AI-generated content classifiers without sacrificing performance. Unlike a universal threshold, which presumes well-calibrated probabilities across varied text distributions, our method effectively captures the variations of the probability distribution across different subgroups, leading to a more robust and fair classification.
While our work focuses primarily on text classifiers, the FairOPT framework can naturally extend to other modalities where detecting AI-generated content is critical for safety and copyright (e.g., image, video, and audio). This method will be highly beneficial in post-processing, guiding the model toward enhanced performance or increased fairness, contingent upon the specific requirements and deployment context. Furthermore, while our primary focus has been on improving fairness in the post-training stage, we hypothesize that the proposed method can also be integrated as an in-training fairness enhancement mechanism. Finally, other approaches for threshold optimization, including Bayesian methods and convex optimization, can also be studied in future research.
Impact Statement
This paper introduces novel theoretical frameworks and methodologies to enhance the robustness of AI-generated text detection systems. By extracting and using features for subgroups such as text length and stylistic traits, our approach addresses disparity in classification outcomes. This technology can play a crucial role in AI detection for identifying misinformation (Chen & Shu, 2024), safeguarding the integrity of publication and academic organizations (Wu et al., 2023; Perkins, 2023), and countering potential cybersecurity threats (Yao et al., 2024). Since it is crucial for these AI-content detection methods to be robust and fair across many potential users, our method takes a major step in this direction by formulating the problem setting and developing a new algorithm.
Acknowledgment
Thanks to Hengzhi Li, Megan Tjandrasuwita, David Dai, and Jeongmin Kwon, Chanakya Ekbote for constructive feedback and discussion.
Appendix A Performance and fairness metrics
A.1 Contingency table
The model can make not only true positive (TP) and true negative (TN) cases but also false positive (FP) and false negative (FN):
Classified as AI ($\hat{Y_{1}}$) Classified as non-AI ($\hat{Y_{0}}$)
AI-generated ($Y_{1}$) TP: Correct identification FN: Failed to detect AI and misidentified as human
Human-developed ($Y_{0}$) FP: Incorrectly marked human work as AI TN: Correctly identified by not classifying human work as AI
Table 3: Confusion matrix for the four possible outcomes when an AI detection system classifies either AI-generated or human-developed content.
A.2 Notation and Metrics
Given the table A.1 that denotes the outcomes in a binary classification task, we can mathematically express the performance and fairness metrics like below:
  • Accuracy (ACC).
    $\mathrm{ACC}\;=\;\frac{\mathrm{TP}+\mathrm{TN}}{\mathrm{TP}+\mathrm{TN}+\mathrm{FP}+\mathrm{FN}}.$
  • F1 Score. Given $\mathrm{Precision}$ and $\mathrm{Recall}$ below,
[…]
    F1 matters because it balances Precision and Recall into a single measure, making it especially useful when the class distribution is imbalanced or when both metrics are important.
  • Precision.
    $\mathrm{Precision}\;=\;\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}}$
    Precision measures the fraction of predicted positives that are actually correct, focusing on how often the model is right when it predicts a positive.
  • Recall.  (Recall is also referred to as TPR in some contexts.)
    $\mathrm{Recall}\;=\;\mathrm{TPR}\;=\;\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}$
    Recall measures the fraction of actual positives that are correctly identified by the model, capturing how many real positives are detected.
  • Demographic Parity (DP). Suppose we have a protected attribute $A\in\{0,1\}$. Demographic Parity requires $P(\hat{Y}=1\mid A=0)=P(\hat{Y}=1\mid A=1)$, or $P(\hat{Y}=1\mid A=0)\approx P(\hat{Y}=1\mid A=1)$. A common metric is the disparity:
    $\Delta_{\mathrm{DP}}\;=\;\bigl{|}\,P(\hat{Y}=1\mid A=0)\;-\;P(\hat{Y}=1\mid A=1)\bigr{|}.$
  • Equalized Odds (EO). Equalized Odds requires equal true positive rates and false positive rates across groups. One way to measure the gap is:
    $\Delta_{\mathrm{EO}}\;=\;\bigl{|}P(\hat{Y}=1\mid Y=1,A=0)\;-\;P(\hat{Y}=1\mid Y=1,A=1)\bigr{|}\;+\;\bigl{|}P(\hat{Y}=1\mid Y=0,A=0)\;-\;P(\hat{Y}=1\mid Y=0,A=1)\bigr{|}.$
    Under the strictest notion of fairness, one requires
[…]
    implying no disparity in true or false positive rates across groups.
    Relaxation in fairness allows
    $\Delta_{\mathrm{EO}}\geq 0,$
    for feasibility.
  • False Positive Rate (FPR).
    $\mathrm{FPR}\;=\;\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}$
  • Balanced Error Rate (BER).
    $\mathrm{BER}\;=\;\frac{1}{2}\left(\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}+\frac{\mathrm{FN}}{\mathrm{TP}+\mathrm{FN}}\right)\;=\;\frac{\mathrm{FPR}+(\mathrm{FNR})}{2}\;=\;\frac{\mathrm{FPR}+(\mathrm{1-TPR})}{2}$
    Balanced Error Rate (BER) averages the error rates across both classes, providing a metric that is insensitive to class imbalance. By considering both the False Positive Rate and the False Negative Rate, BER ensures that the classifier performs equally well on both classes. The concept is especially well-introduced in (Ferrer, 2022)
  • Specificity.
    $\mathrm{Specificity}\;=\;\frac{\mathrm{TN}}{\mathrm{TN}+\mathrm{FP}}$
    Specificity measures the proportion of actual negative cases correctly identified by the model, reflecting its ability to detect true negatives. It is also known as true negative rate (TNR).
A.3 Impossibility theorem
The impossibility theorem typically refers to the result that no single classifier can satisfy all common fairness metrics (e.g., demographic parity, equalized odds, predictive parity) simultaneously, except in trivial cases. This arises because different fairness definitions can be mathematically incompatible when the base rates (prevalences) differ across groups. This research used one fairness metrics in the processing to address this issue.
Appendix B AUROC
B.1 Static threshold and AUROC-based approach
In this research, the AUROC-based approach is employed to optimize the classification threshold while balancing model performance and fairness constraints. This method evaluates the trade-off between the TPR and FPR, as visualized by the ROC curve shown in Figure 6. The optimal threshold is determined under a specified FPR constraint, $\text{FPR}\leq 0.1$, ensuring that the selected threshold aligns with a predefined ethical constraint.
Figure 6: Threshold optimization process using the AUROC-based approach with an FPR constraint of 0.1. The ROC curve (blue line) demonstrates the trade-off between TPR and FPR. The optimal threshold (green marker) is selected to maximize performance under the specified FPR constraint.
Appendix C Applied Thresholds
Figure 2 uses thresholds that is noted on Table 4.
Threshold Type Value
Static 0.5
ROCFPR 0.9984415
Adaptive Thresholds (n=15)
Group Threshold
long_agreeableness 0.2545000
long_conscientiousness 0.3055681
long_extroversion 0.2500000
long_neuroticism 0.2748645
long_openness 0.2549773
medium_agreeableness 0.2445749
medium_conscientiousness 0.2448965
medium_extroversion 0.1000000
medium_neuroticism 0.2429920
medium_openness 0.2436330
short_agreeableness 0.2529650
short_conscientiousness 0.2480741
short_extroversion 0.2528146
short_neuroticism 0.2528743
short_openness 0.2489719
Table 4: Thresholds that used in the study
Appendix D Feature Engineering
In order to mitigate disparities within AI classifiers, a thorough feature engineering process was undertaken. Features were systematically extracted to characterize each text within the dataset. The table 5 includes all extraced features for this research. The primary focus of the experiment encompassed text length and anticipated personality traits. Although formality and sentiment were also extracted, these attributes are utilized to expedite processing, as the computational time increases exponentially with the inclusion of additional features. The subsequent table illustrates the features primarily extracted for the experiment, with an emphasis solely on the utilization of length and personality traits.
This research mainly used length and personality as a main feature in using FairOPT.
The text length attribute measures the number of characters in a text. OpenAI’s evaluations (OpenAI, 2023) revealed performance disparities between short (under 1,000 characters) and long texts when using their AI classifier. While their global threshold approach failed to address group-specific differences, it provided a valuable benchmark. Building on this and using exploratory data analysis (EDA), we refined the thresholds to 1,000 and 2,500 characters, categorizing texts as short (up to 1,000), mid-length (1,000–2,500), or long (over 2,500).
The personality attribute includes five traits: extroversion, neuroticism, agreeableness, conscientiousness, and openness. This attribute contributes to the enforcement of fairness metrics, such as demographic parity and equality of odds. For this feature, we determine the personality trait by selecting the one with the highest probability among the five using the (Minej, 2023) model, which is fine-tuned from BERT-base-uncased (Devlin et al., 2018).
Attribute Characteristics Labels
Text Length Measures the number of characters in a text short, medium, long (n = 3)
Formality Indicates the level of formality in the language formal, informal (n = 2)
Sentiment Reflects the emotional tone. positive, negative (n= 2)
Personality Represents the inferred Big Five personality traits. Openness, Conscientiousness, Extroversion, Agreeableness, Neuroticism (n = 5)
Table 5: Characteristics and labels of each attribute
Outside of the length and sentiment, the formality attribute measures a text’s formality by analyzing the frequency of specific parts of speech. According to (Heylighen & Dewaele, 2002), with formal texts typically containing a higher proportion of nouns, adjectives, articles, and prepositions, whereas informal texts tend to have a greater frequency of pronouns, adverbs, verbs, and interjections. As general rule, we set a threshold of 50 to distinguish between formal (scores above 50) and informal texts.
The sentiment attribute allows for the assessment of emotional tone and attitudes within content. AI-generated texts may display distinct sentiment patterns due to the differences in their underlying algorithms compared to human expression. To classify texts as positive or negative, we used the label output from the (DistilBERT-community, ) model, which is a fine-tuned version of DistilBERT-base-uncased (Sanh et al., 2019), specifically trained on the SST-2 dataset(Socher et al., 2013).
Appendix E Discrepancies by classifiers with additional features
E.1 Discrepancy of Probability in Subgroups
E.1.1 Discrepancy in RoBERTa-large Detector
The analysis revealed significant variability in the probability distributions between subgroup pairs, as depicted in Figure 7. Employing two-sample Kolmogorov-Smirnov (KS) statistic for two-sided tests, we assessed distributional discrepancies across various text groups characterized by distinct feature combinations, under the assumption of a non-parametric distribution. Significant differences were identified in certain subgroup comparisons, indicating notable discrepancies in their probability distributions. For example, comparing texts characterized by conscientiousness, positive sentiment, formal style, and long length with those exhibiting extroversion, negative sentiment, formal style, and long length resulted in a KS statistic of 0.7286 ($p<0.05$). Similarly, evaluating agreeableness, positive sentiment, formal style, and mid-length against conscientiousness, positive sentiment, formal style, and long length yielded a KS statistic of 0.7143 ($p<0.01$).
Conversely, smaller discrepancies were observed between neuroticism, positive sentiment, formal style, and short length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2706, $p<0.05$) and neuroticism, negative sentiment, formal style, and short length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2323, $p<0.05$).
Figure 7: Kernel Density Estimation (KDE) plots comparing probability distributions for subgroup pairs with the top two highest and lowest discrepancies based on the K-S test with statistical validity. The largest discrepancy is observed with a KS statistic of $0.7286$, while the smallest discrepancy has a KS statistic of $0.2323$.
E.1.2 Discrepancy in RoBERTa-base Detector
The analysis revealed a substantial variation in the probability distributions among pairs of subgroups, as shown in Figure 8. The KS test was applied to assess distributional differences across diverse text groupings defined by multiple feature sets, under the assumption of a non-parametric distribution. Specific subgroup comparisons highlighted significant differences, pointing to noticeable disparities in their probability distributions. For example, contrasting texts defined by agreeableness, positive sentiment, formal style, and mid-length with those featuring neuroticism, negative sentiment, informal style, and short length produced a KS statistic of 0.9 ($p<0.01$). Furthermore, the comparison between conscientiousness, negative sentiment, formal style, and long length and neuroticism, negative sentiment, informal style, and short length resulted in a KS statistic of 0.7750 ($p<0.01$).
Figure 8: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity. The largest discrepancy is observed with a KS statistic of $0.9$, while the smallest discrepancy has a KS statistic of $0.5071$. These discrepancies indicate varying levels of divergence between groups based on their characteristics.
Lesser discrepancies were noted between neuroticism, negative sentiment, informal style, and short length versus openness, positive sentiment, formal style, and long length (KS = 0.5522, $p<0.05$) and neuroticism, negative sentiment, informal style, and short length versus openness, positive sentiment, formal style, and short length (KS = 0.5071, $p<0.05$).
E.1.3 Discrepancy in RADAR Detector
The investigation disclosed considerable variability in the probability distributions across subgroup pairs, as illustrated in Figure 9. Through the application of the KS test, we evaluated distributional discrepancies among various text groups defined by distinct feature combinations, operating under the premise of a non-parametric distribution. Significant differences were discerned in particular subgroup comparisons, which highlight notable discrepancies in their probability distributions. For instance, the comparison between texts characterized by conscientiousness, negative sentiment, informal style, and long length and those exhibiting extroversion, positive sentiment, formal style, and short length yielded a KS statistic of 0.8 ($p<0.05$).
Figure 9: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity. The largest discrepancy is observed with a KS statistic of $0.8$, while the smallest discrepancy has a KS statistic of $0.1830$.
Similarly, the examination of conscientiousness, negative sentiment, informal style, and long length against neuroticism, negative sentiment, formal style, and short length resulted in a KS statistic of 0.6667 ($p<0.05$). In contrast, smaller discrepancies were detected between conscientiousness, negative sentiment, formal style, and long length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2110, $p<0.01$) and openness, negative sentiment, formal style, and long length versus openness, positive sentiment, formal style, and mid-length (KS = 0.1830, $p<0.01$).
E.1.4 Discrepancy in GPT4o-mini Detector
The analysis revealed substantial variability in the probability distributions among subgroup pairs, as depicted in Figure 10. Using the KS test, we assessed distributional discrepancies across various text groups defined by distinct feature combinations, assuming a non-parametric distribution. Significant differences were detected in particular subgroup comparisons, suggesting notable divergences in their probability distributions. For example, the comparison between texts characterized by openness, negative sentiment, informal style, and mid-length and those exhibiting openness, positive sentiment, informal style, and mid-length resulted in a KS statistic of 0.6667 ($p<0.001$). Similarly, the evaluation of extroversion, negative sentiment, formal style, and mid-length versus openness, positive sentiment, informal style, and mid-length produced a KS statistic of 0.6522 ($p<0.001$). Conversely, smaller discrepancies were noted between neuroticism, positive sentiment, formal style, and short length versus openness, negative sentiment, formal style, and long length (KS = 0.2853, $p<0.05$) and neuroticism, negative sentiment, formal style, and short length versus openness, negative sentiment, formal style, and long length (KS = 0.2560, $p<0.05$).
Figure 10: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity. The largest discrepancy is observed with a KS statistic of $0.6667$, while the smallest discrepancy has a KS statistic of $0.2560$.
Appendix F Operationalizing FairOPT
F.1 Algorithmic description
1:  Input: Labeled dataset $D=\{(x_{n},y_{n})\}$, predicted probabilities $\{p_{n}\}$, subgroup labels $\{g_{n}\}$ with possible groups $G_{1},\dots,G_{j}$, features $\{S_{1},\dots,S_{n}\}$, initial thresholds $\theta_{init}$, fairness metrics $\{M_{1},\ldots,M_{k}\}$ with acceptable disparity $\delta_{\mathrm{fair}}$, minimum performance thresholds: $\alpha$ (ACC) and $\beta$ (F1), penalty weight $\kappa$, learning rate $\eta$, tolerance $\epsilon_{\mathrm{tol}}$, maximum iterations, finite-difference step $\delta$, minimum probability for threshold $a$, maximum probability for threshold $b$
2:  Output: Subgroup thresholds $\{\theta^{*}(G_{1}),\dots,\theta^{*}(G_{j})\}$
3:  Initialization:
4:  split $D$ to ${(G_{1}),\dots,(G_{j})}$ by $S$
5:  while iteration $<$ maximum iterations do
6:    for each group $G_{j}$ do
7:       $\mathcal{I}_{i}\leftarrow\{\,n\mid g_{n}=G_{j}\}$// all samples in group $G_{j}$
8:       $\hat{y}_{n}\leftarrow[\,p_{n}\geq\theta(G_{j})\,],\;\forall n\in\mathcal{I}_{i}$
9:       Store contingency table for $G_{i}$ using $\hat{y}_{n}$ and ${y}_{n}$
10:       $\Omega\;\leftarrow\;\min_{i}\{\mathrm{ACC}(G_{j})\}\;\geq\;\alpha\;\wedge\;\min_{i}\{\mathrm{F1}(G_{j})\}\;\geq\;\beta$ // boolean output
11:       for each fairness metric $M_{k}$ do
12:         Compute $\{M_{k}(G_{1}),\dots,M_{k}(G_{j})\}$
13:         $\Delta_{k}\leftarrow\max_{i}(M_{k}(G_{j}))-\min_{i}(M_{k}(G_{j}))$
14:         $\Psi\;\leftarrow\;\Bigl{[}\,\max_{k}\Delta_{k}\;\leq\;\delta_{\mathrm{fair}}\Bigr{]}$ // boolean output
15:       for each group $G_{j}$ do
16:         $L_{i}(\theta)=-\mathrm{Acc}_{j}(\theta)+\kappa\,[\,\beta-\mathrm{F1}_{j}(\theta)\,]_{+}$
17:
$\nabla L_{i}(\theta(G_{j}))\;\approx\;\frac{L_{i}(\theta(G_{j})+\delta)\;-\;L_{i}(\theta(G_{j})-\delta)}{2\,\delta}.$
18:
$\theta(G_{j})\;\leftarrow\;\theta(G_{j})\;-\;\eta\;\nabla L_{i}(\theta(G_{j})).$
19:         Clip $\theta(G_{j})$ to stay within $[a,b]$
20:    $\Delta_{\theta}\;\leftarrow\;\max_{G_{j}}\bigl{|}\theta(G_{j})-\text{previousThresholds}[G_{j}]\bigr{|}$
21:    if ($\Delta_{\theta}<\epsilon_{\mathrm{tol}}\;\wedge\;\Omega\;\wedge\;\Psi$then
22:       break
[…]
24:    $\text{iteration}\leftarrow\text{iteration}+1$
25:  return $\{\theta(G_{1}),\dots,\theta(G_{j})\}$
Algorithm 2 FairOPT: Gradient-based adaptive threshold optimization with fairness for subgroups
F.2 Early stopping method for the tradeoff curve
An optimization algorithm was implemented using a function from the scipy.optimize library, aiming to determine probability thresholds that satisfy specific fairness constraints (acceptable_disparity = [1.00, 0.30, 0.25, 0.21. 0.20, 0.19, 0.15, 0.10]) while ensuring minimum performance thresholds for both performance criteria (ACC and F1). The optimization in the minimize function was performed using the L-BFGS-B method. Early stopping is facilitated by the default values of ftol and gtol in L-BFGS-B. Additionally, the learning rate was set to $10^{-3}$ to enable controlled and stable updates during optimization, and a penalty parameter was set to $10$ to strengthen the L2 penalty in the loss function. Furthermore, minimum performance thresholds for accuracy and F1-score were enforced based on the performance observed for each group when applying a static threshold of 0.5 and a ROC-based threshold. As stricter fairness constraints are imposed, these minimum thresholds are gradually relaxed.
F.3 Hyperparameter settings for the FairOPT
The application of the FairOPT is configured with a set of hyperparameters designed to ensure a balance between fairness and performance during the threshold optimization process. The learning rate is set to $10^{-3}$ to enable controlled and stable updates during optimization. The maximum number of iterations is specified as $10^{5}$ to provide sufficient time for the optimization algorithm to converge. An acceptable disparity of 0.2 is defined to regulate subgroup fairness, ensuring that performance differences across subgroups remain within an acceptable range through the relaxed fairness. Minimum thresholds for accuracy (0.25) and F1-score (0.25) are enforced to maintain baseline model performance and stability, particularly in cases of imbalanced datasets. A tolerance level of $10^{-2}$ is introduced to impose convergence criteria, ensuring the optimization process halts only when the changes in the objective function are negligible. Furthermore, a penalty parameter of 20 is applied to enforce performance constraints by penalizing significant performance disparities across subgroups. For fairness criteria, Demographic Parity (DP) was employed to test whether the generated thresholds meets the expected fairness. Also, this experiment used early stopping method for effective processing. The early stopping check through the change of the ACC, F1, and maximum DP discrepancy of each group, and terminates if all of them shows minimal change based on predefined patience.

### v4

@@
 Several AI-text detectors have been developed in response, which typically utilize a fixed global threshold (e.g., $\theta=0.5$) to classify machine-generated text.
-However, we find that one universal threshold can fail to account for subgroup-specific distributional variations.
-For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text than longer, and more positive classifications on neurotic writing styles than open among long text.
-These discrepancies can lead to misclassification that disproportionately affects certain groups.
-We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization in AI-generated content classifiers.
-Our approach partitions data into subgroups based on attributes (e.g., text length and writing style) and learns decision thresholds for each group, which enables careful balancing of performance and fairness metrics within each subgroup.
-In experiments with four AI text classifiers on three datasets, FairOPT enhances overall F1 score and decreases balanced error rate (BER) discrepancy across subgroups.
-Our framework paves the way for more robust and fair classification criteria in AI-generated output detection.
+However, one universal threshold could fail to account for distributional variations by subgroups.
+For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text, and more positive classifications of neurotic writing styles among long texts.
+These discrepancies can lead to misclassifications that disproportionately affect certain groups.
+We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization for probabilistic AI-text detectors.
+We partitioned data into subgroups based on attributes (e.g., text length and writing style) and implemented FairOPT to learn decision thresholds for each group to reduce discrepancy.
+In experiments with nine AI text classifiers on three datasets, FairOPT decreases overall balanced error rate (BER) discrepancy by 12% while minimally sacrificing accuracy by 0.003%.
+Our framework paves the way for more robust classification in AI-generated content detection via post-processing.
 We release our data, code, and project information at URL.
-detection, robustness, fairness, threshold optimization
-[…]
-mO May[#1]
 1 Introduction
-The large-scale adoption of large language models (LLMs) has led to the widespread dissemination of AI-generated text that is challenging to differentiate from human-written content (Wu et al., 2025).
-This has raised concerns around the spread of misinformation (Chen & Shu, 2024), the degradation of publication standards (Wu et al., 2023), potential cybersecurity threats (Yao et al., 2024), and breaches of academic integrity (Perkins, 2023).
-As a result, effective AI content detection systems are essential for mitigating the risks from generative models.
-Recently, there has been substantial development in AI text detection tools, such as RoBERTa-based models (Solaiman et al., 2019), DetectGPT (Mitchell et al., 2023), Fast-DetectGPT (Bao et al., 2023), GLTR (Gehrmann et al., 2019), RADAR (Hu et al., 2023), and efforts to use LLMs as detectors (Bhattacharjee & Liu, 2024).
-Generally, these methods estimate a probability $p$ that the text is generated by AI by comparing the patterns (e.g., token probabilities or logits) of the given text against AI-generated text, and use a single universal threshold of $p>0.5$ for classification (Freeman & Moisen, 2008).
+The copious use of large language models (LLMs) facilitated widespread dissemination of AI-generated text [1, 2] and has raised concerns around the dissemination of misinformation [3], the degradation of publication standards [4], cybersecurity threats [5], and breaches of academic integrity [6].
+In response to the anticipated risks, AI detection techniques have been proposed [1, 7].
+Typically, these methods estimate a probability $p$ that the text is generated by AI by comparing the patterns (e.g., token probabilities) of the given text against AI-generated text, and use a universal threshold (e.g., $\theta>0.5$[8] for a binary classification.
 Figure 1: We find that using one universal probability threshold (often $\theta>0.5$) to classify AI and human-generated text can fail to account for subgroup-specific distributional variations, such as text length or writing style, leading to higher errors for certain demographics.
-We propose FairOPT to learn multiple decision thresholds tailored to each subgroup, which leads to more robust AI-text detection.
-Figure 2: Applying adaptive thresholds to different probability distributions by subgroup, as seen through cumulative density functions (CDFs).
-We partition texts by three length categories (short, medium, long) and five personality traits (extroversion, neuroticism, agreeableness, conscientiousness, openness), and use RADAR (Hu et al., 2023) to infer AI-generated probabilities on the test dataset.
-Observe that a static classification threshold ($0.5$, in black) and a single optimized threshold (in gray at the right end side) does not account for subgroup-specific distributional variations.
-We used the thresholds in Appendix C.
-The static threshold followed conventional $0.5$, AUROC method is noted in Appendix B.
-However, a single universal threshold for AI content classification fails to account for text with different characteristics, such as length or writing style, which can have very different probability distributions.
-That is to say, one fixed threshold does not consider the variation of the probability.
-For example, when using a fixed threshold of 0.5, a writing style characterized by openness and positivity is more likely to be identified as AI-generated than one characterized by neuroticism, as illustrated in Figure 2.
-This can lead to misclassification and disproportionately affect certain groups.
-Similar concerns were raised regarding race (Alghamdi et al., 2022; Bendekgey & Sudderth, 2021) and gender (Weber et al., 2020; Jang et al., 2022; Bendekgey & Sudderth, 2021).
-To address these limitations, we propose FairOPT to learn decision thresholds for distinct subgroups rather than using a standard threshold (see Figure 1).
-Our method balances performance (e.g., ACC, F1, precision) and fairness metrics (e.g., demographic parity, equality of odds, equal opportunity) in learning these thresholds, ensuring that predefined fairness criteria are met across all subgroups in question.
-In experiments with four AI text classifiers on three datasets, FairOPT increases overall F1 score and decreases balanced error rate (BER) discrepancy across subgroups, yielding better overall tradeoffs between performance and fairness.
-To summarize, our main contributions are: (a) identifying discrepancies in AI-generated text classifiers based on length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold method to enhance group performance and fairness, and (c) developing a model with a rule-based relaxed fairness criteria to achieve fast convergence even when perfect fairness is not possible.
+We propose FairOPT to learn multiple decision thresholds tailored to each subgroup, which reduces discrepancy in AI text detection as shown in Figure 7.
+Figure 2: Applying adaptive thresholds to different probability distributions by subgroup in the test dataset, as seen through cumulative density functions (CDFs).
+We partition texts by three length categories and five personality traits as noted in Table 1.
+Observe that a static classification threshold ($0.5$, in black) and a single optimized threshold using AUROC-based method (in gray) do not account for subgroup-specific distributional variations.
+We used fifteen thresholds that have been generated using FairOPT to address the varying probabilistic distribution from AI text classifiers.
+The training is based on the RoBERTa-large detector.
+Details about the thresholds are in Appendix E.
+However, a single universal threshold for AI content classification neglects the heterogeneity of probability distributions by textual characteristics, such as length and writing style.
+To put in another way, one fixed threshold does not consider the variation of the probability from various stylistic features.
+For example, when using a fixed threshold, a writing style characterized by long neurotic writing could be more likely to be identified as AI-generated than one characterized by agreeableness, resulting in disproportionate misclassifications, affecting certain groups.
+Analogous disparities have been documented in other classification tasks, regarding race [9, 10] and gender [11, 12, 10].
+To overcome these limitations, we introduce FairOPT, a group-adaptive thresholding framework that learns distinct decision thresholds for subgroups rather than using methods with one universal threshold (see Figure 1).
+Our method balances performance (e.g., accuracy (ACC), F1, precision) and fairness metrics (e.g., demographic parity, or equality of odds) in learning thresholds.
+Experiment on nine AI text detectors on three benchmark datasets demonstrates that FairOPT substantially reduces BER disparity across subgroups while incurring only a marginal tradeoff in overall performance.
+To summarize, our main contributions are: (a) identifying significant discrepancies in AI-generated text classifiers based on text length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold that calibrates the decision boundaries to reduce group discrepancy, and (c) showing the effectiveness of the FairOPT, achieving a 12% reduction in discrepancy while showing negligible tradeoff in performance by dropping less than 0.1% of ACC.
 2 Related Work
-We review work in AI-generated text detection and methods to threshold probability distributions into classifier labels.
-AI text detectors
-AI text detectors have been developed to flag AI-generated text.
-RADAR utilizes adversarial learning techniques (Hu et al., 2023), GLTR employs statistical metrics, including entropy, distributions, and ranks, to discern anomalies suggestive of AI content (Gehrmann et al., 2019), DetectGPT examines the curvature of log probability distributions, concentrating on local likelihood curvature to detect AI-generated text (Mitchell et al., 2023), and RoBERTa-based detectors fine-tune pretrained language models to proficiently categorize text as either human-produced or AI-generated (Solaiman et al., 2019).
-We employed probabilistic classifiers among them which require a threshold-based decision mechanism.
-Unlike direct discrimination methods, thresholding enables post-processing for the balancing of performance and fairness.
-Fixed universal thresholds for classification:
-The most prevalent approach for making decision thresholds in classification involves employing a fixed universal threshold of $0.5$ to map predicted probabilities to class labels (Freeman & Moisen, 2008).
-It is commonly used in many fields like photogrammetry (Shao et al., 2016), ecology (Manel et al., 1999; Hanberry & He, 2013) and computer science (Lu et al., 2024).
-Nevertheless, as highlighted by Freeman & Moisen (2008), the dependence on $\theta=0.5$ as a default threshold, merely due to its general acceptance, is frequently unreliable owing to variations in data distributions.
-Universal thresholds with optimization:
-Threshold optimization techniques have been formulated to learn optimal thresholds across entire datasets.
-These methods optimize metrics such as the area under the receiver operating characteristic curve (AUROC) (Bradley, 1997) and derive optimized decision thresholds under constraints such as the false positive rate (FPR) (Krishna et al., 2024; Lipton et al., 2014).
-These approaches find applications across disciplines, including econometrics (Staňková, 2023), statistics (Esposito et al., 2021), and machine learning (OpenAI, 2023).
-While these methods yield a globally optimized threshold tailored to the chosen model, they can fail to adapt to the characteristics of individual instances, which can compromise the robustness of detection systems, particularly in significantly divergent data distributions.
-Adaptive thresholds across groups:
-Recent work has highlighted the necessity for adaptive thresholds across groups (Jang et al., 2022; Bakker et al., 2021).
-Menon & Williamson (2018) propose instance-dependent thresholds for inferred probabilities, resulting in optimal classifiers under cost-sensitive fairness objectives and enhancing overall performance.
+AI text detectors:
+AI text detectors have been developed to tackle issues related to misinformation, integrity, and security.
+RADAR utilizes adversarial learning techniques [13], GLTR employs statistical metrics, including entropy, probability distribution over the model’s vocabulary for each next token (i.e.
+the softmax scores across all candidate tokens), and records the rank position of the actual chosen token within that sorted probability to discern [14], DetectGPT examines the curvature of log probability distributions of the referred LLM, concentrating on local likelihood curvature to detect AI-generated text [15], and RoBERTa-based detectors fine-tune pretrained language models [16].
+We employed probabilistic classifiers for an AI-likelihood score that use a threshold for classification rather than architectures that generate a direct discrimination.
+Static thresholding for classification:
+The canonical thresholding in probabilistic classification is using a threshold of $0.5$ to map the probabilities to labels [8].
+This default rule is commonly used in many fields like photogrammetry [17], ecology [18, 19], and machine learning [20].
+Nevertheless, as highlighted by Freeman and Moisen [8], the dependence solely on $\theta=0.5$ as a default threshold, merely due to its conventional acceptance, could be unreliable because of variations in data.
+Universal thresholding with optimization:
+Threshold optimization techniques have been formulated to learn an optimal threshold across the dataset.
+These methods optimize classifiers’ performance metrics, such as the True Positive rate (TPR) and false positive rate (FPR) [21, 22], and the most representative method is optimizing using the receiver operating characteristic curve (AUROC) [23].
+These approaches find applications across disciplines, including econometrics [24], statistics [25], and machine learning [26].
+While the AUROC-based method yields a globally optimized threshold tailored to the chosen model, it can lead to over-fitting or fail to adapt to other models.
+Adaptive thresholding across groups:
+Several studies advocate for group-adaptive decision thresholds to mitigate disparities [12, 27].
+Menon and Williamson [28] propose instance-dependent thresholds for inferred probabilities, resulting in reduced discrepancy.
 Similarly, Corbett-Davies et al.
-(2017) demonstrate that group-specific thresholds, informed by group-level statistics, can facilitate fairness-aware classification.
+[29] demonstrate that group-specific thresholds can facilitate fairness-aware classification.
 Canetti et al.
-(2019) investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity.
+[30] investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity.
 Jang et al.
-(2022) present threshold-adaptation methods to ensure fair classification, and Bakker et al.
-(2021) explore threshold tuning to maintain stable classification performance across subgroups.
-Our method, FairOPT is grounded in the aforementioned principles and introduce a novel approach of adaptive thresholding.
+[12] present threshold-adaptation methods to ensure fair classification, and Bakker et al.
+[27] explore threshold tuning to maintain stable classification performance.
+Building on these foundations, our framework FairOPT introduces a novel group-adaptive thresholding algorithm that jointly optimizes performance and fairness objectives.
+Figure 3: Human-written short texts have a higher FPR compared to other text lengths, given the static threshold, showing the inadequacy of the static threshold to variations.
+Curves reflect the average probability of AI detectors in the training dataset.
 3 Conceptual Framework and Methodology
-This section outlines our method for adaptive thresholds in AI-generated text detection, FairOPT, which iteratively adjusts classification thresholds for different subgroups to balance performance with inter-group fairness.
+This section outlines our method for adaptive thresholding, FairOPT, which iteratively adjusts classification thresholds for different subgroups to reduce discrepancy.
 3.1 Objective
-Considering a set of subgroups $G_{1},\ldots,G_{i}$ characterized by attributes (e.g., length and style) and acknowledging that texts with varying characteristics shows discrepancies in probability distributions, our objective is to refine the binary classification thresholds $\{\theta(G_{1}),\ldots,\theta(G_{i})\}$ applicable to $G_{1},\ldots,G_{i}$.
-By accounting for subgroup-specific variations, these tailored thresholds ensure that decision boundaries accurately reflect the specific probability distributions.
+Consider a set of subgroups $G_{1},\ldots,G_{i}$ with attributes (e.g., length and style).
+Acknowledging that texts with varying characteristics show variations in probability distributions, our objective is to find optimal thresholds $\{\theta(G_{1}),\ldots,\theta(G_{i})\}$ for binary classification.
 3.1.1 Preliminaries of AI Text Detection
-The detection of AI-generated text is commonly formulated as a binary classification task, where the input space $\mathcal{X}$ consists of the sequence of tokenized indices derived from a given text, and the label space $\mathcal{Y}=\{0,1\}$ denotes the two distinct classes of 00 for human-written text and $1$ for AI-generated text.
-To address this task, existing work primarily utilizes a pretrained neural network as a probabilistic classifier $M_{\theta}:\mathcal{X}\rightarrow[0,1]$, which assigns a probability $p$ to the input reflecting the likelihood of it being AI-generated (Solaiman et al., 2019; Hu et al., 2023).
+The detection of AI-generated text is commonly formulated as a probabilistic classification for a binary classification, where the input space $\mathcal{X}$ consists of the sequence of tokens derived from a given text, and the label space $\mathcal{Y}=\{0,1\}$ denotes the two distinct classes of 00 for human-written and $1$ for AI-generated.
+To address this task, existing work primarily utilizes a pretrained neural network as a probabilistic classifier $M_{\theta}:\mathcal{X}\rightarrow[0,1]$, which assigns a probability $p$ to the input reflecting the likelihood of it being AI-generated [16, 13].
 This probability is then converted into a predicted label by applying a threshold, typically set at $0.5$.
-Thus, the final predicted label $\hat{Y}\in\{0,1\}$ is determined as follows:
+Thus, the predicted label, AI-written, $\hat{Y}\in\{0,1\}$ is determined as follows:
 $\hat{Y}=\begin{cases}1,&\text{if }M_{\theta}(x)\geq\text{threshold},\\ 0,&\text{otherwise.}\end{cases}$ (1)
 3.1.2 Performance metrics
-After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$.
-Correctly classifying AI-generated content as AI constitutes a true positive (TP).
-If human text is not classified as AI, it is true negative (TN).
-Flagging human text as AI is a false positive (FP), and flagging AI text as human is a false negative (FN).
-Details about the classification matrix and performance metrics are in Appendix A.1.
+After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Details about the confusion matrix and performance metrics are in Appendix B.1.
 3.1.3 Fairness metrics
-The objective of fair classification is to mitigate disparities while preserving performance.
-To quantify fair performance, we focus on Demographic Parity (DP) (Kim et al., 2020), a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$.
-This condition mandates that the probability of assigning a positive outcome is equitable across all groups, regardless of the true class labels $Y$.
-To measure the extent of disparity, we define the DP metric:
+To quantify disparity, we use Demographic Parity (DP) [31] and Balanced Error Rate (BER) [32].
+DP is a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$.
+This condition mandates that the probability of assigning a positive outcome is equal across all groups.
+We define the DP metric:
 $\Delta_{\text{DP}}$ $=\left|P(\hat{Y}=1\mid S=a)-P(\hat{Y}=1\mid S=b)\right|$ (2)
-By addressing an overall equal rate of positive predictions across groups, DP seeks to measure disparate impact, where one group receives a disproportionate share of positive or negative outcomes due to biases in the classifier model (Feldman et al., 2015).
-In applying DP, the perfect fairness is $\Delta_{\text{DP}}=0$, which means that there is no difference in positive rates between groups.
-Details about fairness and performance metrics are noted in Appendix A.2.
+DP seeks to measure disparate impact, where one group receives disproportionate results [33], by addressing an overall equal rate of positive predictions, encompassing TP and TN, across groups.
+In applying DP, the perfect fairness is $\Delta_{\text{DP}}=0$, which means no difference in overall TP and TN between groups.
+In our subgroup-adaptive evaluation, we use BER to equally measure the type I and II errors, which are FP and FN, while including positive predictions from DP.
+For each subgroup $G_{i}$ we compute its BERi to measure the discrepancy, and the lower discrepancy indicates more equitable classification across groups.
+$\mathrm{BER}=\tfrac{1}{2}\Bigl{(}\underbrace{\tfrac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}}_{\mathrm{FPR}}\;+\;\underbrace{\tfrac{\mathrm{FN}}{\mathrm{FN}+\mathrm{TP}}}_{\mathrm{FNR}}\Bigr{)}$ (3)
+$\text{Discrepancy}=\max\mathrm{BER}_{i}\;-\;\min\mathrm{BER}_{i}$ (4)
+Details about fairness metrics are noted in Appendix B.2.
+Figure 4: The FairOPT algorithm enables us to obtain a balance in the tradeoff between performance and fairness, through a relaxed fairness criterion discussed in Section 3.2.
+Details about the tradeoff are in Appendix H.2.
 3.2 Relaxed Fairness
-While achieving 100% fairness across groups is theoretically ideal, it is often impractical due to computational constraints and the difficulty of ensuring model convergence under strict fairness criteria (Dwork et al., 2012).
-The 80% fairness rule, often called relaxed fairness, offers a pragmatic approach to balance reasonable fairness and performance (Feldman et al., 2015), and can also mitigate overfitting risks and ensure convergence based on the rule-based constraints.
-We formalize a relaxed fairness criterion requiring that the ratio of minimum to maximum value of each fairness metric across all subgroups does not fall below a specified threshold $\tau$, set to $0.8$.
-It can be mathematically represented as:
+While achieving $100\%$ fairness across groups (e.g., $DP=0$ across fifteen groups) is theoretically ideal, it is often impractical due to computational constraints and the difficulty of ensuring model convergence under strict fairness criteria [34].
+The $80\%$, relaxed fairness, rule has been suggested to offer a pragmatic approach (e.g., allowing $DP=0.2$ in achieving fairness) [33], which offers a more feasible fairness criterion and mitigates overfitting risks.
+We formalize a relaxed fairness criterion requiring that the ratio of minimum to maximum value of each fairness metric across all subgroups does not fall below a given flexibility $\tau$, set to $0.8$:
 $\frac{\min\limits_{i\in\mathcal{S}}\,c_{i}^{(k)}(\theta)}{\max\limits_{i\in\mathcal{S}}\,c_{i}^{(k)}(\theta)}\geq\tau,\quad\forall k\in\mathcal{K},$ (5)
@@
   • $c_{i}^{(k)}(\theta)$ computes the $k$-th metric in $\mathcal{K}$ for each group using classifier parameters $\theta$.
-  • $\tau\in[0,1]$ is the fairness criteria, with $\tau=0.8$ reflecting the 80% rule and $\tau=1$ reflects 100% fairness.
-Note that we should consider limiting the number of fairness metrics because of the impossibility theorem (Dwork et al., 2012) as it is computationally impractical to satisfy all fairness criteria simultaneously.
-The generated thresholds for subgroups based on the training dataset is used for the classification criteria in the test dataset that has same subgroups.
-3.3 Group-Adaptive Threshold Optimization
-Process
-Given these preliminaries, we now introduce our FairOPT algorithm, which is summarized in Algorithm 1.
-At a high level, FairOPT uses a gradient-based subgroup threshold update to jointly optimize classification performance (ACC and F1) and group-level fairness.
-Formally, we partition the dataset $D$ into subgroups $\{G_{1},\dots,G_{i}\}$ based on features $\{S_{1},\dots,S_{n}\}$.
-Each subgroup $G_{i}$ is assigned an initial threshold $\theta_{\mathrm{init}}$ (e.g., 0.5 to all groups).
-During each iteration, we compute predictions for each sample $x_{n}\in G_{i}$ by binarizing the predicted probability $p_{n}$ using the current threshold $\theta(G_{i})$.
-The model derives a contingency matrix based on the predictions $\hat{y}_{n}$ and actual values $y_{n}$.
-For each subgroup $G_{i}$, we define a loss function that penalizes both low accuracy and insufficient F1 score:
-$L_{i}(\theta)\;=\;-\,\mathrm{ACC}_{i}(\theta)\;+\;\kappa\,\bigl{[}\beta-\mathrm{F1}_{i}(\theta)\bigr{]}_{+},$ (4)
-where $\beta$ is a minimum F1 threshold, $\kappa>0$ is a penalty weight, $\alpha$ is also given for the minimum ACC, and $[\cdot]_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero.
-The minimum thresholds are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
-To optimize each $\theta(G_{i})$, we employ a finite-difference approximation of the gradient:
-$\nabla L_{i}(\theta(G_{i}))\;\approx\;\frac{L_{i}(\theta(G_{i})+\delta)-L_{i}(\theta(G_{i})-\delta)}{2\,\delta}.$ (5)
-The finite-difference approximation estimates the gradient of each subgroup’s threshold-based loss function by computing the difference between the loss values at $\theta(G_{i})+\delta$ and $\theta(G_{i})-\delta$, then dividing by $2\delta$ (Liu et al., 2020).
-This method is employed for an effective interaction with the discrete contingency matrix, as it allows gradient estimation without requiring an explicit analytic form of the loss function’s derivative.
-We then perform a gradient descent step for each subgroup:
-$\theta(G_{i})\;\leftarrow\;\theta(G_{i})-\eta\,\nabla L_{i}(\theta(G_{i})),$ (6)
-followed by clipping to the allowable interval $\bigl{[}a,b\bigr{]}\subseteq[0,1]$ to make sure it is a valid probability.
+  • $\tau\in[0,1]$ is the fairness criteria, with $\tau=1$ reflects 100% fairness, which is the most strict.
+We limited the number of fairness metrics due to impossibility theorem [34] to make it computationally feasible, as it is impractical to satisfy all fairness criteria simultaneously.
+Algorithm 1 FairOPT: Group-Specific Thresholding
 Algorithm 1 FairOPT: Group-Specific Thresholding
@@
 9:           Adjust $\theta(G_{i})$
-10:  Output: $\{\theta(G_{1}),\dots,\theta(G_{j})\}$.
-Finally, the model calculates domain-relevant fairness metrics $\{M_{1},\ldots,M_{k}\}$ across all subgroups by measuring the maximum inter-group disparity:
+10:        end if
+11:     end for
+12:  end for
+13:  Output: $\{\theta(G_{1}),\dots,\theta(G_{j})\}$.
+3.3 Group-Adaptive Threshold Optimization
+Process
+Given this setup, we introduce FairOPT  which is described in Algorithm 1.
+FairOPT uses a gradient-based subgroup threshold update to jointly optimize group-level fairness and classification performance.
+Formally, we partition the dataset $D$ into subgroups $\{G_{1},\dots,G_{i}\}$ based on features $\{S_{1},\dots,S_{n}\}$.
+Each subgroup $G_{i}$ is assigned an initial threshold $\theta_{\mathrm{init}}$ (e.g., 0.5 to all groups).
+During each iteration, we compute predictions for each sample $x_{n}\in G_{i}$ by binarizing the predicted probability $p_{n}$ using the current threshold $\theta(G_{i})$.
+The model derives a contingency matrix based on the predictions $\hat{y}_{n}$ and actual values $y_{n}$.
+For each subgroup $G_{i}$, we define a loss function that penalizes both low accuracy and insufficient F1 score:
+$L_{i}(\theta)\;=\;-\,\mathrm{ACC}_{i}(\theta)\;+\;\kappa\,\bigl{[}\beta-\mathrm{F1}_{i}(\theta)\bigr{]}_{+},$ (6)
+where $\beta$ is a minimum F1, $\kappa>0$ is a penalty weight, $\alpha$ is given for the minimum ACC, and $[\cdot]_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero.
+These minimum standards are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
+To optimize each $\theta(G_{i})$, we employ a finite-difference approximation of the gradient:
+$\nabla L_{i}(\theta(G_{i}))\;\approx\;\frac{L_{i}(\theta(G_{i})+\delta)-L_{i}(\theta(G_{i})-\delta)}{2\,\delta}.$ (7)
+The finite-difference approximation estimates the gradient of each subgroup’s threshold-based loss function by computing the difference between the two loss values at $\theta(G_{i})+\delta$ and $\theta(G_{i})-\delta$, then dividing by $2\delta$ [35].
+We then perform a gradient descent step for each subgroup:
+$\theta(G_{i})\;\leftarrow\;\theta(G_{i})-\eta\,\nabla L_{i}(\theta(G_{i})),$ (8)
+followed by clipping to $\bigl{[}a,b\bigr{]}\subseteq[0,1]$ to ensure a valid probability.
+Finally, the model calculates the given fairness metric(s) $\{M_{1},\ldots,M_{k}\}$ across all subgroups by measuring the inter-group disparity:
 $\Delta_{k}\;=\;\max_{i}\,M_{k}(G_{i})\;-\;\min_{i}\,M_{k}(G_{i}),$ (9)
 and verifies whether $\Delta_{k}\leq\delta_{\mathrm{fair}}$ for each metric $M_{k}$.
-If any disparity exceeds $\delta_{\mathrm{fair}}$, it is considered unfair.
-The evaluation for fairness is not included in the loss function because it assesses performance in a binary manner, fair or unfair (e.g., pass or fail), similar to zero-one loss, rather than providing a continuous gradient that can guide optimization.
+If any disparity exceeds $\delta_{\mathrm{fair}}$ (e.g., $0.2$), it is considered unfair.
+The fairness evaluation is not included in the main loss function because the discrepancy is evaluated in a binary manner: fair or unfair (e.g., pass or fail), similar to zero-one loss.
 To determine termination, we monitor the largest threshold change among all groups,
@@
 Outcome
-By iterating this process, the algorithm 1 produces an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that strike a balance between high classification performance and bounded inter-group fairness disparities.
-This adaptive thresholding method is particularly effective when probability distributions differ significantly across subgroups, as it tailors the decision boundary to each subgroup’s characteristics rather than relying on a single global threshold.
-Additional details about the algorithm, including learning and implementation details, are included in Appendix F.1.
+By iterating this process, the Algorithm 1 converges to an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that reduces the disparity while satisfying the demanded performance criteria.
+For full implementation details and hyperparameter specifications, refer to Appendix H.1.
+Figure 5: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoBERTa-large detector results in cumulative probability distribution function (CDF; y-axis) based on the training dataset in the range of 0.1 to 0.9 (x-axis).
+The biggest discrepancy is observed with a KS statistic of $0.3081$ ($p<0.01$), while the smallest is $0.0842$ ($p<0.01$).
+Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions by stylistic features.
+Figure 6: Threshold optimization based on the AUROC-based method under an $FPR\leq 10\%$ constraint (see Appendix C).
+The ROC curve (blue line) illustrates the trade-off between TPR and FPR, and the optimal threshold (green marker) maximizes TPR under the specified FPR constraint.
 4 Experiments
-Our experiments aim to assess the performance of FairOPT in comparison to a static threshold and a single universally optimized threshold.
-For training, we use the RAID and MAGE datasets, and for testing, we utilized SemEval24, MAGE, and Human v.
-AI 500K dataset.
-4.1 Datasets
-4.1.1 Train dataset
-The RAID dataset (Dugan et al., 2024) encompasses text samples generated by a wide array of language models, including GPT-4, GPT-3, Llama 2 70B, Cohere, MPT-30B, and Mistral 7B.
-These models produced text across diverse domains such as abstracts, recipes, reddit posts, book summaries, news articles, and poetry.
-MAGE (Li et al., 2024) includes human-written text covering opinion statements, news articles, question answering, stories, and scientific writing, with machine-generated texts produced by 27 LLMs, including OpenAI’s GPT variants, LLaMA series, FLAN-T5, and EleutherAI models.
-We randomly extracted 10,000 samples from each of these two datasets.
-4.1.2 Test dataset
-We also use three datasets for test: SemEval24, MAGE, and Human v.
-AI 500K.
-The SemEval24 dataset (Wang et al., 2024) was employed for its diverse representation of commonly used AI text generators balanced with human-written texts across various topics such as news, social media posts, and creative writing, from which we extracted 3,000 samples.
-We also extracted random 3,000 writings from the MAGE dataset except for the instances that used in the training, and randomly selected 2,000 samples from the Human v.
-AI 500K dataset, the most widely used dataset for AI text detection on Kaggle.
+4.1 Dataset
+Our experiments aim to assess the performance of FairOPT in comparison to the static and the AUROC-based threshold optimization methods The dataset is constructed with the RAID [36], MAGE [37], and SemEval24 [38] with an even balance between human and AI text to avoid class imbalance.
+The generator includes GPT and Text-Davinci variants for their common usage, and the dataset spans a diverse range of writing domains, including social media (e.g., Reddit, ELI5), encyclopedic and academic sources (e.g., Wikipedia, arXiv), instructional and creative content (e.g., WikiHow, poetry, recipes), and reviews and summaries (e.g., Yelp, XSum), ensuring broad coverage of both formal and informal text styles.
+This experiment employed 19,117 instances for the training and 4,779 samples for the testing.
+We conducted the featured engineering noted in 4.3 and AI detections noted in 4.2 to extract features and get inferences from classifiers.
 4.2 AI-Generated Text Classifiers
-To detect AI-generated text, we employed a combination of openly accessible classifiers, RoBERTa-base (Solaiman et al., 2019), RoBERTa-large (Solaiman et al., 2019), and RADAR (Hu et al., 2023).
-These classifiers were selected based on their open-source availability and the ability to produce unified probabilistic outputs ranging from 0 to 1 with seven decimal points.
-GPT4o-mini was incorporated for its promising application in AI content detection tasks (Bhattacharjee & Liu, 2024).
-Unlike binary classification approaches, these models assign a probability score indicating the likelihood that a given text was generated by AI rather than directly making a discrimination.
-To maintain uniformity in our evaluation, GPT4o-mini was explicitly prompted to return probability scores within the same range as other detectors.
-Figure 3: Histograms show how AI-generated probability distributions differ by text length (Short in red, Medium in green, Long in blue).
-If we use the threshold of 0.5 (black dash line), human-written medium-length text shows higher error rate than other lengths.
-Each kernel density estimation (KDE) curve reflects the probability scores assigned by RoBERTa-large on the training dataset.
+To classify AI-generated text, we employed multiple AI text detectors, DeskLib [39], E5LoRA [40], FastDetectGPT [41], RADAR [13], RoBERTa-base and -large [16], and SimpleAI’s detector [42] for open-source availability.
+Among closed-source models, GPT4.1-nano and GPT4o-mini [43, 44] have been used for LLM-based text detection [7].
+These detectors were selected based on their common application and used in both the train and test datasets.
+The closed-source text detection is only conducted in the test dataset.
+Figure 7: FairOPT decreases the Balanced Error Rate (BER) discrepancy of AI detectors compared to other thresholding methods.
+FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$).
+The error bar shows the standard deviation from the mean.
 4.3 Subgroup Attributes
-We extracted subgroup attributes from each dataset to undercover the potential disparities present in AI classifiers.
-These attributes are shown in Table 1 and include both non-sensitive attributes like text length and sensitive attributes inferred from stylistic traits.
+We extracted attributes from each text for sub-grouping to capture heterogeneity in text characteristics.
+These attributes are shown in Table 1.
+We categorized text length into three classes based on the number of characters as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500).
+We also labeled text with stylistic personality based on five traits by selecting the one with the highest probability among the five using a BERT-based personality prediction model [45].
+Details are noted in Appendix F.
+These features were chosen because text length captures structural variation and stylistic personality reflects writing tendencies, both of which critically influence detector stability.
 Attribute Characteristics Labels
-Text Length Number of characters in a text Short, medium, long (n = 3)
-Stylistic personality Inferred Big Five personality traits. Openness, conscientiousness, extroversion, agreeableness, neuroticism (n = 5)
-Table 1: The subgroup attributes studied in our experiments.
-Figure 4: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoBERTa-large detector probabilities.
-The visualization is based on KDE.
-The biggest discrepancy is observed with a KS statistic of $0.2078$ ($p<0.01$), while the smallest is $0.1001$ ($p<0.01$).
-These discrepancies indicate varying levels of divergence between groups based on the characteristics of the given text.
-(a) RoBERTa-large
-(b) RoBERTa-base
-Figure 5: The FairOPT algorithm enables us to obtain a balanced tradeoff between performance and fairness, through a relaxed fairness criterion specified by a given fairness threshold.
-We improve the Pareto frontier of performance and fairness compared to static and ROCFPR baselines.
-Text length measures the number of characters in a text.
-OpenAI’s evaluations (OpenAI, 2023) revealed performance disparities between short (under 1,000 characters) and long text when using their AI classifier.
-Building on this and using exploratory data analysis, we refined the thresholds to 1,000 and 2,500 characters, categorizing texts as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500).
-The stylistic personality attribute includes five traits: extroversion, neuroticism, agreeableness, conscientiousness, and openness (Minej, 2023).
-We determine the personality trait by selecting the one with the highest probability among the five using a BERT-based personality prediction model (Minej, 2023), which is fine-tuned from BERT-base-uncased (Devlin et al., 2018).
-Details are noted in Appendix D.
+Text length Number of characters in text. Short, Medium, Long (k = 3)
+Stylistic traits Inferred Big Five personality traits. Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism (k = 5)
+Table 1: Subgroup attributes from the feature engineering.
+Figure 8: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness rule that is shown in Figure 4.
+Our method incurs minimal loss of ACC, achieving $\mu=0.6344\pm 0.1518$ compared to Static ($\mu=0.6376\pm 0.1975$) and ROCFPR ($\mu=0.6249\pm 0.2057$) methods.
 4.4 Discovery of Subgroup Bias in Text Detection
-Our first experimental analysis uncovers many discrepancies in AI-text detection performance with respect to different text lengths and writing styles.
-For example, as shown in Figure 3, medium-length human-written text is more likely to be flagged as AI-generated than short-length human-written text.
-Similarly, in Figure 2, long text with a neurotic style is also more likely to be detected as AI-generated.
-To obtain a holistic view of these biases, we used the Kolmogorov-Smirnov (KS) test to evaluate distributional discrepancies across various text groups under the assumption of a non-parametric distribution.
-As shown in Figure 4, there are notable differences between certain subgroup pairs.
-For instance, the comparison between texts characterized by neuroticism and medium-length versus those with neuroticism and short-length yielded a KS statistic of 0.2078 ($p<0.01$).
-Extroversion and mid-length and neuroticism and short length, showed a KS statistic of 0.1959 ($p<0.01$).
-On the other hand, relatively smaller discrepancies were observed between groups such as neuroticism and short length versus openness and short length, with a KS statistic of 0.1155 ($p<0.001$), and between neuroticism and short length versus openness and mid-length, with a KS statistic of 0.1001 ($p<0.01$).
-These observed discrepancies show that varying subgroup characteristics can lead to diverging outcomes when using a fixed threshold for classification.
-Addressing these discrepancies through tailored threshold adjustments or other interventions is crucial to ensuring equitable treatment across all subgroups.
-Detector Method SemEval Kaggle MAGE
-F1 ACC FairLen FairPer FPR F1 ACC FairLen FairPer FPR F1 ACC FairLen FairPer FPR
-RoBERTa-large Static 0.4382 0.2847 0.1444 0.1377 0.9892 0.4309 0.2823 0.0607 0.1161 0.9830 0.7147 0.6360 0.1817 0.1760 0.6464
-ROCFPR 0.0740 0.2990 0.2197 0.1792 0.4854 0.3469 0.3900 0.1490 0.5782 0.6364 0.5201 0.6475 0.3471 0.2777 0.1031
-FairOPT 0.4849 0.3223 0.1347 0.1075 0.9892 0.4542 0.2973 0.0291 0.0755 0.9835 0.7088 0.6080 0.1749 0.1209 0.6473
-RADAR Static 0.5592 0.6373 0.3826 0.3399 0.2266 0.7097 0.7373 0.3088 0.2876 0.3360 0.4364 0.3930 0.5979 0.4599 0.6896
-ROCFPR 0.0153 0.5293 0.1162 0.0475 0.0019 0.0351 0.6333 0.1199 0.1412 0.0005 0.3215 0.4935 0.4102 0.4811 0.2662
-FairOPT 0.6307 0.6620 0.3536 0.4939 0.2228 0.6445 0.6183 0.2269 0.1566 0.3732 0.4541 0.3820 0.5705 0.4140 0.6965
-RoBERTa-base Static 0.3345 0.2133 0.0976 0.0978 0.9703 0.3659 0.2457 0.0950 0.1294 0.9553 0.7020 0.6515 0.2023 0.2124 0.5265
-ROCFPR 0.0740 0.2253 0.2463 0.1428 0.7633 0.3374 0.2943 0.0876 0.1276 0.8171 0.5814 0.6580 0.4275 0.3666 0.1739
-FairOPT 0.3548 0.2253 0.1072 0.0866 0.9696 0.3706 0.2447 0.0876 0.1234 0.9553 0.7122 0.6485 0.1966 0.1517 0.5265
-GPT4o-mini Static 0.3737 0.5497 0.1806 0.3137 0.2114 0.2949 0.3083 0.1042 0.2494 0.7390 0.2017 0.4340 0.2921 0.1990 0.2878
-ROCFPR 0.0004 0.5247 0.0053 0.0130 0.0057 0.0000 0.6267 0.0005 0.0020 0.0005 0.0000 0.5080 0.0064 0.0039 0.0020
-FairOPT 0.4371 0.5553 0.1907 0.6844 0.2139 0.3792 0.3223 0.1279 0.3471 0.7459 0.2802 0.4170 0.2578 0.1697 0.3006
-Table 2: Comparison of three thresholding methods (Static, ROCFPR, FairOPT) across four probabilistic AI text classifiers (RoBERTa-large, RoBERTa-base, RADAR, GPT4o-mini) and three data sources (SemEval, Kaggle, MAGE).
-Metrics include F1, ACC, FairLen (BER discrepancy for Length), FairPer (BER discrepancy for Personality), and FPR.
-Compared to the Static method, FairOPT results in an average increase of 2.91% in F1 score and decrease of 1.77% in FairPer, while causing a decrease of 0.58% in ACC and an increase of 1.59% in FairLen., demonstrating its effectiveness in enhancing specific aspect of performance and fairness with minimal trade-offs in classifier performance.
-Although ROCFPR method showed the best result in decreasing the FPR of the model, it overly sacrificed performance.
+We assess detection performance across text-length subgroups under a static threshold $\theta=0.5$.
+As illustrated in Figure 3, short human-authored samples incur a substantially higher FPR than mid-length and long texts.
+This length-dependent disparity underscores the necessity of text length adaptive thresholding to mitigate the discrepancy.
+Next, we examine the combined effects of length and stylistic personality (Figure 5).
+By comparing the empirical cumulative distribution functions (CDFs) of AI-likelihood scores for short, conscientious versus long, neurotic texts, a two-sample Kolmogorov–Smirnov (KS) test reveals a statistically significant divergence in their score distributions ($p<0.05$), indicating that static thresholding disproportionately misclassifies certain stylistic and length combinations (see Appendix G.1.1 for full KS-test details).
+The observed disparity in the combination of text length and stylistic personality further highlights the need for adaptive thresholding to guarantee equitable detection performance across heterogeneous text cohorts.
 4.5 Mitigating Subgroup Bias using FairOPT
-Given these findings, we now describe the experimental setup using FairOPT for more robust AI-text detection across subgroups.
-We extracted the text length and personality features from every item in the dataset, which enabled the partition of 15 subgroups with different characteristics.
-For the training dataset, the AI text classifiers RoBERTa-base, RoBERTa-large, and RADAR were employed to generate predicted probabilities indicating the likelihood that a given text was AI-generated.
-These probabilities ranged from 0 to 1, with no direct classification decisions made at this stage.
-For the test dataset, additional predictions were generated using the GPT4o-mini API.
-We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup.
-We include all hyperparameters and training details in Appendix F.3.
-Threshold optimization was performed on the training dataset based on the results only from RoBERTa-large rather than averaging the probabilities of three detectors.
-The relaxed fairness criteria in 3.2 is employed to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness.
-The RoBERTa-large model was also used to implement the AUROC-based approach (Method based on ROC curve B ) with $FPR<0.1$ constraint for comparison, and the static threshold 2 does not need a model for training in generating a decision threshold.
-The optimized thresholds were then applied to the test dataset.
-Performance was evaluated using F1, ACC, BER, and BER calculated per subgroup.
-FairLen and FairPer metrics were utilized to quantify the maximum discrepancies in BER across groups defined by text length and personality traits.
-These metrics are defined as follows:
-FairLen $=\max_{a\in L}\text{BER}(a)-\min_{a\in L}\text{BER}(g)$ (9)
-FairPer $=\max_{b\in P}\text{BER}(b)-\min_{b\in P}\text{BER}(g)$ (10)
-where $L$ includes three groups based on text length, $P$ includes five groups based on big five personality traits, and BER denotes the balanced error rate for each group.
-A higher value of FairLen or FairPer indicates a greater disparity in classification performance across subgroups.
+Given these findings, we now describe the experimental setup for applying FairOPT for AI-text detection with reduced discrepancy across subgroups.
+We extracted the text length and personality features from every instance in the dataset, partitioning the corpus into 15 subgroups.
+For the dataset, multiple AI text classifiers were employed to generate predicted probabilities indicating the probability that a given text was AI-generated with $p\in[0,1]$.
+We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup and applied them as noted in Figure 2.
+We include all hyperparameters and training details in Appendix H.3.
+Threshold optimization was performed on the training dataset based on the results from RoBERTa-large [16] rather than averaging the probabilities of multiple detectors.
+The relaxed fairness criterion is integrated to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness [46].
+We also applied an AUROC-based thresholding method using the RoBERTa-large [16] to select the decision threshold that maximizes the TPR under the constraint $FPR<0.1$; see Figure 6 for the AUROC-based optimization procedure [26].
+The optimized thresholds were then applied to the test dataset and evaluated based on discrepancy, ACC, and F1 score.
+Metrics are noted in Appendix B.2.
+4.6 Tradeoff Analysis between Performance and Fairness
+Figure 4 shows the performance-fairness tradeoff by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines.
+The relaxed fairness criterion from Section 3.2 to less sacrificing ACC while still satisfying a demanded level of fairness.
+The ROCFPR method, shown in Figure 6, demonstrates a notable ability to minimize the FPR in training.
+However, this reduction in FPR comes at a significant cost to the TPR.
+This trade-off is particularly problematic as it diminishes the detector’s reliability in correctly identifying AI-generated text, although the overall ACC is not overly sacrificed as shown in Figure 8.
 4.7 Results with FairOPT for AI-Text Detection
-Table 2 presents a comprehensive comparison of three thresholding methodologies—Static, ROCFPR, and FairOPT—applied to four AI text detectors (RoBERTa-base, RoBERTa-large, RADAR, GPT4o-mini) across three distinct datasets (MAGE, Kaggle, SemEval).
-The evaluation metrics include F1, ACC, FairLen (maximum BER discrepancy for text length), FairPer (maximum BER discrepancy for personality traits), and FPR.
-Overall across all datasets and tasks tested, using FairOPT yields an average increase of 2.91% in F1 Score and a decrease of 1.77% in FairPer, thereby strictly improving the performance-fairness tradeoff.
-Overall accuracy was only reduced by a marginal 0.58%, indicating that FairOPT does not affect performance accuracy.
-These results indicate that FairOPT effectively enhances the overall classification performance and reduces fairness disparities across subgroups, outperforming the Static threshold approach by achieving higher F1 scores and lower BER discrepancies related to stylistic features.
-We now analyze the results of FairOPT on each specific dataset.
-On the SemEval task, overall F1 scores show significant improvement, with RADAR achieving a notable increase of 7% after running our FairOPT algorithm.
-These findings imply that environments exhibiting patterns similar to SemEval can greatly benefit from this method.
-Additionally, the RoBERTa-large model exhibits a substantial reduction in discrepancy across all datasets that it was tested on - in the MAGE dataset, the discrepancy related to writing style based on personality decreases by 5%.
-This implies that models that work similarly to RoBERTa-large can pursue better fairness through our method.
-For example, RoBERTa-based models also exhibit reduced disparities in text length and style.
-These results provide strong evidence to support the future application of FairOPT to other models, particularly detectors currently under development, as outlined in OpenAI (2023).
-4.7 Tradeoff Analysis between Performance and Fairness
-Figure 5 illustrates the accuracy-fairness tradeoffs achieved by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines.
-The relaxed fairness criterion from Section 3.2 effectively navigates the trade-off space, enhancing the Pareto frontier between classification performance and fairness metrics.
-Especially, the relaxed fairness criterion leads to early termination of the algorithm before overly sacrificing ACC while still satisfying a reasonable level of fairness.
-In contrast, the ROCFPR baseline, while achieving lower FPRs as noted in Table 2, significantly compromises TPR thus harming overall ACC, highlighting the superiority of FairOPT in maintaining a harmonious balance between constraints and model performance.
-The ROCFPR method demonstrates a notable ability to minimize the FPR across various classifiers and datasets.
-However, this reduction in FPR comes at a significant cost to the TPR, leading to a sacrifice of ACC as illustrated in Appendix B.
-This trade-off is particularly problematic as it diminishes the detector’s reliability in correctly identifying AI-generated text.
-Conversely, the adaptive thresholding approach employed by FairOPT offers advantages in F1 and group fairness over both static and ROCFPR methods by tailoring decision boundaries to subgroup-specific probability distributions.
-This strategy ensures that each subgroup, characterized by attributes such as text length and personality traits, is evaluated with an appropriate threshold, thereby balancing the trade-off between fairness and performance.
-FairOPT achieves this balance by improving F1 scores and reducing FairPer discrepancies without substantially sacrificing ACC, thereby enhancing the classifier’s overall performance and addressing fairness.
+Figure 7 demonstrates that group-adaptive thresholding from FairOPT consistently lowers the average discrepancy.
+Our method decreased the discrepancy to $0.4341$ ($\sigma=0.1927$), showing $12\%$ of reduction compared to the static-threshold baseline ($\mu=0.5512$, $\sigma=0.1758$) and outperforms the AUROC-tuned method ($\mu=0.5595$, $\sigma=0.1582$).
+Notably, E5LoRA shows the greatest decrease, with its BER discrepancy dropping by nearly half (from $0.63$ to $0.33$), and GPT4o-mini achieved a substantial reduction (from $0.60$ to $0.37$).
+The Desklib experiences a modest 0.003 reduction.
+Overall, the FairOPT decreased the discrepancy in all AI text classifiers.
+The detailed results with standard deviation by the dataset, detector, and thresholding methods are in Appendix LABEL:tab:overall_by_detector_and_dataset.
+Figure 8 illustrates that FairOPT incurs only a 0.003 of sacrifice in overall ACC and reduction of the standard deviation by 0.04 while outperforming the AUROC-based method.
+These findings demonstrate that FairOPT substantially mitigates fairness disparities with negligible ACC tradeoff, offering a practical application.
+Notably, our method shows great effectiveness to commonly used AI detectors like DeskLib [39], FastDetectGPT [41], RADAR [13], and SimpleAI’s detector [42] by securing ACC over $75\%$ while reducing the discrepancy in all cases.
+Moreover, the E5LoRA detector [40] exemplifies the ability of FairOPT to transcend the fairness–performance tradeoff.
+On the SemEval24 dataset, FairOPT achieves a 0.43 reduction in discrepancy while increasing ACC by 0.23.
+Similarly, on the MAGE, it delivers a 0.25 decrease in discrepancy coupled with a 0.21 ACC increase.
+These outcomes highlight the broad applicability of adaptive thresholding for improving both fairness and performance across diverse AI-text detection models, as well as other kinds of probabilistic classifiers.
 5 Conclusion
-This paper proposed FairOPT, based on the observation that there is a statistically significant difference in AI-generated text classifier outputs across subgroups of text (e.g.
-length and personality) and that this discrepancy can lead to notably higher error rates in specific groups.
-To address this problem, we proposed FairOPT, a new method for group-specific threshold optimization in AI-generated content classifiers without sacrificing performance.
-Unlike a universal threshold, which presumes well-calibrated probabilities across varied text distributions, our method effectively captures the variations of the probability distribution across different subgroups, leading to a more robust and fair classification.
-While our work focuses primarily on text classifiers, the FairOPT framework can naturally extend to other modalities where detecting AI-generated content is critical for safety and copyright (e.g., image, video, and audio).
-This method will be highly beneficial in post-processing, guiding the model toward enhanced performance or increased fairness, contingent upon the specific requirements and deployment context.
-Furthermore, while our primary focus has been on improving fairness in the post-training stage, we hypothesize that the proposed method can also be integrated as an in-training fairness enhancement mechanism.
+This paper proposes FairOPT, a post-processing algorithm for probabilistic classifiers that reduces the discrepancy while preserving overall accuracy.
+Based on the observation that there is a statistically significant difference in AI detector outputs by text characteristics (e.g., length and personality), the need to address the variance has been raised because it can lead to notably higher error rates for specific groups.
+To address this problem, we proposed a new method for group-specific threshold optimization.
+Unlike universal thresholding methods, which presume evenly distributed probability regardless of the stylistic features of the text, our method effectively captures the variations of the probability distribution of AI text classifiers across different subgroups, leading to classification with a lower discrepancy.
+While our work focuses primarily on text classifiers, the FairOPT framework can naturally extend to other modalities where detecting AI-generated content is critical for safety and copyright (e.g., image, video, and audio), including deepfake video and audio detectors.
+This method will be highly beneficial in post-processing, guiding the model toward decreased discrepancy, contingent upon the specific requirements and deployment context.
+We can also apply this method to all other probabilistic classifications when subgroup discrepancy matters (e.g., recidivism risk prediction or hiring decisions) by learning tailored decision thresholds that substantially mitigate bias with negligible loss in overall accuracy.
+Furthermore, while our primary focus has been improving fairness in the post-training stage, we hypothesize that the proposed method can also be integrated as an in-training mechanism.
 Finally, other approaches for threshold optimization, including Bayesian methods and convex optimization, can also be studied in future research.
+Although our experiments focus on text-based detectors, the FairOPT framework is modality-agnostic and readily extends to other domains such as image, video, and audio deepfake detection, where probabilistic outputs require post-processing to ensure equitable performance.
+More generally, FairOPT applies to any probabilistic classification task with subgroup disparities (e.g., recidivism risk prediction, hiring decisions) by learning tailored decision thresholds that mitigate bias with minimal impact on overall accuracy.
+6 Limitations
+While FairOPT demonstrably reduces subgroup discrepancy on our English-text benchmarks, it relies on only two features (character-count bins and model-inferred Big Five traits).
+Other text attributes (e.g., topic, syntax, language) may induce additional biases that our current feature set does not capture.
+Moreover, the stylistic labels are derived from a downstream BERT-based personality predictor, which itself may introduce cultural bias.
+We focus on relaxing demographic parity—extensions to other fairness definitions (equal opportunity or equalized odds) may require different optimization objectives or could be fundamentally incompatible.
 Impact Statement
 This paper introduces novel theoretical frameworks and methodologies to enhance the robustness of AI-generated text detection systems.
-By extracting and using features for subgroups such as text length and stylistic traits, our approach addresses disparity in classification outcomes.
-This technology can play a crucial role in AI detection for identifying misinformation (Chen & Shu, 2024), safeguarding the integrity of publication and academic organizations (Wu et al., 2023; Perkins, 2023), and countering potential cybersecurity threats (Yao et al., 2024).
+By extracting and using features for subgroups such as text length and stylistic traits, our approach addresses disparities in classification outcomes.
+This technology can play a crucial role in AI detection for identifying misinformation [3], safeguarding the integrity of publication and academic organizations [4, 6], and countering potential cybersecurity threats [5].
 Since it is crucial for these AI-content detection methods to be robust and fair across many potential users, our method takes a major step in this direction by formulating the problem setting and developing a new algorithm.
-Acknowledgment
-Thanks to Hengzhi Li, Megan Tjandrasuwita, David Dai, and Jeongmin Kwon, Chanakya Ekbote for constructive feedback and discussion.
-Appendix A Performance and fairness metrics
-A.1 Contingency table
+Appendix A Technical Appendices and Supplementary Material
+Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline.
+There is no page limit for the technical appendices.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3.
+Answer: [Yes] We introduced a novel method that reduces the discrepancy of the probabilistic classifiers, especially in AI text detectors, and the main claim is noted both in the abstract and introduction.
+  4. Justification: We noted the main claims in the abstract and the main body of the paper and provided experimental results as evidence to support the main claim.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes] Limitations are addressed at the end of the paper.
+  9. Justification: Limitations about the AI text detection methods and questions around the inference to other kinds of probabilistic classifiers are addressed.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes] The paper provides the conceptual framework and the computational process
+  14.
+Justification: We provides the confusion matrix, employed performance and fairness metrics, fairness criterion, multiple thresholding methods, and the optimization process based on the finite-difference approximation.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes] The paper includes the URL to the dataset and codes that are used for the main results.
+  19.
+Justification: The paper provides the code and dataset.
+The experiment can be simulated with the provided dataset, and other experiments can also be conducted if they are based on the probabilistic classifiers and split into subgroups by adjusting the column names or adjusting hyperparameters.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes] We disclose all codes and dataset.
+  24. Justification: You can access the data and code through the provided technical materials or via the URL.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes] We disclose the dataset balance, split, and characteristics of data. The employed hyperparameters are included in the appendix. The optimizer is developed in this research.
+  29. Justification: The details about the dataset construction is noted in the dataset part, and appendix includes the hyperparameters.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33.
+Answer: [Yes] We report the error bars, which implies the standard deviation of the results.
+Also, the statistical significance of the probability distribution is addressed using the K-S test in the main body of the paper.
+  34.
+Justification: The reported error bar shows the variance of the result and we conducted an experiment to figure out the p-value of the statistical significance of the KS probability distribution.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [No] This research is based on a local device (CPU) and does not use GPU
+  39. Justification: The model is designed to be able to run in any environment without the constraint of the hardware.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes] This research complies with the Code of Ethics
+  44.
+Justification: Since this research is not a human-subjects research, this project is exempt from the IRB investigation.
+This research does not include human-subject sensitive information, but they are augmented using feature engineering.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes] We have included potential impacts from this research in Section Impact Statement.
+  49.
+Justification: This research generate sensitive features.
+But the extracted features do not infringe privacy because it is generating labels for each writing based on statistical inferences.
+This could lead a negative impacts.
+Regarding the positive impacts, the decreased discrepancy of the classifiers has been discussed.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [Yes] We used the clipping method to restrict the optimization of the probability in the range of 0 to 1
+  54.
+Justification: Since the results outside of the range of 0 and 1 could lead to an error in the system, we used clipping in the model to prevent the exceed of the bound in the gradient descent process.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes] We used multiple AI text classifiers and they are all cited and addressed. Some codes may need the API keys from specific platforms like OpenAI
+  59. Justification: The train and test datasets include multiple AI-generated text detectors. And the test dataset has two more detection results based on the open-source LLMs.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes] All resources used for the paper are provided in the URL
+  64. Justification: We disclose all resources that used for the research.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A] This research is not crowdsourcing research.
+  69.
+Justification: The given multiple types of writings (e.g., essay, twitter, etc.) are constructed through the cited datasets, and the main novelty of the dataset is the application of multiple AI text detectors.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A] This research is not human-subject research
+  74. Justification: Some texts are written by human, but the sources are authorized and publicly accessible datasets.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes] This research used two language models for the AI text detection, which his not a conventional method despite there being research that shows its efficiency.
+  79. Justification: This research used GPT4.1 and GPT4 like other AI text classifiers. This method is explored in other research as discussed in the main part of the paper.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix B Performance and fairness metrics
+B.1 Contingency table
 The model can make not only true positive (TP) and true negative (TN) cases but also false positive (FP) and false negative (FN):
@@
 Table 2: Confusion matrix for the four possible outcomes when an AI detection system classifies either AI-generated or human-developed content.
-A.2 Notation and Metrics
-Given the table A.1 that denotes the outcomes in a binary classification task, we can mathematically express the performance and fairness metrics like below:
+B.2 Notation and Metrics
+Given the table B.1 that denotes the outcomes in a binary classification task, we can mathematically express the performance and fairness metrics like below:
   • Accuracy (ACC).
@@
     Recall measures the fraction of actual positives that are correctly identified by the model, capturing how many real positives are detected.
-  • Demographic Parity (DP).
-Suppose we have a protected attribute $A\in\{0,1\}$.
-Demographic Parity requires $P(\hat{Y}=1\mid A=0)=P(\hat{Y}=1\mid A=1)$, or $P(\hat{Y}=1\mid A=0)\approx P(\hat{Y}=1\mid A=1)$.
-A common metric is the disparity:
+  • Demographic Parity (DP). Suppose we have a protected attribute $A\in\{0,1\}$. Demographic Parity requires $P(\hat{Y}=1\mid A=0)=P(\hat{Y}=1\mid A=1)$, or
+    $P(\hat{Y}=1\mid A=0)\approx P(\hat{Y}=1\mid A=1)$. A common metric is the disparity:
     $\Delta_{\mathrm{DP}}\;=\;\bigl{|}\,P(\hat{Y}=1\mid A=0)\;-\;P(\hat{Y}=1\mid A=1)\bigr{|}.$
   • Equalized Odds (EO). Equalized Odds requires equal true positive rates and false positive rates across groups. One way to measure the gap is:
-    $\Delta_{\mathrm{EO}}\;=\;\bigl{|}P(\hat{Y}=1\mid Y=1,A=0)\;-\;P(\hat{Y}=1\mid Y=1,A=1)\bigr{|}\;+\;\bigl{|}P(\hat{Y}=1\mid Y=0,A=0)\;-\;P(\hat{Y}=1\mid Y=0,A=1)\bigr{|}.$
+    $\Delta_{\mathrm{EO}}\;=\;\bigl{|}P(\hat{Y}=1\mid Y=1,A=0)\;-\;P(\hat{Y}=1\mid Y=1,A=1)\bigr{|}\;+\;\ \$
+    $\bigl{|}P(\hat{Y}=1\mid Y=0,A=0)\;-\;P(\hat{Y}=1\mid Y=0,A=1)\bigr{|}.$
     Under the strictest notion of fairness, one requires
@@
   • Balanced Error Rate (BER).
-    $\mathrm{BER}\;=\;\frac{1}{2}\left(\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}+\frac{\mathrm{FN}}{\mathrm{TP}+\mathrm{FN}}\right)\;=\;\frac{\mathrm{FPR}+(\mathrm{FNR})}{2}\;=\;\frac{\mathrm{FPR}+(\mathrm{1-TPR})}{2}$
+    $\mathrm{BER}\;=\;\frac{1}{2}\left(\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}+\frac{\mathrm{FN}}{\mathrm{TP}+\mathrm{FN}}\right)\;=\;\frac{\mathrm{FPR}+\mathrm{FNR}}{2}\;=\;\frac{\mathrm{FPR}+(\mathrm{1-TPR})}{2}$
     Balanced Error Rate (BER) averages the error rates across both classes, providing a metric that is insensitive to class imbalance.
 By considering both the False Positive Rate and the False Negative Rate, BER ensures that the classifier performs equally well on both classes.
-The concept is especially well-introduced in (Ferrer, 2022)
+The concept is especially well-introduced in [32]
+  • BER Discrepancy.
+    $\text{Discrepancy}=\max\mathrm{BER}_{i}\;-\;\min\mathrm{BER}_{i}$
+    For each subgroup $G_{i}$ we can compute BERi to measure the discrepancy by comparing the maximum BER and minimum BER, and the lower discrepancy indicates more equitable classification across groups.
   • Specificity.
@@
     Specificity measures the proportion of actual negative cases correctly identified by the model, reflecting its ability to detect true negatives. It is also known as true negative rate (TNR).
-A.3 Impossibility theorem
+B.3 Impossibility theorem
 The impossibility theorem typically refers to the result that no single classifier can satisfy all common fairness metrics (e.g., demographic parity, equalized odds, predictive parity) simultaneously, except in trivial cases.
@@
 This research used one fairness metrics in the processing to address this issue.
-Appendix B AUROC
-B.1 Static threshold and AUROC-based approach
+Appendix C AUROC
+C.1 Static threshold and AUROC-based approach
 In this research, the AUROC-based approach is employed to optimize the classification threshold while balancing model performance and fairness constraints.
@@
 The optimal threshold (green marker) is selected to maximize performance under the specified FPR constraint.
-Appendix C Applied Thresholds
+Appendix D Applied Thresholds
 Figure 2 uses thresholds that is noted on Table 3.
@@
 Static 0.5
-ROCFPR 0.9984415
+ROCFPR 0.2866322
 Adaptive Thresholds (n=15)
 Group Threshold
-long_agreeableness 0.2545000
-long_conscientiousness 0.3055681
-long_extroversion 0.2500000
-long_neuroticism 0.2748645
-long_openness 0.2549773
-medium_agreeableness 0.2445749
-medium_conscientiousness 0.2448965
-medium_extroversion 0.1000000
-medium_neuroticism 0.2429920
-medium_openness 0.2436330
-short_agreeableness 0.2529650
-short_conscientiousness 0.2480741
-short_extroversion 0.2528146
-short_neuroticism 0.2528743
-short_openness 0.2489719
-Table 4: Thresholds that used in the study
-Appendix D Feature Engineering
+long_agreeableness 0.988102
+long_conscientiousness 0.983204
+long_extroversion 0.991482
+long_neuroticism 0.998293
+long_openness 0.978291
+medium_agreeableness 0.971192
+medium_conscientiousness 0.980804
+medium_extroversion 0.987482
+medium_neuroticism 0.994291
+medium_openness 0.969759
+short_agreeableness 0.986492
+short_conscientiousness 0.994242
+short_extroversion 0.978118
+short_neuroticism 0.994192
+short_openness 0.970664
+Table 3: Thresholds from three thresholding methods that used in the experiment.
+Appendix E Detailed Results
+We report a table showing the overall discrepancy, ACC, and F1 by dataset source: Table 4, and a table showing the detailed difference by detectors: Table LABEL:tab:overall_by_detector_and_dataset.
+E.1 Overall discrepancy and performance
+The table 4 summarizes the overall discrepancy and performance by each dataset and thresholding method. Detailed results that include the result from each detector is noted in the next section.
+Source Strategy Discrepancy (mean $\pm$ std) Accuracy (mean $\pm$ std) F1 Score (mean $\pm$ std)
+MAGE FairOPT $\mathbf{0.4495\pm 0.1656}$ $0.6339\pm 0.1512$ $0.3877\pm 0.3494$
+ROCFPR $0.5551\pm 0.1664$ $0.6238\pm 0.2113$ $0.5326\pm 0.3170$
+Static $0.5433\pm 0.1781$ $0.6388\pm 0.2015$ $0.5126\pm 0.3352$
+RAID FairOPT $\mathbf{0.4836\pm 0.1781}$ $0.6280\pm 0.1573$ $0.3885\pm 0.3535$
+ROCFPR $0.5644\pm 0.1263$ $0.6253\pm 0.2082$ $0.5363\pm 0.3185$
+Static $0.5562\pm 0.1478$ $0.6364\pm 0.1994$ $0.5136\pm 0.3358$
+SemEval FairOPT $\mathbf{0.3763\pm 0.2166}$ $0.6404\pm 0.1497$ $0.3896\pm 0.3484$
+ROCFPR $0.5598\pm 0.1758$ $0.6258\pm 0.2010$ $0.5303\pm 0.3114$
+Static $0.5548\pm 0.1969$ $0.6376\pm 0.1948$ $0.5081\pm 0.3288$
+Table 4: Overall discrepancy and performance by data source and thresholding strategy (mean $\pm$ std)
+E.2 Comparison by performance and fairness metrics by datasource, detector, and thresholding strategy
+The table LABEL:tab:overall_by_detector_and_dataset shows the results by each detector by the datasource and thresholding method based on fairness and performance metrics.
+Table 5: Discrepancy and performance (mean $\pm$ std) by data source, detector, and thresholding method
+Dataset
+Detector Thresholding $Discrepancy$ $ACC$ $F1score$
+MAGE
+GPT4.1-nano (2025) FairOPT $0.4106\pm 0.2350$ $0.4794\pm 0.0374$ $0.0114\pm 0.0121$
+ROCFPR $0.4636\pm 0.2194$ $0.4274\pm 0.0380$ $0.0694\pm 0.0159$
+Static $0.4647\pm 0.2143$ $0.4393\pm 0.0401$ $0.0507\pm 0.0154$
+GPT4o-mini (2024) FairOPT $0.3820\pm 0.2421$ $0.4866\pm 0.0345$ $0.0046\pm 0.0078$
+ROCFPR $0.6874\pm 0.1882$ $0.4453\pm 0.0288$ $0.2348\pm 0.0399$
+Static $0.6517\pm 0.1661$ $0.4423\pm 0.0253$ $0.1947\pm 0.0304$
+SimpleAI (2023) FairOPT $0.4536\pm 0.0871$ $0.7501\pm 0.0274$ $0.7006\pm 0.0363$
+ROCFPR $0.5682\pm 0.0945$ $0.7573\pm 0.0130$ $0.7520\pm 0.0215$
+Static $0.5527\pm 0.0968$ $0.7525\pm 0.0154$ $0.7360\pm 0.0298$
+DeskLib (2024) FairOPT $0.4896\pm 0.0914$ $0.8900\pm 0.0170$ $0.8809\pm 0.0196$
+ROCFPR $0.4854\pm 0.2388$ $0.8787\pm 0.0212$ $0.8916\pm 0.0224$
+Static $0.4278\pm 0.2561$ $0.9074\pm 0.0206$ $0.9140\pm 0.0219$
+E5LoRA (2024) FairOPT $0.3571\pm 0.2440$ $0.4895\pm 0.0336$ $00.0000\pm 0.0$ 0.0000
+ROCFPR $0.5779\pm 0.1866$ $0.1996\pm 0.0302$ $0.0471\pm 0.0179$
+Static $0.6026\pm 0.2439$ $0.2797\pm 0.0237$ $0.0065\pm 0.0090$
+FastDetectGPT (2023) FairOPT $0.4929\pm 0.0189$ $0.8099\pm 0.0274$ $0.7774\pm 0.0243$
+ROCFPR $0.4881\pm 0.0315$ $0.7776\pm 0.0302$ $0.8114\pm 0.0288$
+Static $0.4911\pm 0.1295$ $0.8398\pm 0.0154$ $0.8520\pm 0.0151$
+RoBERTa-base (2019) FairOPT $0.4893\pm 0.0537$ $0.6742\pm 0.0436$ $0.5381\pm 0.0655$
+ROCFPR $0.5923\pm 0.0884$ $0.7322\pm 0.0311$ $0.6963\pm 0.0320$
+Static $0.6158\pm 0.1606$ $0.7256\pm 0.0364$ $0.6722\pm 0.0423$
+RoBERTa-large (2019) FairOPT $0.3940\pm 0.1992$ $0.5063\pm 0.0364$ $0.0656\pm 0.0372$
+ROCFPR $0.5835\pm 0.0610$ $0.6264\pm 0.0289$ $0.5067\pm 0.0405$
+Static $0.5414\pm 0.1038$ $0.6103\pm 0.0361$ $0.4315\pm 0.0522$
+RADAR (2023) FairOPT $0.5760\pm 0.0755$ $0.6192\pm 0.0216$ $0.5106\pm 0.0388$
+ROCFPR $0.5495\pm 0.2113$ $0.7693\pm 0.0186$ $0.7837\pm 0.0202$
+Static $0.5418\pm 0.1280$ $0.7519\pm 0.0224$ $0.7562\pm 0.0242$
+RAID
+GPT4.1-nano (2025) FairOPT $0.5208\pm 0.2082$ $0.4738\pm 0.0087$ $0.0128\pm 0.0115$
+ROCFPR $0.5820\pm 0.1623$ $0.4243\pm 0.0189$ $0.0795\pm 0.0295$
+Static $0.5781\pm 0.1558$ $0.4306\pm 0.0183$ $0.0465\pm 0.0202$
+GPT4o-mini (2024) FairOPT $0.4781\pm 0.2428$ $0.4710\pm 0.0180$ $00.0000\pm 0.0$ 0.0000
+ROCFPR $0.5762\pm 0.1567$ $0.4557\pm 0.0266$ $0.2292\pm 0.0542$
+Static $0.5816\pm 0.1733$ $0.4529\pm 0.0180$ $0.1875\pm 0.0321$
+SimpleAI (2023) FairOPT $0.5868\pm 0.2080$ $0.7320\pm 0.0408$ $0.6835\pm 0.0548$
+ROCFPR $0.5655\pm 0.1818$ $0.7452\pm 0.0386$ $0.7502\pm 0.0388$
+Static $0.6245\pm 0.2095$ $0.7403\pm 0.0516$ $0.7357\pm 0.0541$
+DeskLib (2024) FairOPT $0.3552\pm 0.1001$ $0.9023\pm 0.0289$ $0.8973\pm 0.0301$
+ROCFPR $0.4578\pm 0.0522$ $0.8667\pm 0.0223$ $0.8843\pm 0.0194$
+Static $0.4292\pm 0.0646$ $0.8988\pm 0.0212$ $0.9076\pm 0.0189$
+E5LoRA (2024) FairOPT $0.4167\pm 0.2041$ $0.4808\pm 0.0153$ $00.0000\pm 0.0$ 0.0000
+ROCFPR $0.6437\pm 0.1177$ $0.2059\pm 0.0384$ $0.0406\pm 0.0272$
+Static $0.6535\pm 0.1981$ $0.2805\pm 0.0323$ $0.0078\pm 0.0097$
+FastDetectGPT (2023) FairOPT $0.4405\pm 0.0560$ $0.8241\pm 0.0203$ $0.7996\pm 0.0230$
+ROCFPR $0.5103\pm 0.0873$ $0.7850\pm 0.0345$ $0.8191\pm 0.0312$
+Static $0.4609\pm 0.1409$ $0.8472\pm 0.0247$ $0.8601\pm 0.0228$
+RoBERTa-base (2019) FairOPT $0.5361\pm 0.0452$ $0.6706\pm 0.0343$ $0.5483\pm 0.0681$
+ROCFPR $0.6246\pm 0.1244$ $0.7348\pm 0.0355$ $0.7030\pm 0.0478$
+Static $0.5978\pm 0.0984$ $0.7180\pm 0.0342$ $0.6682\pm 0.0499$
+RoBERTa-large (2019) FairOPT $0.4282\pm 0.1927$ $0.4990\pm 0.0183$ $0.0698\pm 0.0285$
+ROCFPR $0.6079\pm 0.0816$ $0.6413\pm 0.0218$ $0.5341\pm 0.0394$
+Static $0.5570\pm 0.0658$ $0.6176\pm 0.0175$ $0.4579\pm 0.0319$
+RADAR (2023) FairOPT $0.5900\pm 0.1850$ $0.5980\pm 0.0179$ $0.4853\pm 0.0412$
+ROCFPR $0.5113\pm 0.0602$ $0.7683\pm 0.0263$ $0.7867\pm 0.0233$
+Static $0.5232\pm 0.0644$ $0.7418\pm 0.0265$ $0.7510\pm 0.0257$
+SemEval
+GPT4.1-nano (2025) FairOPT $0.2663\pm 0.2792$ $0.4889\pm 0.0498$ $0.0110\pm 0.0116$
+ROCFPR $0.5139\pm 0.2100$ $0.4459\pm 0.0354$ $0.0912\pm 0.0320$
+Static $0.4889\pm 0.2156$ $0.4495\pm 0.0408$ $0.0503\pm 0.0210$
+GPT4o-mini (2024) FairOPT $0.2763\pm 0.2923$ $0.4943\pm 0.0501$ $0.0029\pm 0.0076$
+ROCFPR $0.5777\pm 0.2485$ $0.4638\pm 0.0350$ $0.2381\pm 0.0484$
+Static $0.5617\pm 0.2843$ $0.4620\pm 0.0418$ $0.2008\pm 0.0364$
+SimpleAI (2023) FairOPT $0.5655\pm 0.2123$ $0.7382\pm 0.0369$ $0.6842\pm 0.0260$
+ROCFPR $0.6726\pm 0.1674$ $0.7430\pm 0.0313$ $0.7375\pm 0.0264$
+Static $0.6519\pm 0.1904$ $0.7436\pm 0.0327$ $0.7283\pm 0.0231$
+DeskLib (2024) FairOPT $0.3711\pm 0.0755$ $0.8942\pm 0.0349$ $0.8844\pm 0.0316$
+ROCFPR $0.4487\pm 0.0727$ $0.8643\pm 0.0169$ $0.8776\pm 0.0241$
+Static $0.3691\pm 0.0970$ $0.8954\pm 0.0139$ $0.9012\pm 0.0184$
+E5LoRA (2024) FairOPT $0.2143\pm 0.2673$ $0.5015\pm 0.0523$ $00.0000\pm 0.0$ 0.0000
+ROCFPR $0.6887\pm 0.2807$ $0.2098\pm 0.0405$ $0.0293\pm 0.0124$
+Static $0.6428\pm 0.2790$ $0.2750\pm 0.0402$ $0.0065\pm 0.0061$
+FastDetectGPT (2023) FairOPT $0.4818\pm 0.0312$ $0.8111\pm 0.0254$ $0.7701\pm 0.0298$
+ROCFPR $0.4689\pm 0.0646$ $0.7490\pm 0.0284$ $0.7851\pm 0.0345$
+Static $0.5232\pm 0.1360$ $0.8159\pm 0.0253$ $0.8268\pm 0.0276$
+RoBERTa-base (2019) FairOPT $0.4922\pm 0.0426$ $0.6940\pm 0.0371$ $0.5676\pm 0.0472$
+ROCFPR $0.5748\pm 0.1175$ $0.7519\pm 0.0241$ $0.7127\pm 0.0362$
+Static $0.5543\pm 0.0866$ $0.7412\pm 0.0263$ $0.6847\pm 0.0435$
+RoBERTa-large (2019) FairOPT $0.2956\pm 0.2093$ $0.5188\pm 0.0513$ $0.0718\pm 0.0086$
+ROCFPR $0.5977\pm 0.1239$ $0.6479\pm 0.0218$ $0.5280\pm 0.0376$
+Static $0.6073\pm 0.1346$ $0.6186\pm 0.0293$ $0.4333\pm 0.0302$
+RADAR (2023) FairOPT $0.4241\pm 0.1244$ $0.6222\pm 0.0358$ $0.5140\pm 0.0407$
+ROCFPR $0.4956\pm 0.0187$ $0.7561\pm 0.0310$ $0.7735\pm 0.0357$
+Static $0.5939\pm 0.1847$ $0.7370\pm 0.0300$ $0.7407\pm 0.0361$
+E.3 Results for Discrepancy, ACC, and F1
+The bar charts illustrate the discrepancy, ACC, and F1 are incorporated in the main body of the paper, and this section encompasses a chart representing the F1 score, accompanied by three figures.
+The sequence of figures is arranged in the order of discrepancy, ACC, and F1.
+Figure 10: FairOPT decreases the discrepancy, Balanced Error Rate (BER) discrepancy, of probabilistic classifiers compared to other thresholding methods.
+FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$).
+Figure 11: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness that is shown in Figure 4.
+Our method incurs only a negligible accuracy sacrifice, achieving $\mu=0.6344\pm 0.1518$ compared to Static $\mu=0.6376\pm 0.1975$ and ROCFPR $\mu=0.6249\pm 0.2057$.
+Figure 12: Overall F1 have been sacrificed in the application of FairOPT.
+FairOPT achieves $\mu=0.389\pm 0.350$, compared to ROCFPR $\mu=0.533\pm 0.316$ and Static $\mu=0.511\pm 0.333$.
+Detectors with low ACC show a very low F1 score.
+Appendix F Feature Engineering
 In order to mitigate disparities within AI classifiers, a thorough feature engineering process was undertaken.
 Features were systematically extracted to characterize each text within the dataset.
-The table 5 includes all extraced features for this research.
 The primary focus of the experiment encompassed text length and anticipated personality traits.
@@
 The text length attribute measures the number of characters in a text.
-OpenAI’s evaluations (OpenAI, 2023) revealed performance disparities between short (under 1,000 characters) and long texts when using their AI classifier.
+OpenAI’s evaluations [26] revealed performance disparities between short (under 1,000 characters) and long texts when using their AI classifier.
 While their global threshold approach failed to address group-specific differences, it provided a valuable benchmark.
@@
 This attribute contributes to the enforcement of fairness metrics, such as demographic parity and equality of odds.
-For this feature, we determine the personality trait by selecting the one with the highest probability among the five using the (Minej, 2023) model, which is fine-tuned from BERT-base-uncased (Devlin et al., 2018).
-Attribute Characteristics Labels
-Text Length Measures the number of characters in a text short, medium, long (n = 3)
-Formality Indicates the level of formality in the language formal, informal (n = 2)
-Sentiment Reflects the emotional tone. positive, negative (n= 2)
-Personality Represents the inferred Big Five personality traits. Openness, Conscientiousness, Extroversion, Agreeableness, Neuroticism (n = 5)
-Table 5: Characteristics and labels of each attribute
-Outside of the length and sentiment, the formality attribute measures a text’s formality by analyzing the frequency of specific parts of speech.
-According to (Heylighen & Dewaele, 2002), with formal texts typically containing a higher proportion of nouns, adjectives, articles, and prepositions, whereas informal texts tend to have a greater frequency of pronouns, adverbs, verbs, and interjections.
-As general rule, we set a threshold of 50 to distinguish between formal (scores above 50) and informal texts.
-The sentiment attribute allows for the assessment of emotional tone and attitudes within content.
-AI-generated texts may display distinct sentiment patterns due to the differences in their underlying algorithms compared to human expression.
-To classify texts as positive or negative, we used the label output from the (DistilBERT-community, ) model, which is a fine-tuned version of DistilBERT-base-uncased (Sanh et al., 2019), specifically trained on the SST-2 dataset(Socher et al., 2013).
-Appendix E Discrepancies by classifiers with additional features
-E.1 Discrepancy of Probability in Subgroups
-E.1.1 Discrepancy in RoBERTa-large Detector
+For this feature, we determine the personality trait by selecting the one with the highest probability among the five using the [45] model, which is fine-tuned from BERT-base-uncased [47].
+Appendix G Discrepancies by classifiers with additional features
+G.1 Discrepancy of Probability in Subgroups
+G.1.1 Discrepancy in RoBERTa-large Detector
 The analysis revealed significant variability in the probability distributions between subgroup pairs, as depicted in Figure 13.
@@
 Significant differences were identified in certain subgroup comparisons, indicating notable discrepancies in their probability distributions.
-For example, comparing texts characterized by conscientiousness, positive sentiment, formal style, and long length with those exhibiting extroversion, negative sentiment, formal style, and long length resulted in a KS statistic of 0.7286 ($p<0.05$).
-Similarly, evaluating agreeableness, positive sentiment, formal style, and mid-length against conscientiousness, positive sentiment, formal style, and long length yielded a KS statistic of 0.7143 ($p<0.01$).
-Conversely, smaller discrepancies were observed between neuroticism, positive sentiment, formal style, and short length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2706, $p<0.05$) and neuroticism, negative sentiment, formal style, and short length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2323, $p<0.05$).
-Figure 7: Kernel Density Estimation (KDE) plots comparing probability distributions for subgroup pairs with the top two highest and lowest discrepancies based on the K-S test with statistical validity.
-The largest discrepancy is observed with a KS statistic of $0.7286$, while the smallest discrepancy has a KS statistic of $0.2323$.
-E.1.2 Discrepancy in RoBERTa-base Detector
-The analysis revealed a substantial variation in the probability distributions among pairs of subgroups, as shown in Figure 8.
-The KS test was applied to assess distributional differences across diverse text groupings defined by multiple feature sets, under the assumption of a non-parametric distribution.
-Specific subgroup comparisons highlighted significant differences, pointing to noticeable disparities in their probability distributions.
-For example, contrasting texts defined by agreeableness, positive sentiment, formal style, and mid-length with those featuring neuroticism, negative sentiment, informal style, and short length produced a KS statistic of 0.9 ($p<0.01$).
-Furthermore, the comparison between conscientiousness, negative sentiment, formal style, and long length and neuroticism, negative sentiment, informal style, and short length resulted in a KS statistic of 0.7750 ($p<0.01$).
-Figure 8: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity.
-The largest discrepancy is observed with a KS statistic of $0.9$, while the smallest discrepancy has a KS statistic of $0.5071$.
-These discrepancies indicate varying levels of divergence between groups based on their characteristics.
-Lesser discrepancies were noted between neuroticism, negative sentiment, informal style, and short length versus openness, positive sentiment, formal style, and long length (KS = 0.5522, $p<0.05$) and neuroticism, negative sentiment, informal style, and short length versus openness, positive sentiment, formal style, and short length (KS = 0.5071, $p<0.05$).
-E.1.3 Discrepancy in RADAR Detector
-The investigation disclosed considerable variability in the probability distributions across subgroup pairs, as illustrated in Figure 9.
-Through the application of the KS test, we evaluated distributional discrepancies among various text groups defined by distinct feature combinations, operating under the premise of a non-parametric distribution.
-Significant differences were discerned in particular subgroup comparisons, which highlight notable discrepancies in their probability distributions.
-For instance, the comparison between texts characterized by conscientiousness, negative sentiment, informal style, and long length and those exhibiting extroversion, positive sentiment, formal style, and short length yielded a KS statistic of 0.8 ($p<0.05$).
-Figure 9: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity.
-The largest discrepancy is observed with a KS statistic of $0.8$, while the smallest discrepancy has a KS statistic of $0.1830$.
-Similarly, the examination of conscientiousness, negative sentiment, informal style, and long length against neuroticism, negative sentiment, formal style, and short length resulted in a KS statistic of 0.6667 ($p<0.05$).
-In contrast, smaller discrepancies were detected between conscientiousness, negative sentiment, formal style, and long length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2110, $p<0.01$) and openness, negative sentiment, formal style, and long length versus openness, positive sentiment, formal style, and mid-length (KS = 0.1830, $p<0.01$).
-E.1.4 Discrepancy in GPT4o-mini Detector
-The analysis revealed substantial variability in the probability distributions among subgroup pairs, as depicted in Figure 10.
-Using the KS test, we assessed distributional discrepancies across various text groups defined by distinct feature combinations, assuming a non-parametric distribution.
-Significant differences were detected in particular subgroup comparisons, suggesting notable divergences in their probability distributions.
-For example, the comparison between texts characterized by openness, negative sentiment, informal style, and mid-length and those exhibiting openness, positive sentiment, informal style, and mid-length resulted in a KS statistic of 0.6667 ($p<0.001$).
-Similarly, the evaluation of extroversion, negative sentiment, formal style, and mid-length versus openness, positive sentiment, informal style, and mid-length produced a KS statistic of 0.6522 ($p<0.001$).
-Conversely, smaller discrepancies were noted between neuroticism, positive sentiment, formal style, and short length versus openness, negative sentiment, formal style, and long length (KS = 0.2853, $p<0.05$) and neuroticism, negative sentiment, formal style, and short length versus openness, negative sentiment, formal style, and long length (KS = 0.2560, $p<0.05$).
-Figure 10: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity.
-The largest discrepancy is observed with a KS statistic of $0.6667$, while the smallest discrepancy has a KS statistic of $0.2560$.
-Appendix F Operationalizing FairOPT
-F.1 Algorithmic description
+For example, comparing texts characterized by conscientiousness and short with those exhibiting neuroticism and long resulted in a KS statistic of 0.3081 ($p<0.01$).
+Similarly, evaluating conscientiousness and short against extroversion and medium yielded a KS statistic of 0.2616 ($p<0.01$).
+Conversely, smaller discrepancies were observed between agreeableness and short versus neuroticism and short (KS = 0.0909, $p<0.05$) and neuroticism and short versus openness and short (KS = 0.0842, $p<0.01$).
+Figure 13: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoB33ERTa-large detector results in cumulative probability distribution function (CDF) based on the training dataset in the range of 0.1 to 0.9.
+The biggest discrepancy is observed with a KS statistic of $0.3081$ ($p<0.01$), while the smallest is $0.0842$ ($p<0.01$).
+Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions
+Appendix H Operationalizing FairOPT
+H.1 Algorithmic description
 1:  Input: Labeled dataset $D=\{(x_{n},y_{n})\}$, predicted probabilities $\{p_{n}\}$, subgroup labels $\{g_{n}\}$ with possible groups $G_{1},\dots,G_{j}$, features $\{S_{1},\dots,S_{n}\}$, initial thresholds $\theta_{init}$, fairness metrics $\{M_{1},\ldots,M_{k}\}$ with acceptable disparity $\delta_{\mathrm{fair}}$, minimum performance thresholds: $\alpha$ (ACC) and $\beta$ (F1), penalty weight $\kappa$, learning rate $\eta$, tolerance $\epsilon_{\mathrm{tol}}$, maximum iterations, finite-difference step $\delta$, minimum probability for threshold $a$, maximum probability for threshold $b$
@@
 6:     for each group $G_{j}$ do
-7:       $\mathcal{I}_{i}\leftarrow\{\,n\mid g_{n}=G_{j}\}$// all samples in group $G_{j}$
+7:        $\mathcal{I}_{i}\leftarrow\{\,n\mid g_{n}=G_{j}\}$  all samples in group $G_{j}$
 8:        $\hat{y}_{n}\leftarrow[\,p_{n}\geq\theta(G_{j})\,],\;\forall n\in\mathcal{I}_{i}$
 9:        Store contingency table for $G_{i}$ using $\hat{y}_{n}$ and ${y}_{n}$
-10:       $\Omega\;\leftarrow\;\min_{i}\{\mathrm{ACC}(G_{j})\}\;\geq\;\alpha\;\wedge\;\min_{i}\{\mathrm{F1}(G_{j})\}\;\geq\;\beta$ // boolean output
+10:        $\Omega\;\leftarrow\;\min_{i}\{\mathrm{ACC}(G_{j})\}\;\geq\;\alpha\;\wedge\;\min_{i}\{\mathrm{F1}(G_{j})\}\;\geq\;\beta$ boolean output
 11:        for each fairness metric $M_{k}$ do
@@
 13:           $\Delta_{k}\leftarrow\max_{i}(M_{k}(G_{j}))-\min_{i}(M_{k}(G_{j}))$
-14:         $\Psi\;\leftarrow\;\Bigl{[}\,\max_{k}\Delta_{k}\;\leq\;\delta_{\mathrm{fair}}\Bigr{]}$ // boolean output
-15:       for each group $G_{j}$ do
-16:         $L_{i}(\theta)=-\mathrm{Acc}_{j}(\theta)+\kappa\,[\,\beta-\mathrm{F1}_{j}(\theta)\,]_{+}$
-17:
+14:           $\Psi\;\leftarrow\;\Bigl{[}\,\max_{k}\Delta_{k}\;\leq\;\delta_{\mathrm{fair}}\Bigr{]}$ boolean output
+15:        end for
+16:        for each group $G_{j}$ do
+17:           $L_{i}(\theta)=-\mathrm{Acc}_{j}(\theta)+\kappa\,[\,\beta-\mathrm{F1}_{j}(\theta)\,]_{+}$
+18:
 $\nabla L_{i}(\theta(G_{j}))\;\approx\;\frac{L_{i}(\theta(G_{j})+\delta)\;-\;L_{i}(\theta(G_{j})-\delta)}{2\,\delta}.$
-18:
+19:
 $\theta(G_{j})\;\leftarrow\;\theta(G_{j})\;-\;\eta\;\nabla L_{i}(\theta(G_{j})).$
-19:         Clip $\theta(G_{j})$ to stay within $[a,b]$
-20:    $\Delta_{\theta}\;\leftarrow\;\max_{G_{j}}\bigl{|}\theta(G_{j})-\text{previousThresholds}[G_{j}]\bigr{|}$
-21:    if ($\Delta_{\theta}<\epsilon_{\mathrm{tol}}\;\wedge\;\Omega\;\wedge\;\Psi$then
-22:       break
+20:           Clip $\theta(G_{j})$ to stay within $[a,b]$
+21:        end for
+22:     end for
+23:     $\Delta_{\theta}\;\leftarrow\;\max_{G_{j}}\bigl{|}\theta(G_{j})-\text{previousThresholds}[G_{j}]\bigr{|}$
+24:     if ($\Delta_{\theta}<\epsilon_{\mathrm{tol}}\;\wedge\;\Omega\;\wedge\;\Psi$then
+25:        break
+26:     end if
 […]
-24:    $\text{iteration}\leftarrow\text{iteration}+1$
-25:  return $\{\theta(G_{1}),\dots,\theta(G_{j})\}$
+28:     $\text{iteration}\leftarrow\text{iteration}+1$
+29:  end while
+30:  return $\{\theta(G_{1}),\dots,\theta(G_{j})\}$
 Algorithm 2 FairOPT: Gradient-based adaptive threshold optimization with fairness for subgroups
-F.2 Early stopping method for the tradeoff curve
-An optimization algorithm was implemented using a function from the scipy.optimize library, aiming to determine probability thresholds that satisfy specific fairness constraints (acceptable_disparity = [1.00, 0.30, 0.25, 0.21.
-0.20, 0.19, 0.15, 0.10]) while ensuring minimum performance thresholds for both performance criteria (ACC and F1).
+H.2 Early stopping method for the tradeoff curve
+The FairOPT algorithm was implemented using functions from the scipy.optimize library, aiming to determine probability thresholds that satisfy specific fairness constraints (acceptable_disparity = [1.00, 0.30, 0.25, 0.21, 0.20, 0.19, 0.15, 0.15]) while ensuring minimum performance thresholds for both performance criteria (ACC and F1).
 The optimization in the minimize function was performed using the L-BFGS-B method.
+The figure is based on the RoBERTa-base model, and it uses different hyperparameters to get eight results with different relaxation methods.
 Early stopping is facilitated by the default values of ftol and gtol in L-BFGS-B.
@@
 As stricter fairness constraints are imposed, these minimum thresholds are gradually relaxed.
-F.3 Hyperparameter settings for the FairOPT
+H.3 Hyperparameter settings for the FairOPT
 The application of the FairOPT is configured with a set of hyperparameters designed to ensure a balance between fairness and performance during the threshold optimization process.
@@
 Minimum thresholds for accuracy (0.25) and F1-score (0.25) are enforced to maintain baseline model performance and stability, particularly in cases of imbalanced datasets.
-A tolerance level of $10^{-2}$ is introduced to impose convergence criteria, ensuring the optimization process halts only when the changes in the objective function are negligible.
+A tolerance level of $1/2*10^{-4}$ is introduced to impose convergence criteria, ensuring the optimization process halts only when the changes in the objective function are negligible.
 Furthermore, a penalty parameter of 20 is applied to enforce performance constraints by penalizing significant performance disparities across subgroups.

### v5

@@
-Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection
+Minseok Jung1, Cynthia Fuertes Panizo2, Liam Dugan3, Yi R. (May) Fung1, Pin-Yu Chen4, Paul Pu Liang1
 Abstract
@@
 We partitioned data into subgroups based on attributes (e.g., text length and writing style) and implemented FairOPT to learn decision thresholds for each group to reduce discrepancy.
-In experiments with nine AI text classifiers on three datasets, FairOPT decreases overall balanced error rate (BER) discrepancy by 12% while minimally sacrificing accuracy by 0.003%.
+FairOPT showed notable discrepancy mitigation across nine detectors and three heterogeneous datasets, and the remarkable mitigation of the minimax problem by decreasing overall discrepancy 27.4% across five metrics while minimally sacrificing accuracy by 0.005%.
 Our framework paves the way for more robust classification in AI-generated content detection via post-processing.
 We release our data, code, and project information at URL.
-1 Introduction
+I Introduction
 The copious use of large language models (LLMs) facilitated widespread dissemination of AI-generated text [47, 20] and has raised concerns around the dissemination of misinformation [8], the degradation of publication standards [48], cybersecurity threats [49], and breaches of academic integrity [39].
@@
 We propose FairOPT to learn multiple decision thresholds tailored to each subgroup, which reduces discrepancy in AI text detection as shown in Figure 2.
-Figure 2: Applying adaptive thresholds to different probability distributions by subgroup in the test dataset, as seen through cumulative density functions (CDFs).
-We partition texts by three length categories and five personality traits as noted in Table 1.
-Observe that a static classification threshold ($0.5$, in black) and a single optimized threshold using AUROC-based method (in gray) do not account for subgroup-specific distributional variations.
-We used fifteen thresholds that have been generated using FairOPT to address the varying probabilistic distribution from AI text classifiers.
-The training is based on the RoBERTa-large detector.
-Details about the thresholds are in Appendix E.
 However, a single universal threshold for AI content classification neglects the heterogeneity of probability distributions by textual characteristics, such as length and writing style.
@@
 Analogous disparities have been documented in other classification tasks, regarding race [1, 4] and gender [46, 25, 4].
-To overcome these limitations, we introduce FairOPT, a group-adaptive thresholding framework that learns distinct decision thresholds for subgroups rather than using methods with one universal threshold (see Figure 1).
-Our method balances performance (e.g., accuracy (ACC), F1, precision) and fairness metrics (e.g., demographic parity, or equality of odds) in learning thresholds.
-Experiment on nine AI text detectors on three benchmark datasets demonstrates that FairOPT substantially reduces BER disparity across subgroups while incurring only a marginal tradeoff in overall performance.
-To summarize, our main contributions are: (a) identifying significant discrepancies in AI-generated text classifiers based on text length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold that calibrates the decision boundaries to reduce group discrepancy, and (c) showing the effectiveness of the FairOPT, achieving a 12% reduction in discrepancy while showing negligible tradeoff in performance by dropping less than 0.1% of ACC.
-2 Related Work
-AI text detectors:
+To overcome these limitations, we introduce FairOPT, a group‐adaptive thresholding framework that learns distinct decision thresholds for subgroups rather than using methods with one universal threshold (see Figure 1).
+Our method balances performance and fairness metrics (e.g., demographic parity, or equality of odds) in learning thresholds.
+Experiment on nine AI text detectors on three benchmark datasets demonstrates that FairOPT substantially reduces discrepancy across subgroups while incurring only a marginal tradeoff in overall performance.
+To summarize, our main contributions are: (a) identifying significant discrepancies in AI-generated text classifiers based on text length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold that calibrates the decision boundaries to reduce group discrepancy, and (c) showing the effectiveness of the FairOPT, achieving overall 27.4% reduction across five metrics in discrepancy while showing negligible tradeoff in performance by dropping less than 0.1% of ACC.
+II Literature review
+AI text detectors
 AI text detectors have been developed to tackle issues related to misinformation, integrity, and security.
@@
 the softmax scores across all candidate tokens), and records the rank position of the actual chosen token within that sorted probability to discern [18], DetectGPT examines the curvature of log probability distributions of the referred LLM, concentrating on local likelihood curvature to detect AI-generated text [35], and RoBERTa-based detectors fine-tune pretrained language models [42].
-We employed probabilistic classifiers for an AI-likelihood score that use a threshold for classification rather than architectures that generate a direct discrimination.
-Static thresholding for classification:
+We employed probabilistic classifiers for an AI‐likelihood score that use a threshold for classification rather than architectures that generate a direct discrimination.
+Static thresholding for classification
 The canonical thresholding in probabilistic classification is using a threshold of $0.5$ to map the probabilities to labels [17].
 This default rule is commonly used in many fields like photogrammetry [41], ecology [32, 21], and machine learning [31].
-Nevertheless, as highlighted by Freeman and Moisen [8], the dependence solely on $\theta=0.5$ as a default threshold, merely due to its conventional acceptance, could be unreliable because of variations in data.
-Universal thresholding with optimization:
+Nevertheless, as highlighted by [17], the dependence solely on $\theta=0.5$ as a default threshold, merely due to its conventional acceptance, could be unreliable because of variations in data.
+Universal thresholding with optimization
 Threshold optimization techniques have been formulated to learn an optimal threshold across the dataset.
@@
 While the AUROC-based method yields a globally optimized threshold tailored to the chosen model, it can lead to over-fitting or fail to adapt to other models.
-Adaptive thresholding across groups:
-Several studies advocate for group-adaptive decision thresholds to mitigate disparities [12, 27].
-Menon and Williamson [28] propose instance-dependent thresholds for inferred probabilities, resulting in reduced discrepancy.
-Similarly, Corbett-Davies et al.
-[29] demonstrate that group-specific thresholds can facilitate fairness-aware classification.
-Canetti et al.
+Adaptive thresholding across groups
+Several studies advocate for group‐adaptive decision thresholds to mitigate disparities [25, 2].
+[33] propose instance-dependent thresholds for inferred probabilities, resulting in reduced discrepancy.
+Similarly, [9] demonstrate that group-specific thresholds can facilitate fairness-aware classification.
 [7] investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity.
-Jang et al.
-[12] present threshold-adaptation methods to ensure fair classification, and Bakker et al.
-[27] explore threshold tuning to maintain stable classification performance.
-Building on these foundations, our framework FairOPT introduces a novel group-adaptive thresholding algorithm that jointly optimizes performance and fairness objectives.
+[25] present threshold-adaptation methods to ensure fair classification, and [2] explore threshold tuning to maintain stable classification performance.
+Building on these foundations, our framework FairOPT introduces a novel group‐adaptive thresholding algorithm that jointly optimizes performance and fairness objectives.
+III Conceptual Framework and Methodology
+Figure 2: FairOPT decreases discrepancy of AI detectors compared to other thresholding methods.
+FairOPT reduces classification discrepancy to $\mu=0.613$, outperforming ROCFPR ($\mu=0.920$) and Static ($\mu=0.887$).
+The error bar shows the standard deviation from the mean.
+Each bar shows average of five types of discrepancies that has been introduced in III-A4.
 Figure 3: Human-written short texts have a higher FPR compared to other text lengths, given the static threshold, showing the inadequacy of the static threshold to variations.
 Curves reflect the average probability of AI detectors in the training dataset.
-3 Conceptual Framework and Methodology
+TABLE I: FairOPTsignificantly reduces subgroup discrepancies across all evaluation metrics while keeping accuracy on par with baselines.
+This table compares thresholding methods: FairOPT, ROCFPR, and Static on ACC and five fairness metrics that measure discrepancies.
+Metric FairOPT ROCFPR Static
+ACC 0.6341 0.6249 0.6376
+MaxBER Discrepancy 0.5852 0.7592 0.7507
+Demographic Parity Difference 0.6772 0.9826 0.9730
+Equalized Odds Difference 0.6379 0.9198 0.8501
+FPR Discrepancy 0.4313 0.9715 0.9471
+Predictive Parity Difference 0.7333 0.9684 0.9130
+TABLE II: FairOPT achieves the lowest subgroup discrepancies among three thresholding methods while maintaining accuracy.
+Comparison of thresholding methods has been made across three datasets: MAGE (MG), RAID (RD), and SemEval (SE) and their overall averages (All).
+MaxBER = BER Discrepancy, MaxDP = Demographic Parity Discrepancy, MaxEO = Equalized Odds Discrepancy, MaxFPR = False Positive Rate Discrepancy, MaxPP = Predictive Parity Discrepancy.
+The discrepancy based on the MiniMax problem has been noted in III-A4.
+Method MaxBER Disc. MaxDP Disc. MaxEO Disc.
+All MG RD SE All MG RD SE All MG RD SE
+Static 0.751 0.730 0.771 0.754 0.973 0.972 0.958 0.987 0.850 0.859 0.838 0.852
+ROCFPR 0.759 0.745 0.783 0.753 0.983 0.978 0.977 0.992 0.920 0.949 0.901 0.906
+FairOPT 0.585 0.582 0.582 0.591 0.677 0.671 0.659 0.700 0.638 0.635 0.629 0.649
+Method MaxFPR Disc. MaxPP Disc. Accuracy
+All MG RD SE All MG RD SE All MG RD SE
+Static 0.947 0.939 0.974 0.932 0.913 0.902 0.920 0.918 0.638 0.639 0.636 0.638
+ROCFPR 0.972 0.971 0.985 0.961 0.968 0.967 0.982 0.958 0.625 0.624 0.625 0.626
+FairOPT 0.431 0.444 0.416 0.432 0.733 0.746 0.713 0.738 0.634 0.633 0.628 0.640
 This section outlines our method for adaptive thresholding, FairOPT, which iteratively adjusts classification thresholds for different subgroups to reduce discrepancy.
-3.1 Objective
+III-A Objective
 Consider a set of subgroups $G_{1},\ldots,G_{i}$ with attributes (e.g., length and style).
 Acknowledging that texts with varying characteristics show variations in probability distributions, our objective is to find optimal thresholds $\{\theta(G_{1}),\ldots,\theta(G_{i})\}$ for binary classification.
-3.1.1 Preliminaries of AI Text Detection
-The detection of AI-generated text is commonly formulated as a probabilistic classification for a binary classification, where the input space $\mathcal{X}$ consists of the sequence of tokens derived from a given text, and the label space $\mathcal{Y}=\{0,1\}$ denotes the two distinct classes of 00 for human-written and $1$ for AI-generated.
-To address this task, existing work primarily utilizes a pretrained neural network as a probabilistic classifier $M_{\theta}:\mathcal{X}\rightarrow[0,1]$, which assigns a probability $p$ to the input reflecting the likelihood of it being AI-generated [16, 13].
-This probability is then converted into a predicted label by applying a threshold, typically set at $0.5$.
-Thus, the predicted label, AI-written, $\hat{Y}\in\{0,1\}$ is determined as follows:
+III-A1 Preliminaries of AI Text Detection
+[…]
 $\hat{Y}=\begin{cases}1,&\text{if }M_{\theta}(x)\geq\text{threshold},\\ 0,&\text{otherwise.}\end{cases}$ (1)
-3.1.2 Performance metrics
-After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Details about the confusion matrix and performance metrics are in Appendix B.1.
-3.1.3 Fairness metrics
-To quantify disparity, we use Demographic Parity (DP) [31] and Balanced Error Rate (BER) [32].
+III-A2 Performance metrics
+After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Details about the confusion matrix and performance metrics are in Appendix -A.
+III-A3 Minimax problem
+We quantify subgroup disparity using five complementary metrics—Demographic Parity (DP) [26], Balanced Error Rate (BER) [16], Equalized Odds, FPR, and Predictive Parity.
+For each metric $k$, let $m_{k,g}$ denote the value on subgroup $g$; we define the inter-group discrepancy $\Delta_{k}\!=\!\max_{g}m_{k,g}\;-\;\min_{g}m_{k,g}$ (lower is better).
+We then adopt a minimax objective,
+$\min_{\theta}\;\max_{k}\;\Delta_{k}(\theta),$
+because fairness guarantees should hold in the worst subgroup rather than only on average. Controlling $\max_{k}\Delta_{k}$ addresses tail risk and improves robustness.
+III-A4 Fairness metrics
 DP is a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$.
@@
 In applying DP, the perfect fairness is $\Delta_{\text{DP}}=0$, which means no difference in overall TP and TN between groups.
-In our subgroup-adaptive evaluation, we use BER to equally measure the type I and II errors, which are FP and FN, while including positive predictions from DP.
+In our subgroup‐adaptive evaluation, we use BER to equally measure the type I and II errors, which are FP and FN, while including positive predictions from DP.
 For each subgroup $G_{i}$ we compute its BERi to measure the discrepancy, and the lower discrepancy indicates more equitable classification across groups.
@@
 $\text{Discrepancy}=\max\mathrm{BER}_{i}\;-\;\min\mathrm{BER}_{i}$ (4)
-Details about fairness metrics are noted in Appendix B.2.
-Figure 4: The FairOPT algorithm enables us to obtain a balance in the tradeoff between performance and fairness, through a relaxed fairness criterion discussed in Section 3.2.
-Details about the tradeoff are in Appendix H.2.
-3.2 Relaxed Fairness
+Discrepancy has been measured based on minimax problem to minimize the maximum subgroup discrepancy (e.g., BER gap, FPR gap), thereby reducing the worst-case disparity.
+Equalized Odds (EO) requires that a classifier’s predictions are conditionally independent of sensitive attributes $S$ given the true label $Y$ [22].
+This means that both the TPR and FPR should be equal across groups.
+In practice, EO measures whether individuals from different subgroups are treated similarly in terms of both favorable outcomes (correctly identifying positives) and unfavorable outcomes (incorrectly labeling negatives as positives).
+The discrepancy is defined as the maximum gap across subgroups in either TPR or FPR:
+$\Delta_{\text{EO}}$ $=\max\Bigl(\max_{g}\text{TPR}-\min_{g}\text{TPR},\;\max_{g}\text{FPR}-\min_{g}\text{FPR}\Bigr).$ (5)
+Perfect fairness under EO is achieved when $\Delta_{\text{EO}}=0$, meaning that subgroup differences in both detection sensitivity and error rates are eliminated.
+False Positive Rate (FPR) Discrepancy focuses specifically on subgroup differences in the rate of false alarms, i.e., when the classifier predicts $\hat{Y}=1$ despite the true label being $Y=0$.
+The discrepancy is formally defined as the difference between the maximum and minimum false positive rates across groups:
+$\Delta_{\text{FPR}}$ $=\max_{g}\text{FPR}_{g}-\min_{g}\text{FPR}_{g}.$ (6)
+This metric captures whether some groups are disproportionately subjected to incorrect positive predictions, which is critical in high-stakes applications such as fraud detection.
+Predictive Parity (PP) Discrepancy measures whether a positive prediction carries the same meaning for different groups, i.e., whether $P(Y=1\mid\hat{Y}=1,S=g)$ is group-independent.
+The discrepancy is defined as:
+$\Delta_{\text{PP}}$ $=\max_{g}\text{PP}_{g}-\min_{g}\text{PP}_{g},$ (7)
+where $\text{PP}_{g}=\frac{\text{TP}_{g}}{\text{TP}_{g}+\text{FP}_{g}}$.
+PP evaluates whether the precision of a classifier, defined as the proportion of true positives among all predicted positives, is consistent across subgroups.
+Addressing the mini–max problem in fairness is important because it explicitly targets the worst-case disparity across subgroups rather than optimizing only for average performance.
+Details about fairness metrics are noted in Appendix -B.
+Figure 4: The FairOPT algorithm enables us to obtain a balance in the tradeoff between performance and fairness, through a relaxed fairness criterion discussed in Section III-B.
+Details about the tradeoff are in Appendix -H.
+III-B Relaxed Fairness
 While achieving $100\%$ fairness across groups (e.g., $DP=0$ across fifteen groups) is theoretically ideal, it is often impractical due to computational constraints and the difficulty of ensuring model convergence under strict fairness criteria [13].
@@
 Algorithm 1 FairOPT: Group-Specific Thresholding
-Algorithm 1 FairOPT: Group-Specific Thresholding
 1:Input: $D$ dataset, $S_{n}$ features, predicted probabilities $p$, performance targets ($\alpha,\beta$), fairness gap limit $\delta_{\text{fair}}$, max iterations, step size $\eta$.
@@
 5:   Convert $p$ in $G_{i}$ to binary predictions using $\theta(G_{i})$ and compute confusion matrix.
-[…]
+6:   if ($\mathrm{ACC}_{i}\geq\alpha$ and $\mathrm{F1}_{i}\geq\beta$ for all ii) and (fairness gap $\leq\delta_{\text{fair}}$) then
 7:    return $\{\theta(G_{1}),\ldots,\theta(G_{j})\}$.
@@
 13:Output: $\{\theta(G_{1}),\dots,\theta(G_{j})\}$.
-3.3 Group-Adaptive Threshold Optimization
+III-C Group-Adaptive Threshold Optimization
 Process
@@
 $L_{i}(\theta)\;=\;-\,\mathrm{ACC}_{i}(\theta)\;+\;\kappa\,\bigl[\beta-\mathrm{F1}_{i}(\theta)\bigr]_{+},$ (9)
-where $\beta$ is a minimum F1, $\kappa>0$ is a penalty weight, $\alpha$ is given for the minimum ACC, and $[\cdot]_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero.
+where $\beta$ is a minimum F1, $\kappa>0$ is a penalty weight, $\alpha$ is given for the minimum ACC, and $\cdot$$_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero.
 These minimum standards are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
@@
 $\Delta_{\theta}\;=\;\max_{G_{i}}\,\bigl|\theta_{\mathrm{new}}(G_{i})-\theta_{\mathrm{old}}(G_{i})\bigr|.$ (13)
-[…]
+If $\Delta_{\theta}$ falls below a tolerance $\epsilon_{\mathrm{tol}}$ and all groups meet the performance criterion ($\mathrm{ACC}_{i}\geq\alpha$, $\mathrm{F1}_{i}\geq\beta$) and fairness disparities ($\Delta_{k}\leq\delta_{\mathrm{fair}}$) hold, the algorithm terminates.
 Outcome
 By iterating this process, the Algorithm 1 converges to an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that reduces the disparity while satisfying the demanded performance criteria.
-For full implementation details and hyperparameter specifications, refer to Appendix H.1.
+For full implementation details and hyperparameter specifications, refer to Appendix -G.
 Figure 5: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoBERTa-large detector results in cumulative probability distribution function (CDF; y-axis) based on the training dataset in the range of 0.1 to 0.9 (x-axis).
@@
 Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions by stylistic features.
-Figure 6: Threshold optimization based on the AUROC-based method under an $FPR\leq 10\%$ constraint (see Appendix C).
-The ROC curve (blue line) illustrates the trade-off between TPR and FPR, and the optimal threshold (green marker) maximizes TPR under the specified FPR constraint.
-4 Experiments
-4.1 Dataset
+Figure 6: Threshold optimization based on the AUROC‐based method under an $FPR\leq 10\%$ constraint (see Appendix AUROC).
+The ROC curve (blue line) illustrates the trade‐off between TPR and FPR, and the optimal threshold (green marker) maximizes TPR under the specified FPR constraint.
+IV Experiment
+IV-A Dataset
 Our experiments aim to assess the performance of FairOPT in comparison to the static and the AUROC-based threshold optimization methods The dataset is constructed with the RAID [12], MAGE [28], and SemEval24 [45] with an even balance between human and AI text to avoid class imbalance.
@@
 This experiment employed 19,117 instances for the training and 4,779 samples for the testing.
-We conducted the featured engineering noted in 4.3 and AI detections noted in 4.2 to extract features and get inferences from classifiers.
-4.2 AI-Generated Text Classifiers
+We conducted the featured engineering noted in IV-C and AI detections noted in IV-B to extract features and get inferences from classifiers.
+IV-B AI-Generated Text Classifiers
 To classify AI-generated text, we employed multiple AI text detectors, DeskLib [10], E5LoRA [50], FastDetectGPT [3], RADAR [24], RoBERTa-base and -large [42], and SimpleAI’s detector [19] for open-source availability.
@@
 These detectors were selected based on their common application and used in both the train and test datasets.
+They are based on probabilistic inference framework that needs a threshold for the final classification.
 The closed-source text detection is only conducted in the test dataset.
-Figure 7: FairOPT decreases the Balanced Error Rate (BER) discrepancy of AI detectors compared to other thresholding methods.
-FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$).
-The error bar shows the standard deviation from the mean.
-4.3 Subgroup Attributes
+IV-C Subgroup Attributes
 We extracted attributes from each text for sub-grouping to capture heterogeneity in text characteristics.
-These attributes are shown in Table 1.
+These attributes are shown in Table III.
 We categorized text length into three classes based on the number of characters as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500).
 We also labeled text with stylistic personality based on five traits by selecting the one with the highest probability among the five using a BERT-based personality prediction model [34].
-Details are noted in Appendix F.
+The third feature is formality that measures the formal expression of the sentence based on Heylighen & Dewaele’s formality score [23] and two categories (formal and informal) are labeled.
+The fourth feature is a binary sentiment label that has been inferred using DistilBERT-based SST-2 classifier[40].
+Details are noted in Appendix Feature Engineering.
 These features were chosen because text length captures structural variation and stylistic personality reflects writing tendencies, both of which critically influence detector stability.
+TABLE III: Subgroup attributes from the feature engineering.
 Attribute Characteristics Labels
@@
 Stylistic traits Inferred Big Five personality traits. Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism (k=5)
-Table 1: Subgroup attributes from the feature engineering.
+Formality Formality of the writing Formal, informal (k=2)
+Sentiment Binary sentiment Positive, negative (k=2)
 Figure 7: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness rule that is shown in Figure 4.
-Our method incurs minimal loss of ACC, achieving $\mu=0.6344\pm 0.1518$ compared to Static ($\mu=0.6376\pm 0.1975$) and ROCFPR ($\mu=0.6249\pm 0.2057$) methods.
-4.4 Discovery of Subgroup Bias in Text Detection
-We assess detection performance across text-length subgroups under a static threshold $\theta=0.5$.
-As illustrated in Figure 3, short human-authored samples incur a substantially higher FPR than mid-length and long texts.
+Our method incurs minimal loss of ACC, achieving $\mu=0.6341$ compared to Static ($\mu=0.6249$) and ROCFPR ($\mu=0.6376$) methods.
+IV-D Discovery of Subgroup Bias in Text Detection
+We assess detection performance across text‐length subgroups under a static threshold $\theta=0.5$.
+As illustrated in Appendix Detailed Results, short human‐authored samples incur a substantially higher FPR than mid‐length and long texts.
 This length-dependent disparity underscores the necessity of text length adaptive thresholding to mitigate the discrepancy.
 Next, we examine the combined effects of length and stylistic personality (Figure 5).
-By comparing the empirical cumulative distribution functions (CDFs) of AI-likelihood scores for short, conscientious versus long, neurotic texts, a two-sample Kolmogorov–Smirnov (KS) test reveals a statistically significant divergence in their score distributions ($p<0.05$), indicating that static thresholding disproportionately misclassifies certain stylistic and length combinations (see Appendix G.1.1 for full KS-test details).
+By comparing the empirical cumulative distribution functions (CDFs) of AI‐likelihood scores for short, conscientious versus long, neurotic texts, a two‐sample Kolmogorov–Smirnov (KS) test reveals a statistically significant divergence in their score distributions ($p<0.05$), indicating that static thresholding disproportionately misclassifies certain stylistic and length combinations (see Appendix -F1 for full KS‐test details).
 The observed disparity in the combination of text length and stylistic personality further highlights the need for adaptive thresholding to guarantee equitable detection performance across heterogeneous text cohorts.
-4.5 Mitigating Subgroup Bias using FairOPT
+IV-E Mitigating Subgroup Bias using FairOPT
 Given these findings, we now describe the experimental setup for applying FairOPT for AI-text detection with reduced discrepancy across subgroups.
-We extracted the text length and personality features from every instance in the dataset, partitioning the corpus into 15 subgroups.
-For the dataset, multiple AI text classifiers were employed to generate predicted probabilities indicating the probability that a given text was AI-generated with $p\in[0,1]$.
+We extracted the textual features from every instance in the dataset, partitioning the corpus into 60 subgroups.
+Across each dataset, multiple AI text classifiers were used to infer the probability $p\in[0,1]$ that a given text was AI-generated.
 We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup and applied them as noted in Figure 9.
-We include all hyperparameters and training details in Appendix H.3.
+We include all hyperparameters and training details in Appendix -I.
 Threshold optimization was performed on the training dataset based on the results from RoBERTa-large [42] rather than averaging the probabilities of multiple detectors.
 The relaxed fairness criterion is integrated to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness [44].
-We also applied an AUROC-based thresholding method using the RoBERTa-large [16] to select the decision threshold that maximizes the TPR under the constraint $FPR<0.1$; see Figure 6 for the AUROC-based optimization procedure [26].
+We also applied an AUROC‐based thresholding method using the RoBERTa‐large [42] to select the decision threshold that maximizes the TPR under the constraint $FPR<0.1$; see Figure 6 for the AUROC‐based optimization procedure [36].
 The optimized thresholds were then applied to the test dataset and evaluated based on discrepancy, ACC, and F1 score.
-Metrics are noted in Appendix B.2.
-4.6 Tradeoff Analysis between Performance and Fairness
+Metrics are noted in Appendix -B.
+IV-F Tradeoff Analysis between Performance and Fairness
 Figure 4 shows the performance-fairness tradeoff by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines.
-The relaxed fairness criterion from Section 3.2 to less sacrificing ACC while still satisfying a demanded level of fairness.
+The relaxed fairness criterion from Section III-B to less sacrificing ACC while still satisfying a demanded level of fairness.
 The ROCFPR method, shown in Figure 6, demonstrates a notable ability to minimize the FPR in training.
@@
 This trade-off is particularly problematic as it diminishes the detector’s reliability in correctly identifying AI-generated text, although the overall ACC is not overly sacrificed as shown in Figure 7.
-4.7 Results with FairOPT for AI-Text Detection
-Figure 7 demonstrates that group-adaptive thresholding from FairOPT consistently lowers the average discrepancy.
-Our method decreased the discrepancy to $0.4341$ ($\sigma=0.1927$), showing $12\%$ of reduction compared to the static-threshold baseline ($\mu=0.5512$, $\sigma=0.1758$) and outperforms the AUROC-tuned method ($\mu=0.5595$, $\sigma=0.1582$).
-Notably, E5LoRA shows the greatest decrease, with its BER discrepancy dropping by nearly half (from $0.63$ to $0.33$), and GPT4o-mini achieved a substantial reduction (from $0.60$ to $0.37$).
-The Desklib experiences a modest 0.003 reduction.
-Overall, the FairOPT decreased the discrepancy in all AI text classifiers.
-The detailed results with standard deviation by the dataset, detector, and thresholding methods are in Appendix LABEL:tab:overall_by_detector_and_dataset.
-Figure 8 illustrates that FairOPT incurs only a 0.003 of sacrifice in overall ACC and reduction of the standard deviation by 0.04 while outperforming the AUROC-based method.
-These findings demonstrate that FairOPT substantially mitigates fairness disparities with negligible ACC tradeoff, offering a practical application.
-Notably, our method shows great effectiveness to commonly used AI detectors like DeskLib [39], FastDetectGPT [41], RADAR [13], and SimpleAI’s detector [42] by securing ACC over $75\%$ while reducing the discrepancy in all cases.
-Moreover, the E5LoRA detector [40] exemplifies the ability of FairOPT to transcend the fairness–performance tradeoff.
-On the SemEval24 dataset, FairOPT achieves a 0.43 reduction in discrepancy while increasing ACC by 0.23.
-Similarly, on the MAGE, it delivers a 0.25 decrease in discrepancy coupled with a 0.21 ACC increase.
-These outcomes highlight the broad applicability of adaptive thresholding for improving both fairness and performance across diverse AI-text detection models, as well as other kinds of probabilistic classifiers.
-5 Conclusion
+IV-G Results with FairOPT for AI-Text Detection
+Table I shows that group-adaptive thresholding with FairOPT consistently reduces subgroup discrepancy across detectors.
+Overall, our method decreased the average discrepancy around $27.4\%$, substantially reducing subgroup gaps compared to static and ROCFPR baselines.
+On average, FairOPT achieves a BER discrepancy of $0.5852$, representing a $22\%$ reduction relative to the static threshold baseline ($0.7507$) and a $23\%$ reduction compared to ROCFPR ($0.7592$).
+Similar improvements are observed across other metrics: demographic parity decreases from $0.9730$ to $0.6772$ ($30\%$), equalized odds from $0.8501$ to $0.6379$ ($25\%$), false-positive rate discrepancy from $0.9471$ to $0.4313$ ($51\%$), and predictive parity from $0.9130$ to $0.7333$ ($18\%$).
+Overall, FairOPT lowers the discrepancy.
+Full results, including means and standard deviations by dataset, detector, and thresholding method, are provided in Appendix LABEL:tab:overall_by_detector_and_dataset.
+Regarding the consistency of FairOPT across datasets, it successfully reduces discrepancy across all fairness metrics compared to both Static and ROCFPR, as shown in Table II.
+On average, the MaxBER discrepancy decreases by $22\%$, dropping from 0.751 (Static) and 0.759 (ROCFPR) to 0.585.
+Similarly, the MaxDP discrepancy decreases by $30\%$, from 0.973 (Static) and 0.983 (ROCFPR) to 0.677, while the MaxEO discrepancy decreases by $25\%$, from 0.850 (Static) and 0.920 (ROCFPR) to 0.638.
+The strongest improvement is observed for MaxFPR, where FairOPT achieves a $55\%$ reduction, falling from 0.947 (Static) and 0.972 (ROCFPR) to 0.431.
+Predictive Parity discrepancy also improves by $20\%$, decreasing from 0.913 (Static) and 0.968 (ROCFPR) to 0.733.
+Importantly, these fairness gains come with negligible accuracy cost: overall accuracy remains stable at 0.634 for FairOPT, essentially unchanged from 0.638 under Static and higher than 0.625 under ROCFPR.
+Across detectors, FairOPT consistently reduces the average discrepancy as shown in 2 while making a negligible sacrifice of ACC 7.
+Across detectors, FairOPT lowers the mean discrepancy to 0.613 versus 0.887 (static) and 0.920 (ROCFPR).
+E5LoRA shows the largest gain, dropping to 0.100 from 0.672 (static) and 0.878 (ROCFPR).
+Among BERT-based classifiers, RoBERTa-base falls to 0.782 from 0.927 (static) and 0.932 (ROCFPR), while RoBERTa-large drops to 0.490 from 0.890 (static) and 0.940 (ROCFPR), both indicating substantial reductions.
+GPT-based detectors also improve meaningfully: GPT4o-mini to 0.253 and GPT4.1-nano to 0.466, compared to 0.9 under both static and AUROC.
+Other models show smaller but positive gains—DeskLib 12%, FastDetectGPT 10%, and SimpleAI 5%.
+The lone exception is RADAR, which increases by 2% under FairOPT; here, AUROC is approximately 1% lower than static.
+Detailed results with standard deviations are included in Appendix LABEL:tab:overall_by_detector_and_dataset.
+Figure 7 shows that the fairness improvements achieved by FairOPT come with only negligible predictive cost.
+Averaged across all datasets, FairOPT achieves an accuracy of $0.6341$, compared to $0.6376$ for Static and $0.6249$ for ROCFPR.
+The difference compared to static is less than $0.1\%$.
+These findings confirm that FairOPT substantially mitigates fairness disparities with negligible accuracy tradeoff, making it a practical approach for fair AI-text detection.
+IV-H Threshold Stability
+The stability of FairOPT were systematically evaluated across three highly diverse and representative benchmark datasets: MAGE, RAID, and SemEval.
+These datasets were intentionally selected to capture a broad spectrum of domains, genres, and writing styles, including academic and encyclopedic texts, social media posts, creative writing, and reviews.
+This diversity ensures that our subgroup definitions, based on features such as length, personality, formality, and sentiment, are well supported and meaningful within each dataset.
+This provides a rigorous testbed for threshold optimization.
+V Conclusion
 This paper proposes FairOPT, a post-processing algorithm for probabilistic classifiers that reduces the discrepancy while preserving overall accuracy.
@@
 Unlike universal thresholding methods, which presume evenly distributed probability regardless of the stylistic features of the text, our method effectively captures the variations of the probability distribution of AI text classifiers across different subgroups, leading to classification with a lower discrepancy.
+The novelty of this work resides within introducing a group-adaptive, post-hoc threshold optimization framework that operates without retraining of the detectors in the given settings by leveraging fairness and performance metrics.
+To best of our knowledge, no prior research has systematically examined threshold optimization in the context of AI text detectors although some research investigated thresholding methods for probabilistic classifiers.
+FairOPT represents a significant step toward AI safety by providing an alternative to the common practice of applying a single fixed threshold (e.g., 0.5, 0.75 or 0.97151).
 While our work focuses primarily on text classifiers, the FairOPT framework can naturally extend to other modalities where detecting AI-generated content is critical for safety and copyright (e.g., image, video, and audio), including deepfake video and audio detectors.
@@
 Finally, other approaches for threshold optimization, including Bayesian methods and convex optimization, can also be studied in future research.
-Although our experiments focus on text-based detectors, the FairOPT framework is modality-agnostic and readily extends to other domains such as image, video, and audio deepfake detection, where probabilistic outputs require post-processing to ensure equitable performance.
-More generally, FairOPT applies to any probabilistic classification task with subgroup disparities (e.g., recidivism risk prediction, hiring decisions) by learning tailored decision thresholds that mitigate bias with minimal impact on overall accuracy.
-6 Limitations
-While FairOPT demonstrably reduces subgroup discrepancy on our English-text benchmarks, it relies on only two features (character-count bins and model-inferred Big Five traits).
+VI Limitations
+While FairOPT demonstrably reduces subgroup discrepancy on our English‐text benchmarks, it relies on only two features (character‐count bins and model‐inferred Big Five traits).
 Other text attributes (e.g., topic, syntax, language) may induce additional biases that our current feature set does not capture.
-Moreover, the stylistic labels are derived from a downstream BERT-based personality predictor, which itself may introduce cultural bias.
+Moreover, the stylistic labels are derived from a downstream BERT‐based personality predictor, which itself may introduce cultural bias.
 We focus on relaxing demographic parity—extensions to other fairness definitions (equal opportunity or equalized odds) may require different optimization objectives or could be fundamentally incompatible.
-Impact Statement
+VII Future Works
+Although our experiments focused on English-language AI-generated text detectors, FairOPT is a modality-agnostic, subgroup-specific thresholding framework that readily extends to new deployment contexts and populations: for previously unseen subgroups, thresholds can be recalibrated from available data or warm-started from similar existing groups (e.g., AI-written text in schools, AI-generated images on a platform, or deepfakes on YouTube).
+Beyond text, FairOPT applies to image, video, and audio deepfake detection—any setting with probabilistic outputs that benefits from post-processing to ensure equitable performance—and more broadly to probabilistic classification tasks with subgroup disparities (e.g., recidivism risk prediction, hiring decisions), where learning tailored decision thresholds can mitigate bias with minimal impact on overall accuracy.
+We recognize that both theoretical analysis and user-centered case studies will be important for validating these extensions across diverse, real-world deployments.
+VIII Impact Statement
 This paper introduces novel theoretical frameworks and methodologies to enhance the robustness of AI-generated text detection systems.
@@
 Since it is crucial for these AI-content detection methods to be robust and fair across many potential users, our method takes a major step in this direction by formulating the problem setting and developing a new algorithm.
-Appendix A Technical Appendices and Supplementary Material
-Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline.
-There is no page limit for the technical appendices.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3.
-Answer: [Yes] We introduced a novel method that reduces the discrepancy of the probabilistic classifiers, especially in AI text detectors, and the main claim is noted both in the abstract and introduction.
-  4. Justification: We noted the main claims in the abstract and the main body of the paper and provided experimental results as evidence to support the main claim.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes] Limitations are addressed at the end of the paper.
-  9. Justification: Limitations about the AI text detection methods and questions around the inference to other kinds of probabilistic classifiers are addressed.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes] The paper provides the conceptual framework and the computational process
-  14.
-Justification: We provides the confusion matrix, employed performance and fairness metrics, fairness criterion, multiple thresholding methods, and the optimization process based on the finite-difference approximation.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes] The paper includes the URL to the dataset and codes that are used for the main results.
-  19.
-Justification: The paper provides the code and dataset.
-The experiment can be simulated with the provided dataset, and other experiments can also be conducted if they are based on the probabilistic classifiers and split into subgroups by adjusting the column names or adjusting hyperparameters.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes] We disclose all codes and dataset.
-  24. Justification: You can access the data and code through the provided technical materials or via the URL.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes] We disclose the dataset balance, split, and characteristics of data. The employed hyperparameters are included in the appendix. The optimizer is developed in this research.
-  29. Justification: The details about the dataset construction is noted in the dataset part, and appendix includes the hyperparameters.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33.
-Answer: [Yes] We report the error bars, which implies the standard deviation of the results.
-Also, the statistical significance of the probability distribution is addressed using the K-S test in the main body of the paper.
-  34.
-Justification: The reported error bar shows the variance of the result and we conducted an experiment to figure out the p-value of the statistical significance of the KS probability distribution.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [No] This research is based on a local device (CPU) and does not use GPU
-  39. Justification: The model is designed to be able to run in any environment without the constraint of the hardware.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes] This research complies with the Code of Ethics
-  44.
-Justification: Since this research is not a human-subjects research, this project is exempt from the IRB investigation.
-This research does not include human-subject sensitive information, but they are augmented using feature engineering.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes] We have included potential impacts from this research in Section Impact Statement.
-  49.
-Justification: This research generate sensitive features.
-But the extracted features do not infringe privacy because it is generating labels for each writing based on statistical inferences.
-This could lead a negative impacts.
-Regarding the positive impacts, the decreased discrepancy of the classifiers has been discussed.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes] We used the clipping method to restrict the optimization of the probability in the range of 0 to 1
-  54.
-Justification: Since the results outside of the range of 0 and 1 could lead to an error in the system, we used clipping in the model to prevent the exceed of the bound in the gradient descent process.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes] We used multiple AI text classifiers and they are all cited and addressed. Some codes may need the API keys from specific platforms like OpenAI
-  59. Justification: The train and test datasets include multiple AI-generated text detectors. And the test dataset has two more detection results based on the open-source LLMs.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes] All resources used for the paper are provided in the URL
-  64. Justification: We disclose all resources that used for the research.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A] This research is not crowdsourcing research.
-  69.
-Justification: The given multiple types of writings (e.g., essay, twitter, etc.) are constructed through the cited datasets, and the main novelty of the dataset is the application of multiple AI text detectors.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A] This research is not human-subject research
-  74. Justification: Some texts are written by human, but the sources are authorized and publicly accessible datasets.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [Yes] This research used two language models for the AI text detection, which his not a conventional method despite there being research that shows its efficiency.
-  79. Justification: This research used GPT4.1 and GPT4 like other AI text classifiers. This method is explored in other research as discussed in the main part of the paper.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
-Appendix B Performance and fairness metrics
-B.1 Contingency table
-The model can make not only true positive (TP) and true negative (TN) cases but also false positive (FP) and false negative (FN):
-Classified as AI ($\hat{Y_{1}}$) Classified as non-AI ($\hat{Y_{0}}$)
-AI-generated ($Y_{1}$) TP: Correct identification FN: Failed to detect AI and misidentified as human
-Human-developed ($Y_{0}$) FP: Incorrectly marked human work as AI TN: Correctly identified by not classifying human work as AI
-Table 2: Confusion matrix for the four possible outcomes when an AI detection system classifies either AI-generated or human-developed content.
-B.2 Notation and Metrics
-Given the table B.1 that denotes the outcomes in a binary classification task, we can mathematically express the performance and fairness metrics like below:
-  • Accuracy (ACC).
-    $\mathrm{ACC}\;=\;\frac{\mathrm{TP}+\mathrm{TN}}{\mathrm{TP}+\mathrm{TN}+\mathrm{FP}+\mathrm{FN}}.$
-  • F1 Score. Given $\mathrm{Precision}$ and $\mathrm{Recall}$ below,
-[…]
-    F1 matters because it balances Precision and Recall into a single measure, making it especially useful when the class distribution is imbalanced or when both metrics are important.
-  • Precision.
-    $\mathrm{Precision}\;=\;\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}}$
-    Precision measures the fraction of predicted positives that are actually correct, focusing on how often the model is right when it predicts a positive.
-  • Recall.  (Recall is also referred to as TPR in some contexts.)
-    $\mathrm{Recall}\;=\;\mathrm{TPR}\;=\;\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}$
-    Recall measures the fraction of actual positives that are correctly identified by the model, capturing how many real positives are detected.
-  • Demographic Parity (DP). Suppose we have a protected attribute $A\in\{0,1\}$. Demographic Parity requires $P(\hat{Y}=1\mid A=0)=P(\hat{Y}=1\mid A=1)$, or
-    $P(\hat{Y}=1\mid A=0)\approx P(\hat{Y}=1\mid A=1)$. A common metric is the disparity:
-    $\Delta_{\mathrm{DP}}\;=\;\bigl{|}\,P(\hat{Y}=1\mid A=0)\;-\;P(\hat{Y}=1\mid A=1)\bigr{|}.$
-  • Equalized Odds (EO). Equalized Odds requires equal true positive rates and false positive rates across groups. One way to measure the gap is:
-    $\Delta_{\mathrm{EO}}\;=\;\bigl{|}P(\hat{Y}=1\mid Y=1,A=0)\;-\;P(\hat{Y}=1\mid Y=1,A=1)\bigr{|}\;+\;\ \$
-    $\bigl{|}P(\hat{Y}=1\mid Y=0,A=0)\;-\;P(\hat{Y}=1\mid Y=0,A=1)\bigr{|}.$
-    Under the strictest notion of fairness, one requires
-[…]
-    implying no disparity in true or false positive rates across groups.
-    Relaxation in fairness allows
-    $\Delta_{\mathrm{EO}}\geq 0,$
-    for feasibility.
-  • False Positive Rate (FPR).
-    $\mathrm{FPR}\;=\;\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}$
-  • Balanced Error Rate (BER).
-    $\mathrm{BER}\;=\;\frac{1}{2}\left(\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}+\frac{\mathrm{FN}}{\mathrm{TP}+\mathrm{FN}}\right)\;=\;\frac{\mathrm{FPR}+\mathrm{FNR}}{2}\;=\;\frac{\mathrm{FPR}+(\mathrm{1-TPR})}{2}$
-    Balanced Error Rate (BER) averages the error rates across both classes, providing a metric that is insensitive to class imbalance.
-By considering both the False Positive Rate and the False Negative Rate, BER ensures that the classifier performs equally well on both classes.
-The concept is especially well-introduced in [32]
-  • BER Discrepancy.
-    $\text{Discrepancy}=\max\mathrm{BER}_{i}\;-\;\min\mathrm{BER}_{i}$
-    For each subgroup $G_{i}$ we can compute BERi to measure the discrepancy by comparing the maximum BER and minimum BER, and the lower discrepancy indicates more equitable classification across groups.
-  • Specificity.
-    $\mathrm{Specificity}\;=\;\frac{\mathrm{TN}}{\mathrm{TN}+\mathrm{FP}}$
-    Specificity measures the proportion of actual negative cases correctly identified by the model, reflecting its ability to detect true negatives. It is also known as true negative rate (TNR).
-B.3 Impossibility theorem
-The impossibility theorem typically refers to the result that no single classifier can satisfy all common fairness metrics (e.g., demographic parity, equalized odds, predictive parity) simultaneously, except in trivial cases.
-This arises because different fairness definitions can be mathematically incompatible when the base rates (prevalences) differ across groups.
-This research used one fairness metrics in the processing to address this issue.
-Appendix C AUROC
-C.1 Static threshold and AUROC-based approach
-In this research, the AUROC-based approach is employed to optimize the classification threshold while balancing model performance and fairness constraints.
-This method evaluates the trade-off between the TPR and FPR, as visualized by the ROC curve shown in Figure 9.
-The optimal threshold is determined under a specified FPR constraint, $\text{FPR}\leq 0.1$, ensuring that the selected threshold aligns with a predefined ethical constraint.
-Figure 9: Threshold optimization process using the AUROC-based approach with an FPR constraint of 0.1.
-The ROC curve (blue line) demonstrates the trade-off between TPR and FPR.
-The optimal threshold (green marker) is selected to maximize performance under the specified FPR constraint.
-Appendix D Applied Thresholds
-Figure 2 uses thresholds that is noted on Table 3.
-Threshold Type Value
-Static 0.5
-ROCFPR 0.2866322
-Adaptive Thresholds (n=15)
-Group Threshold
-long_agreeableness 0.988102
-long_conscientiousness 0.983204
-long_extroversion 0.991482
-long_neuroticism 0.998293
-long_openness 0.978291
-medium_agreeableness 0.971192
-medium_conscientiousness 0.980804
-medium_extroversion 0.987482
-medium_neuroticism 0.994291
-medium_openness 0.969759
-short_agreeableness 0.986492
-short_conscientiousness 0.994242
-short_extroversion 0.978118
-short_neuroticism 0.994192
-short_openness 0.970664
-Table 3: Thresholds from three thresholding methods that used in the experiment.
-Appendix E Detailed Results
-We report a table showing the overall discrepancy, ACC, and F1 by dataset source: Table 4, and a table showing the detailed difference by detectors: Table LABEL:tab:overall_by_detector_and_dataset.
-E.1 Overall discrepancy and performance
-The table 4 summarizes the overall discrepancy and performance by each dataset and thresholding method. Detailed results that include the result from each detector is noted in the next section.
-Source Strategy Discrepancy (mean $\pm$ std) Accuracy (mean $\pm$ std) F1 Score (mean $\pm$ std)
-MAGE FairOPT $\mathbf{0.4495\pm 0.1656}$ $0.6339\pm 0.1512$ $0.3877\pm 0.3494$
-ROCFPR $0.5551\pm 0.1664$ $0.6238\pm 0.2113$ $0.5326\pm 0.3170$
-Static $0.5433\pm 0.1781$ $0.6388\pm 0.2015$ $0.5126\pm 0.3352$
-RAID FairOPT $\mathbf{0.4836\pm 0.1781}$ $0.6280\pm 0.1573$ $0.3885\pm 0.3535$
-ROCFPR $0.5644\pm 0.1263$ $0.6253\pm 0.2082$ $0.5363\pm 0.3185$
-Static $0.5562\pm 0.1478$ $0.6364\pm 0.1994$ $0.5136\pm 0.3358$
-SemEval FairOPT $\mathbf{0.3763\pm 0.2166}$ $0.6404\pm 0.1497$ $0.3896\pm 0.3484$
-ROCFPR $0.5598\pm 0.1758$ $0.6258\pm 0.2010$ $0.5303\pm 0.3114$
-Static $0.5548\pm 0.1969$ $0.6376\pm 0.1948$ $0.5081\pm 0.3288$
-Table 4: Overall discrepancy and performance by data source and thresholding strategy (mean $\pm$ std)
-E.2 Comparison by performance and fairness metrics by datasource, detector, and thresholding strategy
-The table LABEL:tab:overall_by_detector_and_dataset shows the results by each detector by the datasource and thresholding method based on fairness and performance metrics.
-Table 5: Discrepancy and performance (mean $\pm$ std) by data source, detector, and thresholding method
-Dataset
-Detector Thresholding $Discrepancy$ $ACC$ $F1score$
-MAGE
-GPT4.1-nano (2025) FairOPT $0.4106\pm 0.2350$ $0.4794\pm 0.0374$ $0.0114\pm 0.0121$
-ROCFPR $0.4636\pm 0.2194$ $0.4274\pm 0.0380$ $0.0694\pm 0.0159$
-Static $0.4647\pm 0.2143$ $0.4393\pm 0.0401$ $0.0507\pm 0.0154$
-GPT4o-mini (2024) FairOPT $0.3820\pm 0.2421$ $0.4866\pm 0.0345$ $0.0046\pm 0.0078$
-ROCFPR $0.6874\pm 0.1882$ $0.4453\pm 0.0288$ $0.2348\pm 0.0399$
-Static $0.6517\pm 0.1661$ $0.4423\pm 0.0253$ $0.1947\pm 0.0304$
-SimpleAI (2023) FairOPT $0.4536\pm 0.0871$ $0.7501\pm 0.0274$ $0.7006\pm 0.0363$
-ROCFPR $0.5682\pm 0.0945$ $0.7573\pm 0.0130$ $0.7520\pm 0.0215$
-Static $0.5527\pm 0.0968$ $0.7525\pm 0.0154$ $0.7360\pm 0.0298$
-DeskLib (2024) FairOPT $0.4896\pm 0.0914$ $0.8900\pm 0.0170$ $0.8809\pm 0.0196$
-ROCFPR $0.4854\pm 0.2388$ $0.8787\pm 0.0212$ $0.8916\pm 0.0224$
-Static $0.4278\pm 0.2561$ $0.9074\pm 0.0206$ $0.9140\pm 0.0219$
-E5LoRA (2024) FairOPT $0.3571\pm 0.2440$ $0.4895\pm 0.0336$ $00.0000\pm 0.0$ 0.0000
-ROCFPR $0.5779\pm 0.1866$ $0.1996\pm 0.0302$ $0.0471\pm 0.0179$
-Static $0.6026\pm 0.2439$ $0.2797\pm 0.0237$ $0.0065\pm 0.0090$
-FastDetectGPT (2023) FairOPT $0.4929\pm 0.0189$ $0.8099\pm 0.0274$ $0.7774\pm 0.0243$
-ROCFPR $0.4881\pm 0.0315$ $0.7776\pm 0.0302$ $0.8114\pm 0.0288$
-Static $0.4911\pm 0.1295$ $0.8398\pm 0.0154$ $0.8520\pm 0.0151$
-RoBERTa-base (2019) FairOPT $0.4893\pm 0.0537$ $0.6742\pm 0.0436$ $0.5381\pm 0.0655$
-ROCFPR $0.5923\pm 0.0884$ $0.7322\pm 0.0311$ $0.6963\pm 0.0320$
-Static $0.6158\pm 0.1606$ $0.7256\pm 0.0364$ $0.6722\pm 0.0423$
-RoBERTa-large (2019) FairOPT $0.3940\pm 0.1992$ $0.5063\pm 0.0364$ $0.0656\pm 0.0372$
-ROCFPR $0.5835\pm 0.0610$ $0.6264\pm 0.0289$ $0.5067\pm 0.0405$
-Static $0.5414\pm 0.1038$ $0.6103\pm 0.0361$ $0.4315\pm 0.0522$
-RADAR (2023) FairOPT $0.5760\pm 0.0755$ $0.6192\pm 0.0216$ $0.5106\pm 0.0388$
-ROCFPR $0.5495\pm 0.2113$ $0.7693\pm 0.0186$ $0.7837\pm 0.0202$
-Static $0.5418\pm 0.1280$ $0.7519\pm 0.0224$ $0.7562\pm 0.0242$
-RAID
-GPT4.1-nano (2025) FairOPT $0.5208\pm 0.2082$ $0.4738\pm 0.0087$ $0.0128\pm 0.0115$
-ROCFPR $0.5820\pm 0.1623$ $0.4243\pm 0.0189$ $0.0795\pm 0.0295$
-Static $0.5781\pm 0.1558$ $0.4306\pm 0.0183$ $0.0465\pm 0.0202$
-GPT4o-mini (2024) FairOPT $0.4781\pm 0.2428$ $0.4710\pm 0.0180$ $00.0000\pm 0.0$ 0.0000
-ROCFPR $0.5762\pm 0.1567$ $0.4557\pm 0.0266$ $0.2292\pm 0.0542$
-Static $0.5816\pm 0.1733$ $0.4529\pm 0.0180$ $0.1875\pm 0.0321$
-SimpleAI (2023) FairOPT $0.5868\pm 0.2080$ $0.7320\pm 0.0408$ $0.6835\pm 0.0548$
-ROCFPR $0.5655\pm 0.1818$ $0.7452\pm 0.0386$ $0.7502\pm 0.0388$
-Static $0.6245\pm 0.2095$ $0.7403\pm 0.0516$ $0.7357\pm 0.0541$
-DeskLib (2024) FairOPT $0.3552\pm 0.1001$ $0.9023\pm 0.0289$ $0.8973\pm 0.0301$
-ROCFPR $0.4578\pm 0.0522$ $0.8667\pm 0.0223$ $0.8843\pm 0.0194$
-Static $0.4292\pm 0.0646$ $0.8988\pm 0.0212$ $0.9076\pm 0.0189$
-E5LoRA (2024) FairOPT $0.4167\pm 0.2041$ $0.4808\pm 0.0153$ $00.0000\pm 0.0$ 0.0000
-ROCFPR $0.6437\pm 0.1177$ $0.2059\pm 0.0384$ $0.0406\pm 0.0272$
-Static $0.6535\pm 0.1981$ $0.2805\pm 0.0323$ $0.0078\pm 0.0097$
-FastDetectGPT (2023) FairOPT $0.4405\pm 0.0560$ $0.8241\pm 0.0203$ $0.7996\pm 0.0230$
-ROCFPR $0.5103\pm 0.0873$ $0.7850\pm 0.0345$ $0.8191\pm 0.0312$
-Static $0.4609\pm 0.1409$ $0.8472\pm 0.0247$ $0.8601\pm 0.0228$
-RoBERTa-base (2019) FairOPT $0.5361\pm 0.0452$ $0.6706\pm 0.0343$ $0.5483\pm 0.0681$
-ROCFPR $0.6246\pm 0.1244$ $0.7348\pm 0.0355$ $0.7030\pm 0.0478$
-Static $0.5978\pm 0.0984$ $0.7180\pm 0.0342$ $0.6682\pm 0.0499$
-RoBERTa-large (2019) FairOPT $0.4282\pm 0.1927$ $0.4990\pm 0.0183$ $0.0698\pm 0.0285$
-ROCFPR $0.6079\pm 0.0816$ $0.6413\pm 0.0218$ $0.5341\pm 0.0394$
-Static $0.5570\pm 0.0658$ $0.6176\pm 0.0175$ $0.4579\pm 0.0319$
-RADAR (2023) FairOPT $0.5900\pm 0.1850$ $0.5980\pm 0.0179$ $0.4853\pm 0.0412$
-ROCFPR $0.5113\pm 0.0602$ $0.7683\pm 0.0263$ $0.7867\pm 0.0233$
-Static $0.5232\pm 0.0644$ $0.7418\pm 0.0265$ $0.7510\pm 0.0257$
-SemEval
-GPT4.1-nano (2025) FairOPT $0.2663\pm 0.2792$ $0.4889\pm 0.0498$ $0.0110\pm 0.0116$
-ROCFPR $0.5139\pm 0.2100$ $0.4459\pm 0.0354$ $0.0912\pm 0.0320$
-Static $0.4889\pm 0.2156$ $0.4495\pm 0.0408$ $0.0503\pm 0.0210$
-GPT4o-mini (2024) FairOPT $0.2763\pm 0.2923$ $0.4943\pm 0.0501$ $0.0029\pm 0.0076$
-ROCFPR $0.5777\pm 0.2485$ $0.4638\pm 0.0350$ $0.2381\pm 0.0484$
-Static $0.5617\pm 0.2843$ $0.4620\pm 0.0418$ $0.2008\pm 0.0364$
-SimpleAI (2023) FairOPT $0.5655\pm 0.2123$ $0.7382\pm 0.0369$ $0.6842\pm 0.0260$
-ROCFPR $0.6726\pm 0.1674$ $0.7430\pm 0.0313$ $0.7375\pm 0.0264$
-Static $0.6519\pm 0.1904$ $0.7436\pm 0.0327$ $0.7283\pm 0.0231$
-DeskLib (2024) FairOPT $0.3711\pm 0.0755$ $0.8942\pm 0.0349$ $0.8844\pm 0.0316$
-ROCFPR $0.4487\pm 0.0727$ $0.8643\pm 0.0169$ $0.8776\pm 0.0241$
-Static $0.3691\pm 0.0970$ $0.8954\pm 0.0139$ $0.9012\pm 0.0184$
-E5LoRA (2024) FairOPT $0.2143\pm 0.2673$ $0.5015\pm 0.0523$ $00.0000\pm 0.0$ 0.0000
-ROCFPR $0.6887\pm 0.2807$ $0.2098\pm 0.0405$ $0.0293\pm 0.0124$
-Static $0.6428\pm 0.2790$ $0.2750\pm 0.0402$ $0.0065\pm 0.0061$
-FastDetectGPT (2023) FairOPT $0.4818\pm 0.0312$ $0.8111\pm 0.0254$ $0.7701\pm 0.0298$
-ROCFPR $0.4689\pm 0.0646$ $0.7490\pm 0.0284$ $0.7851\pm 0.0345$
-Static $0.5232\pm 0.1360$ $0.8159\pm 0.0253$ $0.8268\pm 0.0276$
-RoBERTa-base (2019) FairOPT $0.4922\pm 0.0426$ $0.6940\pm 0.0371$ $0.5676\pm 0.0472$
-ROCFPR $0.5748\pm 0.1175$ $0.7519\pm 0.0241$ $0.7127\pm 0.0362$
-Static $0.5543\pm 0.0866$ $0.7412\pm 0.0263$ $0.6847\pm 0.0435$
-RoBERTa-large (2019) FairOPT $0.2956\pm 0.2093$ $0.5188\pm 0.0513$ $0.0718\pm 0.0086$
-ROCFPR $0.5977\pm 0.1239$ $0.6479\pm 0.0218$ $0.5280\pm 0.0376$
-Static $0.6073\pm 0.1346$ $0.6186\pm 0.0293$ $0.4333\pm 0.0302$
-RADAR (2023) FairOPT $0.4241\pm 0.1244$ $0.6222\pm 0.0358$ $0.5140\pm 0.0407$
-ROCFPR $0.4956\pm 0.0187$ $0.7561\pm 0.0310$ $0.7735\pm 0.0357$
-Static $0.5939\pm 0.1847$ $0.7370\pm 0.0300$ $0.7407\pm 0.0361$
-E.3 Results for Discrepancy, ACC, and F1
-The bar charts illustrate the discrepancy, ACC, and F1 are incorporated in the main body of the paper, and this section encompasses a chart representing the F1 score, accompanied by three figures.
-The sequence of figures is arranged in the order of discrepancy, ACC, and F1.
-Figure 10: FairOPT decreases the discrepancy, Balanced Error Rate (BER) discrepancy, of probabilistic classifiers compared to other thresholding methods.
-FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$).
-Figure 11: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness that is shown in Figure 4.
-Our method incurs only a negligible accuracy sacrifice, achieving $\mu=0.6344\pm 0.1518$ compared to Static $\mu=0.6376\pm 0.1975$ and ROCFPR $\mu=0.6249\pm 0.2057$.
-Figure 12: Overall F1 have been sacrificed in the application of FairOPT.
-FairOPT achieves $\mu=0.389\pm 0.350$, compared to ROCFPR $\mu=0.533\pm 0.316$ and Static $\mu=0.511\pm 0.333$.
-Detectors with low ACC show a very low F1 score.
-Appendix F Feature Engineering
-In order to mitigate disparities within AI classifiers, a thorough feature engineering process was undertaken.
-Features were systematically extracted to characterize each text within the dataset.
-The primary focus of the experiment encompassed text length and anticipated personality traits.
-Although formality and sentiment were also extracted, these attributes are utilized to expedite processing, as the computational time increases exponentially with the inclusion of additional features.
-The subsequent table illustrates the features primarily extracted for the experiment, with an emphasis solely on the utilization of length and personality traits.
-This research mainly used length and personality as a main feature in using FairOPT.
-The text length attribute measures the number of characters in a text.
-OpenAI’s evaluations [26] revealed performance disparities between short (under 1,000 characters) and long texts when using their AI classifier.
-While their global threshold approach failed to address group-specific differences, it provided a valuable benchmark.
-Building on this and using exploratory data analysis (EDA), we refined the thresholds to 1,000 and 2,500 characters, categorizing texts as short (up to 1,000), mid-length (1,000–2,500), or long (over 2,500).
-The personality attribute includes five traits: extroversion, neuroticism, agreeableness, conscientiousness, and openness.
-This attribute contributes to the enforcement of fairness metrics, such as demographic parity and equality of odds.
-For this feature, we determine the personality trait by selecting the one with the highest probability among the five using the [45] model, which is fine-tuned from BERT-base-uncased [47].
-Appendix G Discrepancies by classifiers with additional features
-G.1 Discrepancy of Probability in Subgroups
-G.1.1 Discrepancy in RoBERTa-large Detector
-The analysis revealed significant variability in the probability distributions between subgroup pairs, as depicted in Figure 13.
-Employing two-sample Kolmogorov-Smirnov (KS) statistic for two-sided tests, we assessed distributional discrepancies across various text groups characterized by distinct feature combinations, under the assumption of a non-parametric distribution.
-Significant differences were identified in certain subgroup comparisons, indicating notable discrepancies in their probability distributions.
-For example, comparing texts characterized by conscientiousness and short with those exhibiting neuroticism and long resulted in a KS statistic of 0.3081 ($p<0.01$).
-Similarly, evaluating conscientiousness and short against extroversion and medium yielded a KS statistic of 0.2616 ($p<0.01$).
-Conversely, smaller discrepancies were observed between agreeableness and short versus neuroticism and short (KS = 0.0909, $p<0.05$) and neuroticism and short versus openness and short (KS = 0.0842, $p<0.01$).
-Figure 13: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoB33ERTa-large detector results in cumulative probability distribution function (CDF) based on the training dataset in the range of 0.1 to 0.9.
-The biggest discrepancy is observed with a KS statistic of $0.3081$ ($p<0.01$), while the smallest is $0.0842$ ($p<0.01$).
-Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions
-Appendix H Operationalizing FairOPT
-H.1 Algorithmic description
-1:  Input: Labeled dataset $D=\{(x_{n},y_{n})\}$, predicted probabilities $\{p_{n}\}$, subgroup labels $\{g_{n}\}$ with possible groups $G_{1},\dots,G_{j}$, features $\{S_{1},\dots,S_{n}\}$, initial thresholds $\theta_{init}$, fairness metrics $\{M_{1},\ldots,M_{k}\}$ with acceptable disparity $\delta_{\mathrm{fair}}$, minimum performance thresholds: $\alpha$ (ACC) and $\beta$ (F1), penalty weight $\kappa$, learning rate $\eta$, tolerance $\epsilon_{\mathrm{tol}}$, maximum iterations, finite-difference step $\delta$, minimum probability for threshold $a$, maximum probability for threshold $b$
-2:  Output: Subgroup thresholds $\{\theta^{*}(G_{1}),\dots,\theta^{*}(G_{j})\}$
-3:  Initialization:
-4:  split $D$ to ${(G_{1}),\dots,(G_{j})}$ by $S$
-5:  while iteration $<$ maximum iterations do
-6:     for each group $G_{j}$ do
-7:        $\mathcal{I}_{i}\leftarrow\{\,n\mid g_{n}=G_{j}\}$  all samples in group $G_{j}$
-8:        $\hat{y}_{n}\leftarrow[\,p_{n}\geq\theta(G_{j})\,],\;\forall n\in\mathcal{I}_{i}$
-9:        Store contingency table for $G_{i}$ using $\hat{y}_{n}$ and ${y}_{n}$
-10:        $\Omega\;\leftarrow\;\min_{i}\{\mathrm{ACC}(G_{j})\}\;\geq\;\alpha\;\wedge\;\min_{i}\{\mathrm{F1}(G_{j})\}\;\geq\;\beta$ boolean output
-11:        for each fairness metric $M_{k}$ do
-12:           Compute $\{M_{k}(G_{1}),\dots,M_{k}(G_{j})\}$
-13:           $\Delta_{k}\leftarrow\max_{i}(M_{k}(G_{j}))-\min_{i}(M_{k}(G_{j}))$
-14:           $\Psi\;\leftarrow\;\Bigl{[}\,\max_{k}\Delta_{k}\;\leq\;\delta_{\mathrm{fair}}\Bigr{]}$ boolean output
-15:        end for
-16:        for each group $G_{j}$ do
-17:           $L_{i}(\theta)=-\mathrm{Acc}_{j}(\theta)+\kappa\,[\,\beta-\mathrm{F1}_{j}(\theta)\,]_{+}$
-18:
-$\nabla L_{i}(\theta(G_{j}))\;\approx\;\frac{L_{i}(\theta(G_{j})+\delta)\;-\;L_{i}(\theta(G_{j})-\delta)}{2\,\delta}.$
-19:
-$\theta(G_{j})\;\leftarrow\;\theta(G_{j})\;-\;\eta\;\nabla L_{i}(\theta(G_{j})).$
-20:           Clip $\theta(G_{j})$ to stay within $[a,b]$
-21:        end for
-22:     end for
-23:     $\Delta_{\theta}\;\leftarrow\;\max_{G_{j}}\bigl{|}\theta(G_{j})-\text{previousThresholds}[G_{j}]\bigr{|}$
-24:     if ($\Delta_{\theta}<\epsilon_{\mathrm{tol}}\;\wedge\;\Omega\;\wedge\;\Psi$then
-25:        break
-26:     end if
-[…]
-28:     $\text{iteration}\leftarrow\text{iteration}+1$
-29:  end while
-30:  return $\{\theta(G_{1}),\dots,\theta(G_{j})\}$
-Algorithm 2 FairOPT: Gradient-based adaptive threshold optimization with fairness for subgroups
-H.2 Early stopping method for the tradeoff curve
-The FairOPT algorithm was implemented using functions from the scipy.optimize library, aiming to determine probability thresholds that satisfy specific fairness constraints (acceptable_disparity = [1.00, 0.30, 0.25, 0.21, 0.20, 0.19, 0.15, 0.15]) while ensuring minimum performance thresholds for both performance criteria (ACC and F1).
-The optimization in the minimize function was performed using the L-BFGS-B method.
-The figure is based on the RoBERTa-base model, and it uses different hyperparameters to get eight results with different relaxation methods.
-Early stopping is facilitated by the default values of ftol and gtol in L-BFGS-B.
-Additionally, the learning rate was set to $10^{-3}$ to enable controlled and stable updates during optimization, and a penalty parameter was set to $10$ to strengthen the L2 penalty in the loss function.
-Furthermore, minimum performance thresholds for accuracy and F1-score were enforced based on the performance observed for each group when applying a static threshold of 0.5 and a ROC-based threshold.
-As stricter fairness constraints are imposed, these minimum thresholds are gradually relaxed.
-H.3 Hyperparameter settings for the FairOPT
-The application of the FairOPT is configured with a set of hyperparameters designed to ensure a balance between fairness and performance during the threshold optimization process.
-The learning rate is set to $10^{-3}$ to enable controlled and stable updates during optimization.
-The maximum number of iterations is specified as $10^{5}$ to provide sufficient time for the optimization algorithm to converge.
-An acceptable disparity of 0.2 is defined to regulate subgroup fairness, ensuring that performance differences across subgroups remain within an acceptable range through the relaxed fairness.
-Minimum thresholds for accuracy (0.25) and F1-score (0.25) are enforced to maintain baseline model performance and stability, particularly in cases of imbalanced datasets.
-A tolerance level of $1/2*10^{-4}$ is introduced to impose convergence criteria, ensuring the optimization process halts only when the changes in the objective function are negligible.
-Furthermore, a penalty parameter of 20 is applied to enforce performance constraints by penalizing significant performance disparities across subgroups.
-For fairness criteria, Demographic Parity (DP) was employed to test whether the generated thresholds meets the expected fairness.
-Also, this experiment used early stopping method for effective processing.
-The early stopping check through the change of the ACC, F1, and maximum DP discrepancy of each group, and terminates if all of them shows minimal change based on predefined patience.
+Acknowledgments
+Thanks to Hengzhi Li, Megan Tjandrasuwita, David Dai, Jeongmin Kwon, Chanakya Ekbote, and Seonghyun Nam for constructive feedback and discussion.